### v5

# D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts

Abstract
This paper achieves noteworthy progress in the realm of abstract reasoning, particularly in addressing Raven’s Progressive Matrices (RPM) and Bongard-Logo challenges. Initially, we introduce Lico-Net, a novel baseline model that resolves RPM problems with remarkable accuracy. Leveraging this foundation, we advance with the D3C approach, which advocates representing the underlying concepts in abstract reasoning problems through distributions. This perspective enhances the performance of both Lico-Net and a baseline model excelling in Bongard-Logo tasks. To bolster the computational efficiency of D3C, we present the D3C-cos variant, offering a streamlined yet precise solution. Furthermore, we propose the D2C method, redefining conceptual boundaries within these domains and bridging the divide between high-level abstractions and their lower-dimensional counterparts. Finally, we extend our methodology to D4C, employing adversarial techniques to refine conceptual boundaries further and demonstrate substantial improvements in both RPM and Bongard-Logo challenges. Overall, our contributions present a fresh outlook and practical advancements in the field of abstract reasoning.
Index Terms:
Abstract reasoning, RPM problem, Bongard-logo problem.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Deep neural networks have achieved remarkable success in various domains, including computer vision[1, 2, 3], natural language processing[4, 5, 6], generative models[7, 8, 9], visual question answering[10, 11], and abstract reasoning[14, 12, 13]. This success is primarily attributed to the remarkable capabilities of deep learning, which can recognize text, images, and sound data by learning the inherent patterns and representational hierarchies within sample data, thereby demonstrating analytical learning abilities similar to humans. Especially when addressing complex pattern recognition problems, the powerful capabilities of deep learning enable it to mimic human activities, further advancing the progress of artificial intelligence-related technologies. Its outstanding performance in speech and image recognition has surpassed previous technologies in terms of recognition accuracy.
However, despite these impressive accomplishments in deep learning, there are still a series of unresolved issues and challenges that require further research and exploration. For instance, research is ongoing on how to utilize the results of upperlayer training as initialization parameters for lowerlayer training, aiming to enhance the efficiency of deep model training. Additionally, layer-wise initialization training with unsupervised learning constitutes a significant strategy within the field of deep learning.
Especially in the domain of graphical abstract reasoning, the advancement of deep learning remains an active area of research. Although deep learning has achieved remarkable success in intelligent vision tasks, elevating machine intelligence to potentially higher levels, the academic community’s research on the abstract reasoning capabilities of deep learning has not ceased. Conversely, this has propelled the emergence of graphical reasoning problems. Graphical reasoning problems pose unique challenges for deep learning, including small and difficult-to-label datasets, the requirement for deep understanding of spatial relationships, shapes, and colors in graphics, demanding strong reasoning and innovation capabilities from models, necessitating a certain level of explanatory ability from models, and posing higher demands on models due to problem complexity and diversity.
To address these challenges, we need to further explore novel model architectures, training methodologies, and optimization strategies. This may encompass directions such as designing more effective deep learning models, refining model training methods, and enhancing dataset quality. Simultaneously, we also need to develop a deeper understanding of the essence and characteristics of graphical reasoning problems so that we can better leverage deep learning techniques to tackle these issues. In summary, while deep learning has made certain strides in graphical abstract reasoning, numerous problems still await our investigation and resolution.
The Ravens Progressive Matrices (RPM)[14] and Bongard problems[12, 13], for example, present learning requirements that span from perceptual processes to inferential capabilities.
I-A RAVEN and PGM
The RAVEN database[15], a distinctive collection of Ravens Progressive Matrices (RPM) problems[14], stands out for its particular design elements. Each RAVEN problem consists of sixteen images, with eight forming the essence of the problem, termed as the stem, while the remaining eight constitute a pool of alternate choices. Subjects are tasked with selecting appropriate images from this pool, which, together with the stem images, form a 3×3 matrix of graphical patterns. The matrix is characterized by a progressive pattern of geometric images in the row direction, encompassing an underlying abstract concept. Figure 1 presents a concrete instantiation of a RAVEN case’s construction, shedding light on the methodology and general attributes of these problems.
Figure 1: RAVEN case
Within the design of RAVEN problems, geometric images often encompass human-predefined concepts such as ”shape” and ”color”. However, these concepts are further abstracted in RAVEN problems into a finite, countable, and well-defined set of ”visual attributes”. The introduction of the term ”rule” facilitates the description of how these ”visual attributes” progress in a particular manner. It’s worth noting that not all visual attributes are constrained by this ”rule”; some may be assigned randomly, posing additional challenges for deep model reasoning.
Constructing a comprehensive RAVEN problem entails initially randomly selecting a rule sample from a predefined rule set. Subsequently, corresponding visual attribute values are devised based on this chosen ”rule”. For attributes free from rules, their values are assigned randomly. Once these designs are in place, images are rendered according to the derived ”attribute” information.
The RAVEN database not only encompasses problems with singular rules but also those with dual rules, elevating the level of difficulty and challenge. Specifically, the RAVEN database includes sub-databases with single rules such as center single, distribute four (G2$\times$2), distribute nine (G3×3), among others. In problems with a singular rule, the progressive variation of entity attributes within the images is governed by one set of rules. Conversely, in problems involving dual rules, like in center single out center single (O-IC), up center single down center single (U-D), etc., the attribute progression must satisfy two independent rule sets.
The design approach of PGM[16] and RAVEN problems demonstrates a striking resemblance, with both frameworks employing a problem format consisting of a stem composed of eight images and an answer pool formed by another eight images. It’s noteworthy that within the context of PGM problems, the concept of ”rule” assumes a more profound significance. This concept not only delineates the progressive pattern of ”visual attributes” along the row direction within the matrix but also imposes constraints on the progressive trend along the column direction. A representative instance of a PGM case is illustrated in the accompanying figure 2.
Figure 2: PGM case
Given this complexity, the challenges associated with RPM problems extend beyond the mere exploration of visual attributes at various levels. They also encompass the crucial task of effectively inducing and learning the progressive patterns exhibited by these ”visual attributes”. Mastering and comprehending these progressive patterns hold profound implications for successfully tackling RPM problems, thereby underscoring their importance in this cognitive assessment domain.
I-B Bongrad-logo
The Bongard problem[12, 13] stands in stark contrast to the RPM problem, falling under the umbrella of small sample learning problems. Typically, Bongard problems consist of multiple images divided into two distinct groups: primary and secondary. Images within the primary group all convey an abstract concept that is bound by a specific rule, whereas images in the secondary group deviate from these rules in varying degrees. The crux of solving Bongard problems lies in the ability of deep learning algorithms to accurately classify ungrouped images into their respective categories. Notably, Bongard-logo[13] problems, a concrete instantiation of Bongard problems within the realm of abstract reasoning and a combination of a lot of Bongard-logo cases, pose a formidable challenge in terms of inferential complexity. Each Bongard-logo case comprises 14 images, including six images from both the primary and secondary groups, along with two images awaiting classification. The grouping criterion primarily relies on the geometric shapes present within the images and their arrangement. Figure 3 illustrates a representative Bongard-logo case, where every Bongard case consists of two sets of images: Set A (primary) and Set B (auxiliary). Set A contains six images, with geometric entities in each image adhering to a particular set of rules, while Set B encompasses six images that do not conform to the rules observed in Set A. The task entails assessing whether the images in the test set align with the rules manifested in Set A. The level of difficulty varies depending on the case’s structural complexity.
Figure 3: Bongard-Logo case
Moreover, Bongard-logo problems can be further subclassified into three categories based on the concept domains they encompass: 1) Free-form problems (ff), where shapes are composed of randomly selected action strokes, with each image potentially housing one or two such shapes; 2) Basic shape problems (ba), which require recognition of a combination of one or two shape categories represented within the shape patterns; and 3) High-level concept problems (hd), designed to test a model’s ability to discover and reason about abstract concepts, such as convexity, symmetry, among others. Figure 1 exemplary presents a specific Bongard-logo case alongside its underlying principles.
II Related work
II-A RPM solver
Discriminative models for image reasoning, such as CoPINet[19], LEN+teacher[18], and DCNet[20], offer diverse solutions with multi-dimensional outputs, focusing on learning differences, latent rules, and comparing disparities. NCD[22], SCL[23], SAVIR-T[24], and neural symbolism systems like PrAE, NVSA and ALANS[25, 26, 27] incorporate various techniques to improve reasoning accuracy and interpretability. By combining effective methods, RS-CNN and RS-TRAN[28] have achieved impressive results on the RPM problem. Triple-CFN implicitly extracts information on concepts and reasoning units, and ”indexes” these reasoning unit information according to the concepts, achieving noteworthy reasoning accuracy. CRAB[29], based on a Bayesian modeling approach, has established a highly customized ”conservatory” that aligns seamlessly with its proprietary methodology, embodied in the creation of a new RAVEN database. This bespoke conservatory, despite forfeiting the inherent core challenges of RAVEN such as solution diversity and uncertainty, has facilitated remarkable breakthroughs for CRAB. Meanwhile, the scientific community eagerly anticipates the profound implications of this innovative approach for future research endeavors.
II-B Bongard solver
Researchers have tackled Bongard problems via three primary methods: language feature models, convolutional neural networks, and generated datasets. Depweg et al. used a formal language and symbolic vocabulary with Bayesian inference, but this struggled with complex problems and had scalability issues[12]. Kharagorgiev and Yun pre-trained on image datasets for feature extraction[30]. Nie et al. experimented with CNNs and meta-learning but had subpar results[13]. These methods have distinct strengths and weaknesses, underscoring the need for more research. Yuan introduced PMoC[31], a probabilistic model that measures the likelihood of auxiliary group samples fitting the main group’s distribution, a novel and effective approach. Notably, Triple-CFN[32] has successfully applied a uniform approach to both Bongard-Logo and RPM problems.
II-C Distribution distance measurement method
The formula for the Wasserstein distance[33] can be expressed as:
$W(P,Q)=\mathbf{inf}(\sum|\,|\,X-Y|\,|\,)$ (1)
where $P$ and $Q$ are two probability distributions, and $X$ and $Y$ are random variables sampled from $P$ and $Q$, respectively. Here, $|\,|\,X-Y|\,|\,$ represents the distance between $X$ and $Y$, which can be the Euclidean distance, Manhattan distance, or other suitable metrics. The infimum (inf) denotes the minimum value among all possible mapping strategies. The Wasserstein distance is a measure of dissimilarity between distributions that accounts for the degree of difference between them, particularly suitable for handling non-uniform distributions and considering the shape and structure of the distributions. In the field of machine learning, the Wasserstein distance is widely used in generative models to evaluate the discrepancy between images generated by the generator and real images.
Sinkhorn distance[34] is a metric based on optimal transport, which approximates the Wasserstein distance through the introduction of entropy regularization, leading to increased computational efficiency. This distance satisfies the triangle inequality and is symmetric, but it is not a true distance metric since it converges to the Wasserstein distance only when the regularization term tends to zero. In each iteration, the Sinkhorn algorithm computes a new transportation plan based on the current one and updates the cost matrix, gradually reducing the cost of the transportation plan and approximating the optimal transport solution.
Under certain conditions, the Sinkhorn distance satisfies Lipschitz continuity. Specifically, if we consider two probability distributions with support sets in a compact metric space satisfying certain geometric conditions, the corresponding Sinkhorn distance can be shown to be Lipschitz continuous. This means that there exists a constant L such that for all probability distributions $\mu^{\prime}$ and $\nu^{\prime}$, as well as $\mu^{\prime}$ and $\nu^{\prime}$, the following holds:
$S(\mu,\nu)\leq L\cdot d(\mu,\nu)+S(\mu^{\prime},\nu^{\prime})$ (2)
where $S$ is the Sinkhorn distance based on the cost function $c$, and $d$ is an appropriate distance between these distributions (such as the Wasserstein distance).
The Lipschitz constant $L$ typically depends on the choice of regularization parameters and cost functions. In practice, the Lipschitz constant may be difficult to compute directly but can be estimated through numerical experiments. It is worth noting that the Lipschitz constant often relates to the choice of cost function $c$. For instance, if $c$ is a metric in the metric space (such as Euclidean distance), and the geometric properties of the space (like diameter or curvature) meet certain criteria, determining bounds for the Lipschitz constant becomes easier. However, for general cost functions and regularization parameters, determining the Lipschitz constant can be a highly complex task.
When the cost function for Sinkhorn distance is Euclidean distance, its Lipschitz constant depends on parameters of the Sinkhorn algorithm used and characteristics of the dataset, such as size, shape, and distribution. Additionally, Sinkhorn algorithm parameters, like the number of iterations and the weight of the regularization term, influence the continuity properties of the resulting Sinkhorn distance.
Therefore, to determine the Lipschitz constant for Sinkhorn distance with Euclidean distance as the cost function, specific dataset experiments and analyses are necessary to evaluate the continuity properties of the algorithm under different parameter settings. In applications involving Sinkhorn distance, particularly in machine learning and optimization problems, the focus is often on numerical stability and computational efficiency rather than strictly proving Lipschitz continuity or determining the Lipschitz constant. If strict theoretical guarantees are indeed required, detailed analyses of specific problem settings may be necessary.
II-D Spectral normalization
Spectral normalization[35] is a stabilization technique for deep neural networks, particularly beneficial for generative adversarial networks (GANs). It controls the Lipschitz constant by normalizing weight matrices, ensuring outputs change proportionally to inputs. This regularizes the network, preventing overfitting and instability. In GANs, it stabilizes adversarial training, improving convergence and sample quality. Spectral normalization is computationally efficient and offers a direct way to control network complexity, making it a valuable tool for deep learning practitioners.
II-E generative model
Wasserstein GAN[36] is a deep learning model that uses the Wasserstein distance as its loss function in Generative Adversarial Networks (GANs) to improve training stability and sample quality. By optimizing this distance, it effectively reduces mode collapse and enhances the realism and diversity of generated data.
DDPM (Denoising Diffusion Probabilistic Model)[9] is a generative model that models the transition from a real image to random Gaussian noise and vice versa to create new images. It involves a forward (diffusion) process adding noise and a reverse (denoising) process restoring the original image. The aim is to generate images (x’) that are highly similar to real ones (x), emphasizing realism and authenticity.
III Methodology
Both Bongard-logo and RPM problems[14] belong to the category of abstract reasoning problems, which require solvers to possess the ability to extract specific concepts from abstract entities. These specific concepts, compared to lower-level concepts such as “pixel arrangement and configuration styles”, possess a higher level of abstraction and incorporate prior knowledge from human perspectives to varying degrees, encompassing aspects such as shape, size, color, positional relationships, graph concavity and convexity, and graph closure. Bongard-logo and RPM problems employ different questioning approaches aimed at evaluating the solver’s proficiency in extracting and learning these abstract concepts.
This paper presents Lico-Net, a novel baseline model specifically designed for the RPM problem, which achieves remarkable performance in addressing this cognitive challenge. Subsequently, we introduce a fundamental argument that representing abstract and complex concepts as distributions, rather than vectors, is a more suitable approach. Based on this insight, we develop the D3C method. This technique not only enhances Lico-Net’s performance and interpretability on RPM tasks but also elevates the reasoning accuracy of Triple-CFN, a notable deep model that has excelled in Bongard-Logo problems. Recognizing the sources of reasoning difficulties in Bongard-Logo problem, we devise the D2C method. This method is broadly applicable to both RPM and Bongard-Logo challenges, demonstrating improved performance over several baseline models in these domains. Finally, we introduce the D4C method, which targets the essential aspects of abstract reasoning. This approach contributes significantly to advancing our understanding of RPM and Bongard-Logo problems, paving the way for future research in abstract reasoning.
IV Discriminator of Case (DC)
In this section, to enhance our understanding and tackle the RPM (Raven’s Progressive Matrices) problem more effectively, we introduce a new baseline architecture, named Lico-Net. This innovative network is designed to address the unique challenges posed by RPM, leveraging advanced techniques and methodologies to achieve superior performance. Through a rigorous evaluation, we demonstrate the effectiveness of Lico-Net in solving RPM problems, paving the way for future research in this domain. Furthermore, we delve deeper into the previously established baseline network for the Bongard-Logo problem, critically examining its performance and limitations.
IV-A A novel baseline, Lico-Net, for RPM problem
The RPM problem requires participants to select answers from an option pool in order to complete a $3\times 3$ progressive matrix of images. If the completed matrix adheres to a specific image progressive pattern that can be inferred from the incomplete matrix (problem statement), it is considered a correct answer selection[39]. Therefore, a successful RPM discriminator is expected to be able to identify the rationality of the progression pattern within a complete $3\times 3$ progressive matrix.
We require an outstanding baseline, leading us to propose the Lico-Net (link-conception net) as a novel network for RPM solvers. Lico-Net is designed to assess the rationality of the image progression patterns within a complete $3\times 3$ RPM matrix, and to represent this rationality through a score. Accordingly, we can sequentially employ all the options provided in an RPM case to complete the problem matrix, leverage Lico-Net to evaluate the progressive rationality of the matrix, and assign scores to the options based on this rationality. This approach allows us to reason and solve RPM problems effectively. The design details of Lico-Net are described as follows.
IV-A1 Perceptron module of Lico-Net
Aligned with the current consensus among RPM solvers, which emphasizes the need for multi-scale or multi-viewpoint feature extraction from RPM images[28, 23, 24], Lico-Net incorporates the Vision Transformer (ViT) as a perceptron for feature extraction, while preserving the entire output vector, attention results, as multi-view point features. In this paper, we denote the number of viewpoints as $L$, a practice similar to that employed in Savirt[24], RS-Tran[28] and Triple-CFN. Lico-Net subsequently processes each viewpoint of extracted feature equally. The perceptron of Lico-Net is illustrated in figure 4.
Figure 4: Perceptron of Lico-Net
IV-A2 Reasoning module of Lico-Net
The Reasoning module of Lico-Net processes each view point feature generated by the perception module equally and in parallel, resulting in $L$ sets of inference outcomes. Subsequently, the mean value of these inference results is utilized as the global inference outcome for Lico-Net. The detailed process is described as follows.
For the isometric viewpoint of an completed $3\times 3$ progressive matrix, we employ an MLP (Multi-layers perceptron) with a bottleneck structure to extract information from all minimal reasoning units (three images within a row for RAVEN and three images within a row or column for PGM) and obtain unit vectors. Specifically, we concatenate the representations of all images within a minimal reasoning unit in their inherent order and input them into the MLP. Compared to other structures, the MLP more directly preserves the order of the input representations and its structure is sufficiently simple. The extraction process of unit vector is illustrated in figure 5.
Figure 5: The extraction process of unit vector
Due to the design of RPM problems, typically multiple minimal reasoning units must be combined to arrive at the image progression pattern of a RPM matrix. For instance, a RAVEN case requires at least two row unites to discern the case’s pattern, while PGM necessitates at least two row unites and two column unites[39]. It’s worth noting that determining the RPM matrix’s progression pattern doesn’t necessitate all minimal reasoning units; therefore, numerous combination methods of minimal reasoning units exist that are sufficient to identify the pattern. We refer to the combination requiring the least unit vectors as the minimal concept group.
Given that RPM problems might be constrained by multiple independent and decouple patterns, we embed $N$ optimizable vectors within each minimal concept group (in this paper, we embed two optimizable vectors, with $N$ being a hyperparameter that can be freely set) and input them into a standard Transformer-Encoder as tokens to compute $N$ progression pattern vectors for the progressive matrix in the case. Assuming there are $M$ minimal concept groups, we have thus obtained $N\times M$ progression pattern vectors under one viewpoint. We denote the progressive pattern vector from a single perspective as $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$. These vectors have additional purposes to fulfill. This calculation process can be expressed in figure 6.
Figure 6: The extraction process of pattern vector
This paper independently calculates the scores of each option in the problem under $N$ distinct pattern vectors.
The $M$ minimal concept groups can be divided into two categories: those involving options and those not involving options. In this paper, we select the pattern vectors calculated from the concept groups not involving options as queries, and the pattern vectors involving options as keys and values. We then compute the multi-head cross-attention results between them and use a new MLP to score the attention outputs. Since we intentionally introduced $N$ patterns, we obtain $N$ scores, which are summed to yield the final score for one viewpoint. This process is illustrated in figure 7.
Figure 7: The calculation process of final score for one viewpoint
Finally, we calculate the overall score of one option by averaging the scores from all viewpoints. Aligned with the answer logic, we employ the Cross-Entropy loss function to constrain these scores for optimizing Lico-Net. It’s worth noting that we treat each dimension of the unit vector as sample results from a distribution and use covariance matrices as additional loss function term to control the correlation between these distributions, aiming to encourage the unit vectors to contain more information. The loss function $\ell_{cov}$ attached to a batch of the minimal reasoning unit vectors $\{U_{ij}|\,i\in[1,b],j\in[1,a]\}$ can be expressed as:
$M_{cov}$ ($(\{u_{ij}|\,i\in[1,n],j\in[1,m]\})$
$=\frac{1}{n-1}\sum_{i=1}^{n}\sum_{j=1}^{m}({u}_{ij}-\bar{{u}})({u}_{ij}-\bar{{u}})^{\top},\,$
$\text{where}\,\,\bar{{u}}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{m}{u_{i}}$ (3)
$\ell_{cov}$ ($(\{U_{ij}|\,i\in[1,b],j\in[1,a]\})$
$=\frac{1}{d}\sum M_{cov}(\{U_{ij}|\,i\in[1,b],j\in[1,a]\})^{2}\cdot(1-I)$ (4)
Where $b$ represents the batchsize for training the Lico-Net, and $a$ denotes the count of minimal reasoning units within a single RPM case, which amounts to three in RAVEN and six in PGM. $I$ represents the identify matrix. And $d$ represents the dimension of unit vectors $U_{ij}$. The introduction of the covariance matrix as an additional loss function is an effective technique inherited from Triple-CFN, embodying its essence. The entire feedforward process can be depicted in figure 8.
Figure 8: Feedforward process of Lico-Net
In summary, after completing the RPM problem matrix with any given option, Lico-Net employs its perception module to encode each image within the problem matrix from multiple viewpoints, resulting in $L$ sets of $3\times 3$ image representations. The subsequent reasoning module of Lico-Net processes these $L$ sets of representations in parallel and equally. For any given set of representations, Lico-Net combines them to form all possible $M$ minimal concept groups. Subsequently, Lico-Net extracts $N$ progressive patterns from each minimal concept group, totaling $N\times M$ progressive patterns. Finally, Lico-Net utilizes an cross-attention mechanism to measure the consistency of each progressive pattern expressed by the $M$ minimal concept groups and reflects this consistency through scores, resulting in $N$ consistency scores. The final score for the option is determined by averaging these $L\times N$ consistency scores.
IV-B The baseline for Bongard-Logo problem
The cases designers of Bongard-logo drew from high-dimensional human concepts and knowledge, carefully categorizing the Bongard-logo database into three distinct problem types: FF, BA, and HD. This classification allows us to abstractly represent the distribution of the primary group in each Bongard-logo case as $p_{i}(x|\,y)$, while the distribution of the auxiliary group is denoted as $q_{i}(x|\,y)$. Here, $y$ denotes the reasoning type of the case, which can take values from the set FF, BA, HD, and $i$ represents the case index, ranging from $1$ to $n$, where $n$ is the total number of cases. Additionally, we denote the Bongard-Logo images as $x_{ij}$. Specifically, $\{x_{ij}|\,j\in[1,6]\}$ represents images in the $i$-th primary group, while $\{x_{ij}|\,j\in[8,13]\}$ represents images in the $i$-th auxiliary group. Additionally, $x_{i7}$ represents the test image to be potentially assigned to the $i$-th primary group, and $x_{i14}$ represents the test image to be potentially assigned to the $i$-th auxiliary group.
We copy a baseline from Triple-CFN, which refers to the baseline as $f_{\theta}(z|\,x)$ that maps samples $x_{ij}$ to latent variables $z_{ij}$. The baseline employs the InfoNCE loss function, which results in the $f_{\theta}(z|\,x)$ being able to solve Bongard-logo to a certain degree. The InfoNCE loss function[41] can be mathematically expressed as:
${\ell_{\text{InfoNCE}}}({z_{pri}},{\tilde{z}_{pri}},\{{z_{au{x_{m}}}}\}_{m=1}^{M})$
$=-\log\frac{{{e^{({z_{pri}}\cdot{\tilde{z}_{pri}})/t}}}}{{{e^{({z_{pri}}\cdot{\tilde{z}_{pri}})/t}}+\sum\nolimits_{m=1}^{M}{{e^{({z_{pri}}\cdot{z_{au{x_{m}}}})/t}}}}}$ (5)
Where $z_{pri}$ and $\tilde{z}_{pri}$ represent distinct samples belonging to the primary group, and $z_{au{x_{m}}}$ encompasses all $m$ samples within the auxiliary group. $m$ denotes the number of negative samples within a minimum learning unit, while $t$ represents the temperature coefficient, and $t$ is set $10^{-3}$ in this paper.
Through the strategic employment of the InfoNCE function, we can circumvent the need to explicitly estimate the distributions $p_{i}^{\prime}(z|\,y)$ and $q_{i}^{\prime}(z|\,y)$. Concurrently, this approach encourages enhanced similarity among the representations of the primary group encoded by the network $f_{\theta}(z|\,x)$ and guarantees a greater degree of mutual exclusion between the representations of the primary and auxiliary groups. This process aligns with the underlying logic of Bongard-logo. InfoNCE can be construed as a loss function devised for the clustering of vector groups.
This paper reproduces the experiments for the baseline in Triple-CFN conducted based model ResNet50, which are individually implemented on the Bongard-logo database based on ResNet18. It focuses on the three concepts: ‘FF’, ‘BA’, and ‘HD’ for training, while concentrate on four test concept: ‘FF’, ‘BA’, ‘NV’ and ‘CM’. Subsequently, a combined database was created by merging these three individual databases, and experiments were performed on this combined database. The experimental results are presented in Table I. The feedforward process is shown in figure 9.
The findings indicate that when presented with a combination of the three databases, the convolutional deep model experiences confusion among the three distinct human concepts. The varying degrees of reduction in the model’s reasoning accuracy on the FF and BA problems, as well as the increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion. Furthermore, this confusion is not attributable to insufficient network capacity, as even the ResNet50 employed in this study, which is more capacious than the ResNet18 utilized in Triple-CFN’ s baseline, falls victim to the same predicament.
Figure 9: Feedforward process baseline
TABLE I: Reasoning Accuracies of DC based ResNet based ResNet50 on Bongard-logo.
Test Accuracy(%)
Model and Data Set FF BA CM NV
ResNet18 on Bongard-logo 88.1 97.9 76.0 75.8
ResNet18 on Separated Bongard-logo 97.9 99.0 75.0 72.8
ResNet50 on Bongard-logo 88.9 98.1 76.1 76.0
ResNet50 on Separated Bongard-logo 98.3 99.2 75.3 72.2
This paper posits that, in the Bongard-Logo problem, the determination of image attributes relies heavily on the primary and auxiliary group samples within each case, which poses a fundamental challenge. It is precisely this multifaceted attribution of image properties that confuses deep networks when dealing with the three high-dimensional concepts inherent in the Bongard-Logo problem. Conversely, the RPM problem exhibits consistency in its recognition of image attributes. For instance, a white medium-sized triangle retains its attribute identification as such, regardless of its appearance in any RPM case option pool.
V Distribution Distance-based Discriminator of Case (D3C)
Using distributions to describe human concepts in abstract reasoning problems for the deep networks is preferable to using vectors. Distributions offer a more comprehensive and nuanced approach to capturing the diversity and uncertainty inherent in human concepts. Human concepts are often ambiguous, polysemous, and subject to varying degrees and fluctuations, which can be challenging to represent using static vectors. In contrast, distributions provide a dynamic and probabilistic framework that can accommodate this inherent variability.
By leveraging distributions to represent human concepts, we can furnish deep networks with richer and more flexible information. This enhanced representational capacity enables the networks to better understand and address the complexity and uncertainty encountered in abstract reasoning problems. For instance, probability distributions can be utilized to depict the range of possibilities and variations associated with a concept, while Gaussian distributions can model the continuity, smoothness and even discreteness of a concept’s underlying structure.
V-A D3C for Bongard-Logo
The ideal approach to address the Bongard-Logo problem involves training an encoder capable of encoding images from a Bongard-Logo case into representations, denoted as $z$. In this context, the representations of positive instances (samples in primary group) belong to the distribution $p(z)$, while the representations of negative instances (samples in primary group) belong to $q(z)$. Importantly, there should be minimal overlap between $p(z)$ and $q(z)$ in terms of cross-measure. Therefore, this paper proposes the D3C method to enhance the performance of encoders in the form of $f(z|x)$ on Bongard-Logo problems. Notably, this brings improvements for models such as Triple-CFN and ResNet50.
Specifically, this paper hypothesizes that the representation of the primary group samples in a Bongard-Logo case follows a Gaussian distribution, while the representation of the auxiliary group samples follows a mixture of Gaussian distributions. We utilize an arbitrary backbone network $f(z|\,x)$ to approximate the distribution of image representations within Bongard-Logo cases. To achieve this, we modify the network to output parameters $\mu$ and $\sigma$ instead of a single vector, effectively converting $f(z|\,x)$ into $f(\mu,\sigma|\,x)$. The encoded vector is then divided into two equal-length segments, which we interpret as the mean and the logarithm of the variance, respectively. This approach normalizes the output vectors into Gaussian distributions and facilitates the distribution of image representations in Bongard-Logo cases. Our method bears resemblance to the Variational Autoencoder (VAE) framework. We argue that using distributions, rather than vectors, to describe image representations offers greater flexibility and accuracy in capturing the underlying complexity and uncertainty of abstract reasoning problems.
With the estimated distributions of representations for each image in the Bongard-Logo cases, we compute the Sinkhorn distance between these distributions using reparameterization technique. Following the logical structure of each Bongard-Logo case, we impose constraints on these distances for the backbone network $f(\mu,\sigma|\,x)$, aiming to train an effective discriminator. Specifically, we optimize the network by utilizing a binary cross-entropy loss function to enforce constraints on the exponential function of the negative Sinkhorn distance with base $e$ ($e^{-D}\in[0,1]$, where $D\in[0,+\infty]$ is the values of Sinkhorn distance), aligning with the case’s underlying logic mentioned in DC. This process can be expressed as figure 10. The D3C Method, applied to a batch of Bongard-Logo case distributions encoded by the backbone network $f(\mu,\sigma|\,x)$, can be expressed as follows:
${\ell_{\text{D3C}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
$=\sum_{i=1}^{b}\sum_{j=1}^{7}\sum_{\tilde{j}=j+1}^{7}{\text{Sinkhorn}({z_{ijs}},\,{{z}_{i\tilde{j}s}})}$
$+\sum_{i=1}^{b}\sum_{{j}=1,}^{7}\sum_{\tilde{j}=7}^{14}\log({1-{e^{-\text{Sinkhorn}({z_{ijs}},\,{{z}_{i\tilde{j}s}})}}})$ (6)
Where $b$ is the batchsize of training. $S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
Figure 10: Feedforward process of D3C
The design of D3C allows for its application to any image encoder that can be abstracted as $f(z|x)$. This requirement is met by a wide range of encoders, including those from the ResNet family, ViT, and specialized ones like Triple-CFN, which is particularly suitable for Bongard-Logo problems.
V-B D3C on Lico-Net for RPM
Notably, in contrast to the Bongard-Logo database, both RAVEN and PGM possess well-defined and comprehensive Meta data in the form of (auxiliary) rule labels. These rule descriptions facilitate the construction of a rational concept space, with boundary concepts within the rule labels. This paper introduces the rule labels from the RPM database into the Lico-Net, aiming to enhance the reasoning accuracy and human-perspective rule interpretability of Lico-Net. This process will be implemented by the D3C method.
Within the frameworks of RAVEN and PGM, there is a many-to-one mapping between image progression patterns and their corresponding rules. For instance, both single-step and double-step increments, as well as single-step and double-step decays in image attributes, are all categorized under the ‘progressive’ rule. Therefore, this paper proposes the concept of rules from the RPM database and represent them as distributions while also vectorizing the image progression patterns. We can leverage this approach to naturally constrain the image progression patterns within their corresponding rule distributions, thereby optimizing the deep model’s learning of rules and clarifying the boundaries among various concepts within the data.
In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual descriptions of Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data. Specifically, the initial portion of the feature vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the rule descriptions and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$. Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors termed as progression pattern vectors. Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns should exhibit a high likelihood of adhering to the distribution outlined by their corresponding textual rules, and conversely, demonstrate a low probability of deviation.
A specific RPM case is encoded into $N\times M$ progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ from $L$ viewpoints by Lico-Net as aforementioned. To align Lico-Net with the D3C framework, the number of encoding patterns $N$ in Lico-Net is set according to the instance of the RPM problem, depending on the number of decoupled concepts present in the Meta data of that instance. For example, in PGM, there are two such concepts: shape pattern and line pattern, and $N$ is set to 2. The special setting of $N$ in Lico-Net can be viewed as $N$ slots, which can be aligned with the $N$ decoupled concepts described in the Meta data, respectively. By synthesizing all perspectives—specifically, by calculating the mean of the progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ across all $L$ viewpoints—Lico-Net encodes $N\times M$ vectors, denoted as $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$, representing the full-viewpoint progression patterns. This process can be expressed as:
$\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$
$=\frac{1}{L}\sum_{l=1}^{L}\{P_{nm}|\,n\in[1,N],m\in[1,M]\}_{l}$ (7)
The $l$ represents the index of viewpoints. The $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$ can be treated as $N$ sets of $M$ sampling results. This enables the optimization of distribution distances based on the Sinkhorn distance, thereby enhancing the overall performance through D3C framework. Apparently, this paper utilizes the Sinkhorn distance to constrain the distance between $M$ progression pattern vectors and their corresponding distributions’ sampling results. The $N$ decoupling concept presented in Meta data is employed to concurrently impose constraints on the $N$ sets of $M$ progression pattern vectors, and the process of aligning $\{\overline{P}_{nm}\}$ with the distributional concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ can be illustrated in figure 11.
Figure 11: Feedforward process of D3C
The calculate porcess can be exprecessed as:
${\ell_{\text{D3C}}}(\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\},\{r_{ks}|\,k\in[1,K],s\in[1,S]\})$
$=\sum_{n=1,\tilde{k}|\,n}^{N}{\text{Sinkhorn}({\overline{P}_{nm}},\,{r_{\tilde{k}s}})}$
$+\sum_{k=1,k\neq\tilde{k}}^{K}\sum_{n=1,\tilde{k}|\,n}^{N}\log({1-{e^{-\text{Sinkhorn}({\overline{P}_{nm}},\,{r_{{k}s}})}}})$ (8)
The $\overline{P}_{nm}$ represents the full-viewpoint progressive pattern vectors encoded by Lico-Net, and $\{r_{ks}\}$ represents the $S$ sampling results from the distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of the rule description. It is evident that the specific value of $\tilde{k}$ is contingent upon the description of Meta data corresponding to the particular case of the RPM problem. symbol “$\tilde{k}|\,n$” represents that the specific value of $\tilde{k}$ is depended on $n$. This dependency realizes the fact that each slots in Lico-Net are aligned with their corresponding concepts. The $\text{Sinkhorn}(\cdot,\cdot)$ represents the Sinkhorn distance operator. Other logic of symbol in function (V-B) has mentioned. It becomes evident that we are employing the Binary Cross Entropy (BCE) loss function to impose constraints on the exponential negative Sinkhorn distance result. The feedforward process is illustrated in the figure 12.
Figure 12: Feedforward process of Lico-Net aligned with D3C
V-C A low-cost loss function for the D3C, the Hermes for the Zeus
Sampling $S$ times (where $S$ is arbitrarily set, and $S$ is set to $20$ for Sinkhorn distance) from the constructed distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of concepts in Meta data and yielding $\{r_{ks}|\,k\in[1,K],s\in[1,S]\}$, we calculate the Sinkhorn distance between the progression pattern vectors and these sampling results $\{r_{ks}\}$. This Sinkhorn distance approximates the Wasserstein distance, which is currently a linear programming algorithm also known as the Earth Mover’s Distance (EMD) algorithm. Typically, this approach requires a substantial number of samples and incurs a considerable computational cost. Even if we treat the $M$ progression pattern vectors as $M$ sampling results, it is still difficult to satisfy the prerequisite of extensive sampling in the context of the RPM problem, where $M$ is set to 9 in PGM and 3 in RAVEN. To reduce computational cost and sampling frequency, this paper employs cosine similarity between vectors as an alternative distance metric norm to the Sinkhorn distance. Obviously, the cosine similarity satisfacte with the $Lipschitz$ continuity condition. For the inner product, treating it as a function with one fixed vector, we observe that it exhibits $Lipschitz$ continuity. Specifically, given a fixed vector $v$, the function $f(u)=u\cdot v$ is $Lipschitz$ continuous with respect to $u$, with the $Lipschitz$ constant $L$ being equivalent to the norm of $v$. This property is attributed to the inner product’s adherence to the Cauchy-Schwarz inequality: $|u\cdot v|\leq||u||\cdot||v||$. Leveraging this inequality, we are able to deduce the underlying inequality:
$f(u_{1})-f(u_{2})=|u_{1}\cdot v-u_{2}\cdot v|=|(u_{1}-u_{2})\cdot v|\leq||u_{1}-u_{2}||\cdot||v||$ (9)
In our study, we opted for cosine similarity, considering normalized vectors $u$ and $v$, where $||v||$ and $||u||$ are equal to 1, and Lipschitz constant $L=||v||=1$. It is reasonable to choose the mean of cosine similarity between sampling vectors as the distance metric function instead of Sinkhorn distance. The calculate process of new method applied on Bongard-Logo can be expressed as:
${\ell_{\text{D3C-cos}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
$=-\sum_{i=1}^{b}\sum_{j=1}^{7}\sum_{\tilde{j}=j+1}^{7}\sum_{s=1}^{S}\log\frac{{{e^{({z_{ijs}}\cdot{z_{i\tilde{j}s}})/t}}}}{{{e^{({z_{ijs}}\cdot{z_{i\tilde{j}s}})/t}}+\sum\nolimits_{k=8}^{14}{{e^{({z_{ijs}}\cdot{z_{i{k}s}})/t}}}}}$ (10)
The calculate process of new method aligned with Lico-Net can be expressed as:
${\ell_{\text{D3C-cos}}}(\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\},\{r_{ks}|\,k\in[1,K],s\in[1,S]\})$
$=-\sum_{n=1,\tilde{k}|\,n}^{N}\sum_{m=1}^{M}\sum_{s=1}^{S}\log\frac{{{e^{({\overline{P}_{nm}}\cdot{r_{\tilde{k}s}})/t}}}}{{{e^{({\overline{P}_{nm}}\cdot{r_{\tilde{k}s}})/t}}+\sum\nolimits_{k=1,k\neq\tilde{k}}^{K}{{e^{({\overline{P}_{nm}}\cdot{r_{{k}s}})/t}}}}}$ (11)
The $t$ is set to $10^{-6}$ and $S$ is set to $5$ in function (V-C) and (V-C). The logic of symbols in function (V-C) and (V-C) has mentioned. It is evident that, compared to function (V-A) and (V-B), the demand for sampling frequency $S$ in D3C-cos has been significantly reduced.
From a certain perspective, estimating the distribution distance based on cosine similarity between sampling results can be regarded as a variant of the Wasserstein distance based on cosine similarity that relinquishes the lower bound estimation and possesses a relatively low, fixed Lipschitz constant. Therefore, this method, we called “D3C-cos”, cannot precisely quantify the distance between distributions. It serves as a low-cost loss function for constrained inter-distribution distance, serving as a surrogate for the Sinkhorn distance, the Hermes for the Zeus.
The success of such simplification can be attributed, in part, to our modeling of the distribution of pattern descriptions as a Gaussian distribution.
VI Dynamic Delineation of the Concept (D2C)
In this section, the paper initially endeavors to introduce novel constraints for the Bongard-Logo discriminators, aiming to enhance their performance. Motivated by these constraints, we incorporate innovative designs into the proposed Lico-Net architecture.
VI-A D2C on Bongard-Logo problem
In the Bongard-Logo problem, the determination of image attributes is significantly influenced by specific contextual concepts associated with each individual case. Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem.
To achieve this objective, we have redefined the concepts within the Bongard-Logo database, representing them as conditional probability distributions $p_{i}(x|\,Y_{i})$ and $q_{i}(x|\,Y_{i})$. Here, $i$ ranges from $1$ to $n$, denoting each index of the Bongard-Logo cases. Additionally, we introduce the notation $\{Y^{\prime}_{i}|\,i\in[1,n]\}$ to represent the set of reorganized concepts. It is evident that the number of reorganized concepts corresponds to the number of Bongard-Logo cases. Through this concept reorganization, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases. In other words, each case is uniquely categorized based on its distinct set of concepts. By establishing a one-to-one correspondence between image styles and these redefined concepts, we aim to enable the network to seamlessly infer concepts from the pixel arrangements of images. This, in turn, aims to mitigate the impact of multiple attribute identifications on the solution of the Bongard-Logo problem.
To foster concept antagonism between different problem domains, we have revised the data sampling techniques employed in mini-batch training methodologies. Furthermore, an additional loss function term has been devised to facilitate the reassignment of concepts within the network architecture. This approach, termed “D2C” (Dynamic Delineation of Concepts), endeavors to cultivate a state of concept contrast amidst diverse problem sets.
VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
By augmenting the traditional sampling methods used for mini-batch training with an additional ancestral sampling procedure, this study identifies mutually exclusive pairs of Bongard-logo problems for training purposes. This constitutes a hot-swappable approach for training the Bongard-Logo solver, which is designated as phase 1 of the D2C method in this paper.
Specifically, we denote a batch of Bongard-logo cases as $\{x_{ij}|\,i\in[1,b],j\in[1,14]\}$, where $i$ denotes the index of the Bongard-logo case within a batch, the batch size is $b$, and $j$ represents the sample index within a single Bongard-logo case, with $j\in[1,7]$ being the primary samples and $j\in[8,14]$ being the auxiliary samples. From this batch of cases $\{x_{ij}\}$, we deliberately choose two batches of cases $\{x_{tj}|\,j\in[1,14]\}$ and $\{x_{\tilde{t}j}|\,j\in[1,14]\}$, ensuring that $t$, $\tilde{t}$, and $i$ are all distinct. This process can be viewed as ancestral sampling. The extracted primary samples $\{x_{tj}|\,j\in[1,7]\}$ and $\{x_{\tilde{t}j}|\,j\in[1,7]\}$ are then incorporated into $\{x_{ij}\}$ as auxiliary samples, yielding an expanded set $\{x_{i\tilde{j}}\}$ where $\tilde{j}\in[1,28]$. Subsequently, The new sample groups $\{x_{i\tilde{j}}\}$, which oppose the concepts from different cases compared to their original form, can be fed into any discriminator used to solve Bongard-Logo problems, such as Triple-CFN[32] and PMoC[31]. The feedforward process, as illustrated in the accompanying diagram, involves a strategic misalignment of Bongard-logo problem batches to facilitate ancestral sampling during implementation. The alignment of the D2C phase 1 process with the baseline in DC is illustrated in figure 13.
Figure 13: Feedforward process of D2C phase 1
It is worth emphasizing that the improvement in model performance observed through the aforementioned approach is not solely attributed to an increase in the number of negative examples within each contrastive group. This paper posits that the incorporation of negative examples, which assist in delineating concept boundaries, holds the key to enhancing a model’s proficiency in abstract reasoning tasks. Merely augmenting the dataset with insignificant negative examples does not lead to substantial gains in model performance. In an attempt to enrich the dataset, this study generated a substantial amount of Bongard-logo virtual data using DDPM and integrated these as auxiliary data points within the original problems. However, subsequent analysis revealed that this methodology did not tangibly enhance model accuracy.
In summary, the phase 1 of the D2C approach embodies a straightforward yet effective concept: it introduces positive examples from other cases as negative instances in the current case. This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts. As a result, the first phase of D2C is widely applicable to various Bongard-Logo discriminators, including Triple-CFN and PMoC.
VI-A2 D2C phase 2: The concept delineation method based on a new loss function term
This paper posits that, in addition to altering the manner of sampling training data, establishing an extra loss function to constrain the similarity between the representations of the primary group samples across different cases within a batch of Bongard-Logo training data can implicitly achieve the fundamental objective of this paper.
Specifically, in this paper, we exclusively retain the representation of primary group within the set $\{z_{ij}|\,i\in[1,b],j\in[1,14]\}$, which specifically correspond to $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$. Subsequently, we randomly partition the primary representation of each case within $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ into two distinct subsets: $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, where $\{z_{ih}|\,i\in[1,b]\}\cup\{z_{i\tilde{h}}|\,i\in[1,b]\}=\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ and $\{z_{ih}\}\cap\{{z}_{i\tilde{h}}\}=\emptyset$. By taking the mean of the vectors within $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, we derive the corresponding vectors $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$. Ultimately, we compute InfoNCE loss function for constraining the similarity among $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$ as the additional loss function item. The specific calculation process can be explained by:
${\ell_{\text{D2C}}}(\{z_{ij}|\,\,i\in[1,b],j\in[1,7]\})$
$=-\sum^{b}_{i=1}\log\frac{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}}}{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}+\sum\nolimits_{\tilde{i}=1,\tilde{i}\neq i}^{b}{{e^{({v_{i}}\cdot{v_{\tilde{i}}})/t}}}}}$
$\text{where}\,\,v_{i}=\sum\{z_{ih}\},\,\tilde{v}_{i}=\sum\{z_{i\tilde{h}}\}\,\,$
$\text{and}\,\,\{z_{ih}\}\cup\{z_{i\tilde{h}}\}=\{z_{ij}|\,j\in[1,7]\}$ (12)
This loss term is to decrease the similarity of primary group representations across different cases. The schematic representation of the feedforward procedure is provided in the figure 14.
Figure 14: Feedforward process of D2C phase 2
In essence, the phase 2 of D2C attempts to achieve a similar goal as its first phase by formulating a new loss function, approaching the task from a different angle. Nevertheless, this strategy renders the phase 2 of D2C primarily compatible with backbone that can be abstracted as $f(z|x)$. This adaptability enables a majority of image encoders, such as ResNet, ViT, and Triple-CFN[32], to be compatible with the complete D2C methodology. However, the application of PMoC[31], which relies on probability-based measures to assess image identity, is restricted to the first phase of D2C.
VI-B D2C on RPM probelm
Furthermore, the spirit of D2C method can be integrated with Lico-Net by introducing images from other case statements as negative examples within the cases. However, the application of D2C in RPM discriminators is limited to its phase 1, as it becomes challenging to forcibly apply the phase 2 to these discriminators. This process is illustrated in figure 15.
Figure 15: Feedforward process of Lico-Net aligned with D2C
In summary, D2C is a straightforward approach. In the context of Bongard-Logo, where the determination of image attributes heavily relies on the concepts outlined by the primary and auxiliary groups within each case, D2C addresses this challenge by decoupling the concepts across different cases. Nevertheless, regarding the RPM problem, which emphasizes consistency in image attribute identification, this study maintains a pessimistic outlook on the efficacy of D2C in tackling RPM. It is worth emphasizing that the D2C and D3C methodologies are not mutually exclusive. Specifically, D2C endeavors to redefine concept boundaries for reasoning problems. According to D3C, representing concepts in the form of distributions is a more plausible approach compared to utilizing vectors.
VII Defense-Driven Discrimination of Distributed Concepts (D4C)
This paper aims to delineate more appropriate concept boundaries for the RPM and Bongard-Logo problems.
VII-A Negative examples are not merely intended to confound the discriminator
This paper posits that it is worthwhile to consider the significance of positive and negative examples in abstract reasoning problems. Suppose that the correct solutions to each case in a reasoning problem follow a Gaussian distribution. Given that some reasoning problems may have multiple correct solutions, it is also possible that the distribution is a mixed Gaussian distribution. This is exemplified in datasets such as OIG and D9 from Raven database. One of the core challenges in RPM problems lies in the existence of multiple correct answers. While CRAB has eliminated this core difficulty through data cleaning in the RAVEN dataset, it has achieved noteworthy results on the cleaned version.
Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern where positive examples yield high probabilities when substituted into the case’s problem statement, while negative examples yield low probabilities. We can view this pattern in the following way: the discriminator constructs a solution distribution for the cases presented in abstract reasoning problems. Within this distribution, positive examples are assigned higher probabilities, whereas negative examples are assigned lower ones. The endeavor to train the discriminator can be conceptualized as a process of optimizing and fine-tuning the solution distribution, with the ultimate objective of aligning it more closely with the underlying distribution of correct solutions for the given cases. Using positive and negative examples to estimate a Gaussian distribution through the cross-entropy loss function can be seen as leveraging them to optimize the following function:
$-\log(p(x,\mu,\sigma))=\frac{{(x_{i}-\mu_{i})^{2}}}{{2\sigma^{2}_{i}}}-\log(|\,\sigma_{i}|\,)-\frac{1}{2}\log(2\pi)$ (13)
Where $x$ represents the positive or negative samples, $\mu$ and $\sigma$ represent the mean and standard deviation of the Gaussian distribution separately. It can be observed that when positive instances are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the mean squared loss between sample $x$ and the mean of the correct solution $\mu$. Conversely, when negative instances are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$ of the respective segment.
Therefore, this paper posits that in reasoning problems, the establishment of positive examples aids deep models in estimating the mean of the correct solution distribution, while negative examples serve to estimate the variance. The variance represents the distance between the boundaries of the distribution and its center. Correspondingly, we posit that negative examples play a crucial role in abstract reasoning tasks, serving to delimit the concepts presented in each case. This underscores the significance of establishing both positive and negative examples in abstract reasoning problems. This significance reflects a despairing fact that the upper limit of the discriminator’s reasoning accuracy in this abstract reasoning problem is determined by the quality of positive and negative examples in the question.
VII-B High-quality negative examples lead to more rigorous concept boundaries
The D2C methodology proposed in this study has established a novel problem boundary for deep networks by means of resampling negative instances and incorporating a fresh loss function. Experimental findings tentatively indicate that this boundary is better suited for scenarios characterized by the absence of conceptual supervision signals, as compared to the boundaries delineated by the original negative instances in Bongard-Logo. Nonetheless, the precise tightness of the new boundary remains indeterminate.
In essence, D2C introduces a new variance to the solution distribution of the network, the magnitude of which relative to the variance of the correct solution distribution is indeterminate. In the context of the Bongard-Logo challenge, the variance provided by D2C appears closer to that of the correct solution distribution in the absence of conceptual supervision signals, as compared to the variance described by the original negative instances.
Evidently, the efficacy of D2C is contingent upon specific dataset attributes. As anticipated, empirical evidence suggests that this approach may be insufficient for delineating precise boundaries within the framework of the RPM problem. Contrary to the Bongard-Logo paradigm, RPM problems postulate a shared set of core concepts across distinct instances, ensuring discreteness and stability. This shared nature precludes the utility of methods like D2C in addressing RPM challenges. Specifically, image properties within RPM instances are rigidly governed by definitions of “rules” or “progressive patterns,” maintaining consistent attributions across all instances.
The Bongard-Logo problem exhibits varying attributions of image properties due to differing clustering bases among instances, thereby augmenting problem complexity and difficulty. Incorporating positive instances from other cases as negative examples can, to some extent, delineate conceptual boundaries between distinct cases, adhering to a cautious approach aimed at minimizing conceptual infringement. However, this strategy proves inapplicable in addressing challenges presented by RAVEN and PGM due to substantial disparities across multiple dimensions.
In summary, the success of D2C in addressing the Bongard-Logo problem can be attributed to its indirect enhancement of the quality of negative instances within Bongard-Logo cases. Conversely, its limitations in tackling the RPM problem can also be ascribed to its insignificant improvement in the quality of negative examples within RPM scenarios. To address the limitations inherent in the D2C approach, this paper introduces the D4C (Defense-Driven Discrimination of Distributed Concepts) method. The underlying idea of D4C is to generate higher-quality negative instances and samples within abstract reasoning problems. This approach aims to delineate a more appropriate problem boundary for the discriminator as compared to the original problem setup in abstract reasoning. By leveraging novel mechanisms and strategies, D4C aims to deliver more precise and reliable definitions of conceptual boundaries and foster further advancements in the realm of abstract reasoning.
VII-C Impelmention of D4C method
The D4C methodology proposed in this paper emphasizes on devising a more suitable concept boundary for both RPM and Bongard-Logo discriminators, aiming to enhance their problem-solving capabilities. To achieve this objective, a distinctive approach centered on the generation of confusing samples is introduced. These samples are intentionally designed to challenge and improve the discriminator’s accuracy. A dedicated generator has been devised to carry out this generation process. Through iterative defense against attacks launched by this generator, the discriminator progressively refines its discrimination and classification skills, leading to overall performance enhancement in both areas. Specifically, within this study, a generator has been tailored to augment the capabilities of two distinct Bongard-Logo discriminators: PMoC and Triple-CFN. Additionally, this generator can also be directly utilized for the refinement of Lico-Net. Nevertheless, acknowledging the inherent disparities between RPM problems and Bongard-Logo challenges, additional designs have been incorporated into the generator to cater more effectively to the refinement requirements of Lico-Net.
In this section, we delve into the design of the discriminator and generator within the D4C framework as it pertains to the Bongard-Logo and RPM problems.
VII-C1 discriminator for Bongard-logo
Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC and Triple-CFN, achieving notable successes. These discriminators seamlessly integrate with the D2C and D3C methodologies proposed in this study, and pertinent experiments have been conducted to explore their combinations. These discriminators excel in accurately identifying and classifying Bongard-Logo instances, significantly contributing to the advancement of research in this domain.
Regarding discriminators capable of handling adversarial tasks, the Wasserstein GAN, a prominent adversarial generative network, poses a noteworthy requirement. The discriminator must undertake the task of measuring the distance between samples generated by the generator and real samples. This necessitates that the discriminator satisfies Lipschitz continuity to a certain extent.
In this study, the choice of PMoC, constrained by spectral normalization, as the discriminator for the D4C method on Bongard-logo is motivated by its explicit measurement of the distance between auxiliary and primary group sample distributions. Conversely, the task of Triple-CFN is relatively more intricate as it considers not only the distance between primary and auxiliary samples but also the distance among primary samples. Consequently, compared to PMoC, this study entails more elaborate designs for the generator of Triple-CFN.
VII-C2 generator for Bongard-logo
In this paper, we argue that instead of forcing the generator to produce samples that can confuse the discriminator, it is more effective to generate representations of confusing samples, referred to as confusing representations. This approach reduces the computational burden on the network, as measured by intuitive metrics such as computational complexity.
For the Bongard-Logo problem, we model the overall distribution as a mixture of Gaussian distributions. Within this framework, each individual primary group in a Bongard-Logo case is represented as a separate Gaussian distribution. Consequently, the D4C generator designed for the Bongard-Logo problem exhibits a relatively straightforward structure. The generator aims to encode a distribution based on the representations of the primary groups in a given Bongard-Logo case, ensuring that any representation sampled from this distribution is capable of effectively confusing the discriminator.
Specifically, we employ a standard Transformer encoder to encode a Gaussian distribution based on the representations of the primary samples in a Bongard-Logo case, along with an optimizable vector. This encoding approach bears similarity to D3C, wherein we extract the optimizable vector from the attention results and process it through a linear feed-forward layer to obtain the mean and logarithm of variance for the Gaussian distribution. Subsequently, we sample multiple representations from this distribution as confusing representations to challenge the discriminator, which must exert maximum effort to counter them. The generator’s training objective is to generate representations for Bongard-Logo cases that achieve higher scores in the discriminator compared to the representations of the original negative examples in the case. Conversely, the discriminator’s training objective is unambiguous: to classify these generated representations as negative examples. The aforementioned approach can be seamlessly applied to the PMoC model, providing a direct and effective means of enhancing its performance. The alignment between the generator and PMoC is illustrated in Figure 16.
Figure 16: Feedforward process of D4C aligned with PMoC on Bongard-Logo
In this figure, the “$f_{\theta}(z_{ij}|x_{ij})$” and “$g_{\omega}(p^{\prime}_{i}(z_{ij}|y)|\{z_{ij}\}_{j=1}^{6},z_{ij^{\prime}})$” represent the perceptron and logic calculation modules in PMoC, respectively[31].
For the generator in Triple-CFN, it is essential to attenuate the generated representations. Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group. After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to attenuate the confusing representation vector. This attenuation process is necessary due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation. This makes it easier for a generator based on the Transformer encoder to generate realistic confusing representations, potentially causing Triple-CFN to collapse. Therefore, we have designed the aforementioned attenuation process. The alignment between the generator and Triple-CFN is illustrated in Figure 17. In this figure, “$f_{\theta}(z_{ij}|x_{ij})$” represents the encoder in Triple-CFN.
Figure 17: Feedforward process of D4C aligned with Triple-CFN on Bongard-Logo
VII-C3 discriminator of RPM
In addressing Raven’s Progressive Matrices (RPM) problems, this study places significant confidence in the Lico-Net model, considering it as an exceptional benchmark. It is anticipated that the application of spectral normalization will enhance the performance of Lico-Net, positioning it as a highly suitable discriminator for the D4C methodology.
However, Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix. To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations. Any inconsistencies or irregularities identified within these representations will be flagged as negative instances, indicating incorrect responses. Through this rigorous process, Lico-Net strives to accurately identify and reject perplexing representations, thereby augmenting its overall precision and reliability.
VII-C4 generator of RPM
Unlike the Bongard-Logo problem, the answer to a single case in the RPM problem exhibits non-uniqueness, and thus can be represented as a mixture of Gaussian distributions. Therefore, our aspiration is to tailor a generator for the D4C method within the RPM context, capable of encoding a distribution that reflects this mixed Gaussian nature. In this paper, we employ a glove-shaped network to simulate the process of sampling from such mixed Gaussian distributions.
Specifically, we utilize a standard Transformer-Encoder to process the representations of all problem images in an RPM case, alongside $W$ optimizable vectors (where $W$ is a hyperparameter) and a stochastic promoter. Notably, when these $W+1$ vectors are arranged together, they resemble the shape of a glove. The promoter is a vector sampled from a multivariate Gaussian distribution, wherein the mean is optimizable while the variance is fixed at 1.
It is worth noting that we have devised a novel positional encoding embedding approach for the Transformer-Encoder, which serves as the backbone of our RPM generator. This encoding aims to restore the $3\times 3$ image arrangement characteristic of the RPM problem matrix. For the PGM (Progressive Graph Matching) problem, our positional encoding scheme assigns unique encodings to images located on the diagonal and upper triangular regions within the problem matrix of a given PGM case. Correspondingly, the positional encodings for images in the lower triangular region are symmetrical with respect to the diagonal, mirroring the encodings of their counterparts in the upper triangular region. However, it should be acknowledged that this paper does not demonstrate the superiority of D4C on the relatively less challenging RAVEN dataset.
Moving on to the main topic, after obtaining attention results from the Transformer-Encoder, we extract $W$ optimizable vectors from these results and process them into $W$ distributions using a linear layer, as mentioned earlier. Following the extraction of the promoter, it is processed into a probability output using a linear layer and hard-gumbel-softmax activation, yielding highlighted probabilities for the $W$ distributions. Subsequently, we sample representations, referred to as confusing representations, from the highlighted distributions each time. In the technical implementation, we employ the reparameterization trick to draw samples from each distribution, resulting in $W$ sampling outcomes. We then perform a weighted sum of these sampling results using the hard-highlighted probabilities processed by the hard-gumbel-softmax, thus completing the entire mixed-Gaussian sampling process. The methodology outlined in this paper for modeling and training the network is coined as ”D4C glove-train”. A more detailed illustration of the forward process is provided in the accompanying figure referenced as 18.
The training objective of the glove-shaped generator remains consistent: to produce confusing representations that achieve higher scores in the discriminator compared to the original negative examples provided by the dataset. Conversely, the discriminator aims to resist the interference posed by these confusing representations.
Figure 18: Feedforward process of D4C on RPM problem
Finally, in the realm of abstract reasoning problems, particularly those involving generating answer samples, this paper believes that the D4C approach can also make valuable contributions. This opens up new avenues of exploration for future work, presenting challenges that are yet to be tackled.
VIII Experiment
All experiments conducted in this paper were programmed in Python, utilizing the PyTorch[43] framework.
VIII-A Experiment on RAVEN
Initially, we conducted experiments with Lico-Net on both RAVEN[15] and I-RAVEN[37] datasets to demonstrate its inherent superiority. To ensure a fair comparison between excellent models, all experimental settings and equipment in this study were kept consistent with those of RS-Tran and Triple-CFN, including data volume, optimizer parameters, data augmentation methods (such as reducing the RAVEN image size to $80\times 80$), batch size, and other relevant hyperparameters.
TABLE II: Reasoning Accuracies on RAVEN and I-RAVEN.
Test Accuracy(%)
Model Average Center 2 $\times$ 2 Grid 3 $\times$ 3 Grid L-R U-D O-IC O-IG
SAVIR-T [24] 94.0/98.1 97.8/99.5 94.7/98.1 83.8/93.8 97.8/99.6 98.2/99.1 97.6/99.5 88.0/97.2
SCL [23, 24] 91.6/95.0 98.1/99.0 91.0/96.2 82.5/89.5 96.8/97.9 96.5/97.1 96.0/97.6 80.1/87.7
MRNet [21] 96.6/- -/- -/- -/- -/- -/- -/- -/-
RS-TRAN[28] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
Triple-CFN[32] 99.6/99.8 100.0/100.0 99.7/99.8 98.8/99.4 99.9/100.0 99.9/100.0 99.9/99.9 99.2/99.2
Lico-Net 99.7/99.8 100.0/100.0 99.7/99.8 98.9/99.5 99.9/100.0 99.9/100.0 99.9/99.9 99.4/99.6
VIII-B Experiment on Bongard-Logo
Consequently, we conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN as backbones for D2C, D3C, and D4C and using PMoC as backbone for D2C and D4C. The experimental results are presented in the table III. In the experiment with D3C, we conducted an additional experiment on its lightweight version, D3C-cos, to verify that replacing D3C with D3C-cos would not significantly affect performance. As can be seen in the table III, we did not conduct an experiment with Triple-CFN+D3C due to the large model size of Triple-CFN. This made it impractical to deploy the model using conceptual distribution and logical optimization with the Sinkhorn distance on our existing experimental equipment. This was the driving force behind proposing D3C-cos as an alternative to D3C in this paper. In addition, the combination of D4C and PMoC+Straw Pose-Transformer has not been implemented due to limitations in computing power.
TABLE III: Reasoning Accuracies of models on Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
SNAIL 59.2 56.3 60.2 60.1 61.3
ProtoNet 73.3 64.6 72.4 62.4 65.4
MetaOptNet 75.9 60.3 71.6 65.9 67.5
ANIL 69.7 56.6 59.0 59.6 61.0
Meta-Baseline-SC 75.4 66.3 73.3 63.5 63.9
Meta-Baseline-MoCo[42] 81.2 65.9 72.2 63.9 64.7
WReN-Bongard 78.7 50.1 50.9 53.8 54.3
SBSD 83.7 75.2 91.5 71.0 74.1
PMoC 92.0 92.6 97.7 78.3 75.0
PMoC+Straw Pose-Transformer 96.2 96.0 98.5 78.5 75.3
Triple-CFN 93.2 92.0 99.2 80.8 79.1
ResNet50 92.0 88.9 98.1 76.1 76.0
ResNet50+D2C 92.4 90.7 99.0 77.0 77.5
ResNet50+D3C 91.2 91.0 98.5 75.5 76.9
ResNet50+D2C+D3C 96.6 93.2 99.6 78.8 78.8
ResNet50+D3C-cos 95.2 90.2 98.8 79.0 78.1
ResNet50+D2C+D3C-cos 96.0 92.3 98.8 79.8 79.1
Triple-CFN+D2C 96.5 94.8 99.4 79.5 81.6
Triple-CFN+D3C-cos 97.0 96.2 99.4 82.0 79.7
Triple-CFN+D2C+D3C-cos 97.2 96.2 99.4 82.8 82.7
Triple-CFN+D4C 97.7 93.5 99.6 82.8 83.1
PMoC+D2C 93.9 95.0 98.1 78.5 78.4
PMoC+Straw Pose-Transformer+D2C 98.7 96.8 98.9 81.3 75.6
PMoC+D4C 96.0 95.2 99.2 84.8 82.8
Upon observing the results presented in the table III, we found that D2C, D3C, and D3C-cos made considerable contributions in the Bongard-Logo task, with D3C-cos emerging as an effective alternative to D3C. Additionally, D4C achieved remarkable results, despite not obtaining the best performance across all test data, possibly due to the inherent limitations of the PMoC model. However, compared to its backbone, PMoC, D4C demonstrated significant improvement and contribution, validating the effectiveness of defining tighter problem boundaries for the model. This approach warrants further investigation and development in future research.
VIII-C Experiment on PGM
In this paper, we conducted experiments related to Lico-Net on PGM. And the results are recorded in Tabel IV.
TABLE IV: Reasoning Accuracies of models on PGM.
Model Test Accuracy(%)
SAVIR-T [24] 91.2
SCL [23, 24] 88.9
MRNet [21] 94.5
RS-CNN[28] 82.8
RS-TRAN[28] 97.5
Triple-CFN[32] 97.8
Triple-CFN+Re-space layer[32] 98.2
Triple-CFN+Re-space layer+D2C 98.3(98.25)
Lico-Net 98.5(98.52)
Lico-Net+D2C 98.5(98.53)
D4C (with Lico-Net as the backbone) 98.6(98.64)
D4C glove-train(with Lico-Net as the backbone) 98.7(98.73)
Meta Triple-CFN+Re-space layer[32] 99.3
Triple-CFN+Re-space layer+D3C 99.4(99.36)
Triple-CFN+Re-space layer+D3C-cos 99.4(99.37)
Meta Lico-Net 99.3(99.30)
Lico-Net+D3C 99.4(99.35)
Lico-Net+D3C-cos 99.4(99.38)
In the experimental results recorded in Tabel IV, we can observe various phenomena. Lico-Net, in essence, addresses the imperfections within the Triple-CFN framework. In Triple-CFN, concepts and minimal reasoning units serve as queries and key-value pairs, respectively, for computing cross-attention outcomes. However, a fundamental flaw in its design lies in the logical hierarchy mismatch between concepts and minimal reasoning units. Specifically, concepts often require the combined expression of multiple minimal reasoning units. Despite this shortcoming, Triple-CFN remains a remarkable model. Lico-Net, born from the rectification of these imperfections, has achieved a little improvements in accuracy.
The combination of Lico-Net and D2C, and the combination of Triple-CFN and D2C, did not bring significant improvement, with the enhancement only reflected in the percentile of reasoning accuracy. This phenomenon was anticipated in this paper, as we believe it is due to the difference between RPM and Bongard-Logo problems. In RPM problems, different cases share the same concept space and attribute recognition method. However, the combination of D4C and D4C glove-train significantly improved Lico-Net’s performance. Additionally, the glove-shaped generator proved to be more suitable for generating negative representations of RPM problems, challenging Lico-Net and enabling it to achieve further refinement.
The combination of Lico-Net and D3C is noteworthy. To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted experiments using Meta Lico-Net. Unlike “Lico-Net+D3C”, which encodes the description of image progressive patterns into distributions using a Transformer-Encoder, Meta Lico-Net encodes them into vectors. To simplify, instead of dividing the encoded representation into mu and logvar, we directly utilize the representation as a vector. Furthermore, the cosine distance between progressive pattern vectors $\{\overline{P}_{nm}\}$ and the vectorized pattern description in Lico-Net is constrained using infoNCE. In fact, the process mentioned above is an emulation of how Triple-CFN evolved into Meta Triple-CFN. The structure of Meta Lico-Net is illustrate in figure 19. The intention behind establishing Meta Lico-Net and Triple-CFN+D3C is to demonstrate that, compared to vectors, distributions are more suitable for describing abstract concepts. To further validate the theme of D3C, we applied both D3C and D3C-cos to Triple-CFN and Lico-Net. Through experimentation, we found that D3C and D3C-cos brought improvements to Triple-CFN, elevating its performance to the level of Lico-Net.
Figure 19: Feedforward process of Meta Lico-Net
TABLE V: Progressive Pattern Reasoning Accuracies and of models on PGM.
Accuracy(%)
Model shape line answer
Meta Triple-CFN[32] 99.5 99.9 98.4
Meta Triple-CFN+Re-space layer[32] 99.7 99.9 99.3
Meta Lico-Net 99.7 99.9 99.3
Lico-Net+D3C 99.7 99.9 99.4
Lico-Net+D3C-cos 99.7 99.9 99.4
The D3C’s reasoning accuracys of progressive pattern are recorded in Table V, highlighting the strong interpretability prior to training possessed by both D3C and Lico-Net. This exceeds that of Rs-Tran and its derivative, Tran-clip. It must be acknowledged that the experiment of applying D3C, which is based on the Sinkhorn distance, to Lico-Net was not conducted to its fullest extent in this paper. Once the experimental results of Lico-Net+D3C were sufficient to illustrate our point, we halted the model training. This was due to the significant computational resources required, even with the experimental equipment used in this paper: four A100s graphics cards. The training involved numerous epochs and tedious individual training epochs. This was the intended purpose behind the design of D3C-cos as a replacement for D3C, as the power of Zeus needs to be conveyed by Hermes.
VIII-D Ablation experiment on D2C
In the final of this section, we devised an experiment to validate the D2C approach. Utilizing a 500-step DDPM model, we generated a substantial quantity of Bongard-Logo data, specifically 3,180 Bongard-Logo images. According to the DDPM theory, these images belong to the same distribution as the real data, albeit with unknown semantics and concepts. During the training of ResNet50 and Triple-CFN, for each feedforward pass, we introduced 14 randomly selected DDPM-generated images as auxiliary group examples for the Bongard-Logo cases. The experimental results were recorded in “D2C-” item of the Table VI.
TABLE VI: Reasoning Accuracies of models on Bongard-logo.
Accuracy(%)
Model Train FF BA CM NV
ResNet50 92.0 88.9 98.1 76.1 76.0
Triple-CFN 93.2 92.0 99.2 80.8 79.1
ResNet50+D2C 92.4 90.7 99.0 77.0 77.5
Triple-CFN+D2C 96.5 94.8 99.4 79.5 81.6
ResNet50+D2C- 93.5 84.2 99.5 77.0 76.2
Triple-CFN+D2C- 95.7 93.2 98.9 79.3 80.9
It can be observed from the experiment results recorded in Table VI that introducing generated samples as negative examples is difficult to fundamentally alter the network’s performance, as such behavior struggles to reorganize the concepts. In other words, D2C is restructuring the conceptual space, and the new concept space is highly coupled with the image style, with mutual exclusion between different cases of concepts. When the images generated by DDPM are introduced as auxiliary images into the Bongard-Logo case, they cannot achieve the above goal. Because D2C does not stop at the surface, that is, adding more auxiliary images to the Bongard-Logo case.
IX Conclusion
In summary, D3C advocates for representing concepts in abstract reasoning problems using distributions. It posits that, compared to vectors, distributions are better suited to capture the high-dimensionality, abstractness, and complexity inherent in human concepts, and experimental evidence underscores the superiority of this approach. Additionally, D2C presents a novel approach to tackling Bongard-Logo and RPM problems by resampling training data and introducing new loss function terms, effectively outlining new boundaries for solutions to these challenges. However, the extent to which these boundaries adapt to the specific problems remains uncertain.
D4C builds upon the principles of both D2C and D3C, exploring the roles of positive and negative examples in abstract reasoning problems from a perspective on distribution. Specifically, positive examples are utilized to depict the mean for deep networks, while negative examples contribute to describing the variance. Based on this understanding, D4C employs an adversarial approach, pitting the generator against the discriminator in repeated challenges. The generator continually produces negative examples that push the boundaries of the problem, while the discriminator refines its skills through repeated defense.
This iterative process not only enhances the discriminator’s ability to distinguish between genuine and generated examples but also improves the generator’s capacity to generate more sophisticated negative examples. Ultimately, this adversarial framework promotes the evolution of both components, leading to more robust and reliable solutions in abstract reasoning tasks.

### v9

@@
 This perspective enhances the performance of both Lico-Net and a baseline model excelling in Bongard-Logo tasks.
-To bolster the computational efficiency of D3C, we present the D3C-cos variant, offering a streamlined yet precise solution.
-Furthermore, we propose the D2C method, redefining conceptual boundaries within these domains and bridging the divide between high-level abstractions and their lower-dimensional counterparts.
-Finally, we extend our methodology to D4C, employing adversarial techniques to refine conceptual boundaries further and demonstrate substantial improvements in both RPM and Bongard-Logo challenges.
+To bolster the computational efficiency of D3C, we present the D3C-cos variant, offering a streamlined solution.
+Furthermore, we propose the D2C method, redefining concept boundaries within these domains and bridging the divide between high-level abstractions and their lower-dimensional counterparts.
+Finally, we extend our methodology to D4C, employing adversarial techniques to refine concept boundaries further and demonstrate substantial improvements in both RPM and Bongard-Logo challenges.
 Overall, our contributions present a fresh outlook and practical advancements in the field of abstract reasoning.
 Index Terms:
-Abstract reasoning, RPM problem, Bongard-logo problem.
+Abstract reasoning, Raven’s Progressive Matrices, Bongard-logo problem, Adversarial Learning.
 publicationid: pubid: $0$–$00$ 2021 IEEE
@@
 And $d$ represents the dimension of unit vectors $U_{ij}$.
-The introduction of the covariance matrix as an additional loss function is an effective technique inherited from Triple-CFN, embodying its essence.
-The entire feedforward process can be depicted in figure 8.
+The entire feedforward process of Lico-Net can be depicted in figure 8.
 Figure 8: Feedforward process of Lico-Net
@@
 The cases designers of Bongard-logo drew from high-dimensional human concepts and knowledge, carefully categorizing the Bongard-logo database into three distinct problem types: FF, BA, and HD.
+It is worth noting that the concept HD is decomposed into NV and CM during testing.
 This classification allows us to abstractly represent the distribution of the primary group in each Bongard-logo case as $p_{i}(x|\,y)$, while the distribution of the auxiliary group is denoted as $q_{i}(x|\,y)$.
@@
 Additionally, $x_{i7}$ represents the test image to be potentially assigned to the $i$-th primary group, and $x_{i14}$ represents the test image to be potentially assigned to the $i$-th auxiliary group.
-We copy a baseline from Triple-CFN, which refers to the baseline as $f_{\theta}(z|\,x)$ that maps samples $x_{ij}$ to latent variables $z_{ij}$.
+We copy a baseline from Triple-CFN [32], which refers to the baseline as $f_{\theta}(z|\,x)$ that maps samples $x_{ij}$ to latent variables $z_{ij}$.
 The baseline employs the InfoNCE loss function, which results in the $f_{\theta}(z|\,x)$ being able to solve Bongard-logo to a certain degree.
@@
 $=-\log\frac{{{e^{({z_{pri}}\cdot{\tilde{z}_{pri}})/t}}}}{{{e^{({z_{pri}}\cdot{\tilde{z}_{pri}})/t}}+\sum\nolimits_{m=1}^{M}{{e^{({z_{pri}}\cdot{z_{au{x_{m}}}})/t}}}}}$ (5)
-Where $z_{pri}$ and $\tilde{z}_{pri}$ represent distinct samples belonging to the primary group, and $z_{au{x_{m}}}$ encompasses all $m$ samples within the auxiliary group.
-$m$ denotes the number of negative samples within a minimum learning unit, while $t$ represents the temperature coefficient, and $t$ is set $10^{-3}$ in this paper.
-Through the strategic employment of the InfoNCE function, we can circumvent the need to explicitly estimate the distributions $p_{i}^{\prime}(z|\,y)$ and $q_{i}^{\prime}(z|\,y)$.
-Concurrently, this approach encourages enhanced similarity among the representations of the primary group encoded by the network $f_{\theta}(z|\,x)$ and guarantees a greater degree of mutual exclusion between the representations of the primary and auxiliary groups.
+Where $z_{pri}$ and $\tilde{z}_{pri}$ represent distinct samples belonging to the representations in primary group, and $z_{au{x_{m}}}$ encompasses all $m$ samples within the auxiliary group.
+$m$ denotes the number of negative samples within a minimum learning unit, while $t$ represents the temperature coefficient, and $t$ is set $10^{-3}$.
+This approach fosters a higher degree of similarity between $z_{pri}$ and $\tilde{z}_{pri}$ compared to the similarity between $z_{pri}$ and $z_{au{x_{m}}}$, thereby emphasizing the distinction between primary and auxiliary representations.
 This process aligns with the underlying logic of Bongard-logo.
 InfoNCE can be construed as a loss function devised for the clustering of vector groups.
-This paper reproduces the experiments for the baseline in Triple-CFN conducted based model ResNet50, which are individually implemented on the Bongard-logo database based on ResNet18.
-It focuses on the three concepts: ‘FF’, ‘BA’, and ‘HD’ for training, while concentrate on four test concept: ‘FF’, ‘BA’, ‘NV’ and ‘CM’.
-Subsequently, a combined database was created by merging these three individual databases, and experiments were performed on this combined database.
-The experimental results are presented in Table I.
-The feedforward process is shown in figure 9.
-The findings indicate that when presented with a combination of the three databases, the convolutional deep model experiences confusion among the three distinct human concepts.
-The varying degrees of reduction in the model’s reasoning accuracy on the FF and BA problems, as well as the increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion.
-Furthermore, this confusion is not attributable to insufficient network capacity, as even the ResNet50 employed in this study, which is more capacious than the ResNet18 utilized in Triple-CFN’ s baseline, falls victim to the same predicament.
+This paper reproduces the baseline experiments in Triple-CFN [32], which were originally conducted on the Bongard-Logo using the ResNet18 model.
+The baseline focuses on learning four concepts in Bongard-Logo — ’FF’, ’BA’, ’NV’, and ’CM’ — separately or in combination [32].
+To assess the impact of model size on the baseline experimental findings, we have replaced the backbone of the baseline experiments from ResNet18 to ResNet50 and repeated the experiments.
+The experimental results are presented in Table I, while the feedforward process is illustrated in Figure 9.
+The findings indicate that when presented with a combination of databases, both ResNet18 and ResNet50 experience confusion among the four distinct human concepts.
+The varying degrees of decrease in the models’ reasoning accuracy on the FF and BA problems, coupled with the increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion.
+Furthermore, this confusion cannot be attributed to insufficient network capacity, as even ResNet50, which is more capacious than the ResNet18 used in Triple-CFN’s baseline, is subject to the same challenge.
 Figure 9: Feedforward process baseline
@@
 ResNet18 on Separated Bongard-logo 97.9 99.0 75.0 72.8
-ResNet50 on Bongard-logo 88.9 98.1 76.1 76.0
+ResNet50 on Bongard-logo 88.9 98.0 76.1 76.0
 ResNet50 on Separated Bongard-logo 98.3 99.2 75.3 72.2
-This paper posits that, in the Bongard-Logo problem, the determination of image attributes relies heavily on the primary and auxiliary group samples within each case, which poses a fundamental challenge.
-It is precisely this multifaceted attribution of image properties that confuses deep networks when dealing with the three high-dimensional concepts inherent in the Bongard-Logo problem.
-Conversely, the RPM problem exhibits consistency in its recognition of image attributes.
-For instance, a white medium-sized triangle retains its attribute identification as such, regardless of its appearance in any RPM case option pool.
+After conducting backbone substitution experiments, this paper aligns with the perspective of Triple-CFN [32].
+Triple-CFN argues that the identification of image attributes in the Bongard-Logo problem relies heavily on primary and auxiliary samples in specific cases [32], posing challenges due to this context-dependent attribute determination.
+In other words, the reasoning identity of these images is contingent upon context-specific concepts, leading to multifaceted interpretations of image attributes.
+Consequently, deep networks encounter confusion when processing high-dimensional Bongard-Logo concepts [32].
+In contrast, the RPM problem exhibits consistency in its recognition of image attributes, such that a white medium-sized triangle retains its identity regardless of its appearance in any RPM case option pool.
 V Distribution Distance-based Discriminator of Case (D3C)
@@
 V-A D3C for Bongard-Logo
-The ideal approach to address the Bongard-Logo problem involves training an encoder capable of encoding images from a Bongard-Logo case into representations, denoted as $z$.
-In this context, the representations of positive instances (samples in primary group) belong to the distribution $p(z)$, while the representations of negative instances (samples in primary group) belong to $q(z)$.
-Importantly, there should be minimal overlap between $p(z)$ and $q(z)$ in terms of cross-measure.
-Therefore, this paper proposes the D3C method to enhance the performance of encoders in the form of $f(z|x)$ on Bongard-Logo problems.
+D3C proposes that distributions are more suitable as abstract representations of concepts compared to vectors.
+Therefore, D3C decides to estimate the concept in each image of a Bongard-Logo case using a deep network in the form of $f(z|\,x)$ and represent the concept with a distribution.
+Finally, the distances between these distributions are measured and used as the distances between concepts.
+These distances can be naturally applied to solve the Bongard-Logo problem.
+D3C is a plug-and-play method targeted at solving Bongard-Logo and similar problems for any network in the form of $f(z|\,x)$.
 Notably, this brings improvements for models such as Triple-CFN and ResNet50.
-Specifically, this paper hypothesizes that the representation of the primary group samples in a Bongard-Logo case follows a Gaussian distribution, while the representation of the auxiliary group samples follows a mixture of Gaussian distributions.
-We utilize an arbitrary backbone network $f(z|\,x)$ to approximate the distribution of image representations within Bongard-Logo cases.
-To achieve this, we modify the network to output parameters $\mu$ and $\sigma$ instead of a single vector, effectively converting $f(z|\,x)$ into $f(\mu,\sigma|\,x)$.
-The encoded vector is then divided into two equal-length segments, which we interpret as the mean and the logarithm of the variance, respectively.
-This approach normalizes the output vectors into Gaussian distributions and facilitates the distribution of image representations in Bongard-Logo cases.
-Our method bears resemblance to the Variational Autoencoder (VAE) framework.
+In practice, this paper hypothesizes that the concepts of the image samples in a Bongard-Logo case follows a Gaussian distribution.
+We utilize an arbitrary backbone network $f(z|\,x)$ to excract the concepts of images within Bongard-Logo cases in form of approximating distribution.
+To achieve this, we modify the network to output two vectors $\mu$ and $\sigma$ instead of a single vector $z$, effectively converting $f(z|\,x)$ into $f(\mu,\sigma|\,x)$.
+The encoded vector $z$ is then divided into two equal-length segments $\mu$ and $\log(\sigma^{2})$, which we interpret as the mean and the logarithm of the variance, respectively.
+This approach normalizes the output vectors into Gaussian distributions and represents concepts as distributions in Bongard-Logo cases.
+Our method bears resemblance to the Variational Autoencoder (VAE) framework[8].
 We argue that using distributions, rather than vectors, to describe image representations offers greater flexibility and accuracy in capturing the underlying complexity and uncertainty of abstract reasoning problems.
-With the estimated distributions of representations for each image in the Bongard-Logo cases, we compute the Sinkhorn distance between these distributions using reparameterization technique.
+With the estimated concepts for each image in the Bongard-Logo cases as distributions, we compute the Sinkhorn distance between these distributions using reparameterization technique[8].
 Following the logical structure of each Bongard-Logo case, we impose constraints on these distances for the backbone network $f(\mu,\sigma|\,x)$, aiming to train an effective discriminator.
@@
 $+\sum_{i=1}^{b}\sum_{{j}=1,}^{7}\sum_{\tilde{j}=7}^{14}\log({1-{e^{-\text{Sinkhorn}({z_{ijs}},\,{{z}_{i\tilde{j}s}})}}})$ (6)
-Where $b$ is the batchsize of training. $S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
+Where $b$ is the batchsize of training.
+The $\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\}$ represents $S$ samples drawn from the distribution $\{\mu_{ij},\sigma_{ij}|\,i\in[1,b],j\in[1,14]\}$ using the reparameterization trick, where the distribution $\{\mu_{ij},\sigma_{ij}|\,i\in[1,b],j\in[1,14]\}$ is obtained by mapping the Bongard-Logo image $\{x_{ij}|\,i\in[1,b],j\in[1,14]\}$ through the network $f(\mu,\sigma|\,x)$.
+The network $f(\mu,\sigma|x)$ is derived from the network $f(z|x)$ by D3C Method.
+$S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
 Figure 10: Feedforward process of D3C
@@
 Therefore, this paper proposes the concept of rules from the RPM database and represent them as distributions while also vectorizing the image progression patterns.
+This approach is precisely the original idea behind D3C, which posits that describing concepts using distributions is more appropriate than using vectors.
 We can leverage this approach to naturally constrain the image progression patterns within their corresponding rule distributions, thereby optimizing the deep model’s learning of rules and clarifying the boundaries among various concepts within the data.
@@
 Specifically, the initial portion of the feature vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the rule descriptions and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$.
-Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors termed as progression pattern vectors.
+Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ termed as progression pattern vectors.
 Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns should exhibit a high likelihood of adhering to the distribution outlined by their corresponding textual rules, and conversely, demonstrate a low probability of deviation.
@@
 VI-A D2C on Bongard-Logo problem
-In the Bongard-Logo problem, the determination of image attributes is significantly influenced by specific contextual concepts associated with each individual case.
+In the Bongard-Logo problem, the determination of image attributes is influenced by the specific contextual concepts associated with each individual case [32].
+Distinct concepts require the network to encode specific pixel configurations into differentiated attribute representations, potentially compelling the network to maintain mutually exclusive representation encoding strategies and methods [32].
+This scenario often leads to concept confusion within the network, rendering it uncertain about which concept to rely on when processing a Bongard-Logo case [32].
+In a nutshell, from the perspective of network, there exists mutual infringement and erroneous coupling among the high-dimensional human concepts in the Bongard-Logo.
 Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem.
+We posit that by successfully decoupling such that distinct cases rely on distinct concepts we may alleviate, to a certain extent, the issue of the network losing concept matching when confronted with a Bongard-Logo problem.
 To achieve this objective, we have redefined the concepts within the Bongard-Logo database, representing them as conditional probability distributions $p_{i}(x|\,Y_{i})$ and $q_{i}(x|\,Y_{i})$.
@@
 Through this concept reorganization, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases.
-In other words, each case is uniquely categorized based on its distinct set of concepts.
-By establishing a one-to-one correspondence between image styles and these redefined concepts, we aim to enable the network to seamlessly infer concepts from the pixel arrangements of images.
-This, in turn, aims to mitigate the impact of multiple attribute identifications on the solution of the Bongard-Logo problem.
-To foster concept antagonism between different problem domains, we have revised the data sampling techniques employed in mini-batch training methodologies.
+In other words, each case is categorized and solved based on its distinct concepts.
+By establishing a one-to-one correspondence between problem cases and these redefined concepts, we aim to enable the network to avoid uncertainty in determining which concept to rely on when solving a specific case.
+This approach aims to mitigate the impact of concept confusion on the solution for the Bongard-Logo problem.
+To foster concept antagonism between different Bongard-Logo cases, we have revised the data sampling techniques employed in mini-batch training methodologies.
 Furthermore, an additional loss function term has been devised to facilitate the reassignment of concepts within the network architecture.
-This approach, termed “D2C” (Dynamic Delineation of Concepts), endeavors to cultivate a state of concept contrast amidst diverse problem sets.
+This approach, termed “D2C” (Dynamic Delineation of Concepts), endeavors to cultivate a state of concept contrast amidst diverse Bongard-Logo cases.
 VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
-By augmenting the traditional sampling methods used for mini-batch training with an additional ancestral sampling procedure, this study identifies mutually exclusive pairs of Bongard-logo problems for training purposes.
+By incorporating an additional ancestral sampling procedure into the traditional mini-batch training sampling methods, this study constructs a Bongard-logo problem with reorganized concept space for training purposes.
 This constitutes a hot-swappable approach for training the Bongard-Logo solver, which is designated as phase 1 of the D2C method in this paper.
@@
 However, subsequent analysis revealed that this methodology did not tangibly enhance model accuracy.
-In summary, the phase 1 of the D2C approach embodies a straightforward yet effective concept: it introduces positive examples from other cases as negative instances in the current case.
+In summary, the phase 1 of the D2C approach embodies a straightforward yet effective concept: it introduces positive examples from other cases as negative samples in the current case.
 This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts.
@@
 Nevertheless, regarding the RPM problem, which emphasizes consistency in image attribute identification, this study maintains a pessimistic outlook on the efficacy of D2C in tackling RPM.
-It is worth emphasizing that the D2C and D3C methodologies are not mutually exclusive.
+In a nutshell, D2C is suitable for addressing the Bongard-Logo problem, where there exists mutual infringement and erroneous coupling of concepts across cases, but it is challenging to apply D2C to the RPM problem, which involves sharing a common set of concepts across cases.
+It is evident that D2C employs a precautionary approach, aiming to minimize mutual infringement and erroneous coupling among concepts, even at the cost of potential overcorrection.
+Finally, the D2C and D3C methodologies are not mutually exclusive.
 Specifically, D2C endeavors to redefine concept boundaries for reasoning problems.
@@
 This paper aims to delineate more appropriate concept boundaries for the RPM and Bongard-Logo problems.
-VII-A Negative examples are not merely intended to confound the discriminator
+VII-A Negative samples are not merely intended to confound the discriminator
 This paper posits that it is worthwhile to consider the significance of positive and negative examples in abstract reasoning problems.
@@
 Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern where positive examples yield high probabilities when substituted into the case’s problem statement, while negative examples yield low probabilities.
+In a nutshell, they are models with probabilistic labeling scheme.
 We can view this pattern in the following way: the discriminator constructs a solution distribution for the cases presented in abstract reasoning problems.
@@
 Where $x$ represents the positive or negative samples, $\mu$ and $\sigma$ represent the mean and standard deviation of the Gaussian distribution separately.
-It can be observed that when positive instances are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the mean squared loss between sample $x$ and the mean of the correct solution $\mu$.
-Conversely, when negative instances are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$ of the respective segment.
+It can be observed that when positive samples are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the mean squared loss between sample $x$ and the mean of the correct solution $\mu$.
+Conversely, when negative samples are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$ of the respective segment.
 Therefore, this paper posits that in reasoning problems, the establishment of positive examples aids deep models in estimating the mean of the correct solution distribution, while negative examples serve to estimate the variance.
@@
 This significance reflects a despairing fact that the upper limit of the discriminator’s reasoning accuracy in this abstract reasoning problem is determined by the quality of positive and negative examples in the question.
-VII-B High-quality negative examples lead to more rigorous concept boundaries
-The D2C methodology proposed in this study has established a novel problem boundary for deep networks by means of resampling negative instances and incorporating a fresh loss function.
-Experimental findings tentatively indicate that this boundary is better suited for scenarios characterized by the absence of conceptual supervision signals, as compared to the boundaries delineated by the original negative instances in Bongard-Logo.
+VII-B High-quality negative samples lead to more rigorous concept boundaries
+The D2C methodology proposed in this study has established a novel problem boundary for deep networks by means of resampling negative samples and incorporating a fresh loss function.
+Experimental findings tentatively indicate that this boundary is better suited for scenarios characterized by the absence of concept supervision signals, as compared to the boundaries delineated by the original negative samples in Bongard-Logo.
 Nonetheless, the precise tightness of the new boundary remains indeterminate.
 In essence, D2C introduces a new variance to the solution distribution of the network, the magnitude of which relative to the variance of the correct solution distribution is indeterminate.
-In the context of the Bongard-Logo challenge, the variance provided by D2C appears closer to that of the correct solution distribution in the absence of conceptual supervision signals, as compared to the variance described by the original negative instances.
+In the context of the Bongard-Logo challenge, the variance provided by D2C appears closer to that of the correct solution distribution in the absence of concept supervision signals, as compared to the variance described by the original negative samples.
 Evidently, the efficacy of D2C is contingent upon specific dataset attributes.
 As anticipated, empirical evidence suggests that this approach may be insufficient for delineating precise boundaries within the framework of the RPM problem.
-Contrary to the Bongard-Logo paradigm, RPM problems postulate a shared set of core concepts across distinct instances, ensuring discreteness and stability.
+Contrary to the Bongard-Logo paradigm, RPM problems postulate a shared set of core concepts across distinct cases, ensuring discreteness and stability.
 This shared nature precludes the utility of methods like D2C in addressing RPM challenges.
-Specifically, image properties within RPM instances are rigidly governed by definitions of “rules” or “progressive patterns,” maintaining consistent attributions across all instances.
-The Bongard-Logo problem exhibits varying attributions of image properties due to differing clustering bases among instances, thereby augmenting problem complexity and difficulty.
-Incorporating positive instances from other cases as negative examples can, to some extent, delineate conceptual boundaries between distinct cases, adhering to a cautious approach aimed at minimizing conceptual infringement.
+Specifically, image properties within RPM cases are rigidly governed by definitions of “rules” or “progressive patterns,” maintaining consistent attributions across all cases.
+The Bongard-Logo problem exhibits varying identifications and determinations of image attribute due to differing clustering bases among cases, thereby augmenting problem complexity and difficulty.
+Incorporating positive samples from other cases as negative examples can, to some extent, delineate concept boundaries between distinct cases, adhering to a cautious approach aimed at minimizing concept infringement.
 However, this strategy proves inapplicable in addressing challenges presented by RAVEN and PGM due to substantial disparities across multiple dimensions.
-In summary, the success of D2C in addressing the Bongard-Logo problem can be attributed to its indirect enhancement of the quality of negative instances within Bongard-Logo cases.
+In summary, the success of D2C in addressing the Bongard-Logo problem can be attributed to its indirect enhancement of the quality of negative samples within Bongard-Logo cases.
 Conversely, its limitations in tackling the RPM problem can also be ascribed to its insignificant improvement in the quality of negative examples within RPM scenarios.
 To address the limitations inherent in the D2C approach, this paper introduces the D4C (Defense-Driven Discrimination of Distributed Concepts) method.
-The underlying idea of D4C is to generate higher-quality negative instances and samples within abstract reasoning problems.
+The underlying idea of D4C is to generate higher-quality negative samples and samples within abstract reasoning problems.
 This approach aims to delineate a more appropriate problem boundary for the discriminator as compared to the original problem setup in abstract reasoning.
-By leveraging novel mechanisms and strategies, D4C aims to deliver more precise and reliable definitions of conceptual boundaries and foster further advancements in the realm of abstract reasoning.
+By leveraging novel mechanisms and strategies, D4C aims to deliver more precise and reliable definitions of concept boundaries and foster further advancements in the realm of abstract reasoning.
 VII-C Impelmention of D4C method
@@
 These discriminators seamlessly integrate with the D2C and D3C methodologies proposed in this study, and pertinent experiments have been conducted to explore their combinations.
-These discriminators excel in accurately identifying and classifying Bongard-Logo instances, significantly contributing to the advancement of research in this domain.
+These discriminators excel in accurately identifying and classifying Bongard-Logo cases, significantly contributing to the advancement of research in this domain.
 Regarding discriminators capable of handling adversarial tasks, the Wasserstein GAN, a prominent adversarial generative network, poses a noteworthy requirement.
@@
 In this figure, the “$f_{\theta}(z_{ij}|x_{ij})$” and “$g_{\omega}(p^{\prime}_{i}(z_{ij}|y)|\{z_{ij}\}_{j=1}^{6},z_{ij^{\prime}})$” represent the perceptron and logic calculation modules in PMoC, respectively[31].
-For the generator in Triple-CFN, it is essential to attenuate the generated representations.
+For the generator in Triple-CFN, it is essential to decay the generated representations.
 Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group.
-After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to attenuate the confusing representation vector.
-This attenuation process is necessary due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation.
-This makes it easier for a generator based on the Transformer encoder to generate realistic confusing representations, potentially causing Triple-CFN to collapse.
-Therefore, we have designed the aforementioned attenuation process.
+After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to decay the confusing representation.
+This decay process is necessary due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation.
+This mechanism makes it easier for a generator based on the Transformer encoder to generate realistic confusing representations, potentially causing Triple-CFN to collapse.
+Therefore, we have designed the aforementioned decay process.
 The alignment between the generator and Triple-CFN is illustrated in Figure 17.
@@
 To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations.
-Any inconsistencies or irregularities identified within these representations will be flagged as negative instances, indicating incorrect responses.
+Any inconsistencies or irregularities identified within these representations will be flagged as negative samples, indicating incorrect responses.
 Through this rigorous process, Lico-Net strives to accurately identify and reject perplexing representations, thereby augmenting its overall precision and reliability.
@@
 VIII-A Experiment on RAVEN
-Initially, we conducted experiments with Lico-Net on both RAVEN[15] and I-RAVEN[37] datasets to demonstrate its inherent superiority.
-To ensure a fair comparison between excellent models, all experimental settings and equipment in this study were kept consistent with those of RS-Tran and Triple-CFN, including data volume, optimizer parameters, data augmentation methods (such as reducing the RAVEN image size to $80\times 80$), batch size, and other relevant hyperparameters.
+Initially, we conducted experiments with Lico-Net on both RAVEN[15] and I-RAVEN[37] datasets to demonstrate its inherent potential.
+To ensure a fair comparison between excellent models, all experimental settings and equipment in this study were kept consistent with those of RS-Tran [28] and Triple-CFN [32], including data volume, optimizer parameters, data augmentation methods (such as reducing the RAVEN image size to $80\times 80$), batch size, and other relevant hyperparameters.
 TABLE II: Reasoning Accuracies on RAVEN and I-RAVEN.
@@
 RS-TRAN[28] 98.4/98.7 99.8/100.0 99.7/99.3 95.4/96.7 99.2/100.0 99.4/99.7 99.9/99.9 95.4/95.4
-Triple-CFN[32] 99.6/99.8 100.0/100.0 99.7/99.8 98.8/99.4 99.9/100.0 99.9/100.0 99.9/99.9 99.2/99.2
-Lico-Net 99.7/99.8 100.0/100.0 99.7/99.8 98.9/99.5 99.9/100.0 99.9/100.0 99.9/99.9 99.4/99.6
+Triple-CFN[32] 98.9/99.1 100.0/100.0 99.7/99.8 96.2/97.5 99.8/99.9 99.8/99.9 99.9/99.9 97.0/97.3
+Triple-CFN+Re-space[32] 99.4/99.6 100.0/100.0 99.7/99.8 98.0/99.1 99.9/100.0 99.9/100.0 99.9/99.9 98.5/99.0
+Lico-Net 99.5/99.7 100.0/100.0 99.7/99.8 98.2/99.2 99.9/100.0 99.9/100.0 99.9/99.9 98.8/99.2
+An interesting observation from the Table II indicates a consensus among existing excellent RPM solvers, including Lico-Net.
+Specifically, these solvers exhibit differing values of reasoning accuracy when tackling the RAVEN [15] and I-RAVEN [37] datasets.
+The I-RAVEN dataset was designed to address certain limitations in the option pool configuration of RAVEN by introducing improvements to its negative example setup.
+These enhancements include, but are not limited to, increasing the number of attributes that are altered in the negative examples and augmenting the magnitude of attribute deviations within them.
+However, from a human perspective, such additional setups do not alter the difficulty level of reasoning, and thus do not make I-RAVEN any harder or easier compared to RAVEN.
+This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the superior quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN.
+Such phenomenons echo the D4C’s spirit.
 VIII-B Experiment on Bongard-Logo
-Consequently, we conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN as backbones for D2C, D3C, and D4C and using PMoC as backbone for D2C and D4C.
+We conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN [32] as backbones for D2C, D3C(-cos), and D4C and using PMoC [31] as backbone for D2C and D4C.
+PMoC, a probabilistic labeling model, does not meet the requirements of D3C(-cos), as D3C(-cos) demands a backbone in the form of $f(z|x)$.
 The experimental results are presented in the table III.
-In the experiment with D3C, we conducted an additional experiment on its lightweight version, D3C-cos, to verify that replacing D3C with D3C-cos would not significantly affect performance.
-As can be seen in the table III, we did not conduct an experiment with Triple-CFN+D3C due to the large model size of Triple-CFN.
-This made it impractical to deploy the model using conceptual distribution and logical optimization with the Sinkhorn distance on our existing experimental equipment.
-This was the driving force behind proposing D3C-cos as an alternative to D3C in this paper.
-In addition, the combination of D4C and PMoC+Straw Pose-Transformer has not been implemented due to limitations in computing power.
+Upon observing the results presented in the table III, we found that D2C, D3C, and D3C-cos made considerable contributions in the Bongard-Logo task.
+As can be seen in the table III, replacing D3C with D3C-cos to integrate with ResNet50 would not significantly affect performance.
+We argue that, from a results-oriented perspective, D3C-cos can be considered as a effective alternative to D3C.
+Notably, we did not conduct an experiment with Triple-CFN+D3C and Triple-CFN+D2C+D3C.
+This is because the large number of Triple-CFN parameters makes integrating the D3C approach impractical on our current experimental equipment.
+This was the driving force that led us to propose D3C-cos as an alternative to D3C in this paper.
 TABLE III: Reasoning Accuracies of models on Bongard-logo.
@@
 SBSD 83.7 75.2 91.5 71.0 74.1
-PMoC 92.0 92.6 97.7 78.3 75.0
-PMoC+Straw Pose-Transformer 96.2 96.0 98.5 78.5 75.3
-Triple-CFN 93.2 92.0 99.2 80.8 79.1
-ResNet50 92.0 88.9 98.1 76.1 76.0
-ResNet50+D2C 92.4 90.7 99.0 77.0 77.5
-ResNet50+D3C 91.2 91.0 98.5 75.5 76.9
-ResNet50+D2C+D3C 96.6 93.2 99.6 78.8 78.8
-ResNet50+D3C-cos 95.2 90.2 98.8 79.0 78.1
-ResNet50+D2C+D3C-cos 96.0 92.3 98.8 79.8 79.1
-Triple-CFN+D2C 96.5 94.8 99.4 79.5 81.6
-Triple-CFN+D3C-cos 97.0 96.2 99.4 82.0 79.7
-Triple-CFN+D2C+D3C-cos 97.2 96.2 99.4 82.8 82.7
-Triple-CFN+D4C 97.7 93.5 99.6 82.8 83.1
-PMoC+D2C 93.9 95.0 98.1 78.5 78.4
-PMoC+Straw Pose-Transformer+D2C 98.7 96.8 98.9 81.3 75.6
+PMoC 92.0 90.6 97.7 77.3 76.0
+Triple-CFN 93.2 92.0 98.2 78.0 78.1
+ResNet50 (infoNCE baseline) 92.0 88.9 98.0 76.1 76.0
+ResNet50+D2C 92.6 90.5 98.2 77.0 77.2
+ResNet50+D3C 92.3 90.6 98.3 76.5 76.8
+ResNet50+D2C+D3C 93.0 91.0 98.5 77.5 77.5
+ResNet50+D3C-cos 92.1 90.2 98.2 76.9 76.4
+ResNet50+D2C+D3C-cos 92.9 90.8 98.4 77.5 77.4
+ResNet50+D4C 93.2 90.8 98.5 77.2 77.3
+Triple-CFN+D2C 93.9 92.7 98.5 78.2 78.5
+Triple-CFN+D3C-cos 93.7 92.5 98.3 78.2 78.2
+Triple-CFN+D2C+D3C-cos 94.1 93.0 98.5 78.2 78.6
+Triple-CFN+D4C 94.3 92.8 98.8 78.7 79.0
+PMoC+D2C 92.5 91.2 98.0 77.7 76.7
+PMoC+D4C 93.0 92.6 98.2 78.0 77.5
+Additionally, D4C has imparted varying degrees of improvement to Triple-CFN, ResNet50, and PMoC, with its effectiveness rivaling that of the combined application of D2C and D3C.
+However, D4C did not yield an ideal improvement on the FF concept for the Bongard-logo problem when applied to Triple-CFN and ResNet50.
+This study hypothesizes that the suboptimal performance may be attributed to the fact that D4C was specifically designed and tailored for discriminative models with a probabilistic labeling scheme.
+Conversely, Triple-CFN and ResNet50 do not inherently adhere to such a framework.
+It has been mentioned that when integrating D4C with Triple-CFN and ResNet50, a decay mechanism was incorporated.
+This mechanism is a compromise made by D4C to accommodate the non-probabilistic labeling scheme of these models.
+From the perspective of experimental results, this decay mechanism does not appear to be a fully developed mechanism.
+Exploring more advanced designs to facilitate the adaptation of D4C to non-probabilistic labeling models, such as Triple-CFN, deserves attention as future work.
+In contrast to Triple-CFN, PMoC [31] serves as an intuitive probabilistic labeling model for the Bongard-Logo problem.
+Correspondingly, D4C provides a more significant enhancement when applied to PMoC, owing to its scheme.
+D4C demonstrated improvement and contribution, validating the effectiveness of defining tighter problem boundaries for the model with a probabilistic labeling scheme.
+To integrate D4C with Triple-CFN, we devised a representation decay mechanism that, despite its imperfections, is sufficiently effective.
+This approach warrants further investigation and development in future research.
+Finally, this paper notes that augmentation experiments were conducted for both PMoC [31] and Triple-CFN [32].
+These experiments involved expanding the data volume of the Bongard-Logo dataset using data augmentation techniques, specifically including rotation and flipping.
+It should be emphasized that these data augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [31].
+Rather, the aim of these experiments was to investigate the potential of the models with a larger dataset.
+Since PMoC posits that Transformer-based visual processing models require substantial data as a foundation, a point also echoed by ViT [17] and RS-Tran [28].
+Furthermore, both PMoC and Triple-CFN agree that the utilization of these data augmentation methods serves to enrich the expressions of concepts at the pixel level, thereby challenging the network’s abilities to summarize, derive, accommodate, and invert concepts [31, 32].
+This paper conducted experiments to verify the effectiveness of D4C with the expanded dataset.
+The results are recorded in Table IV.
+TABLE IV: Reasoning Accuracies of D4C on Augmented Bongard-logo.
+Accuracy(%)
+Model Train FF BA CM NV
+PMoC 94.5 92.6 98.0 78.3 76.5
+Triple-CFN 94.9 93.0 99.2 80.8 79.1
 PMoC+D4C 96.0 95.2 99.2 84.8 82.8
-Upon observing the results presented in the table III, we found that D2C, D3C, and D3C-cos made considerable contributions in the Bongard-Logo task, with D3C-cos emerging as an effective alternative to D3C.
-Additionally, D4C achieved remarkable results, despite not obtaining the best performance across all test data, possibly due to the inherent limitations of the PMoC model.
-However, compared to its backbone, PMoC, D4C demonstrated significant improvement and contribution, validating the effectiveness of defining tighter problem boundaries for the model.
-This approach warrants further investigation and development in future research.
+Triple-CFN+D4C 97.7 94.8 99.6 82.8 83.1
+The results presented in Table IV demonstrate that the integration of D4C into PMoC, which is under a probabilistic labeling scheme, has led to a significant enhancement in performance.
+The increase in data volume has enabled D4C to further narrow the gap between PMoC and Triple-CFN.
 VIII-C Experiment on PGM
-In this paper, we conducted experiments related to Lico-Net on PGM. And the results are recorded in Tabel IV.
-TABLE IV: Reasoning Accuracies of models on PGM.
+In this paper, we conducted experiments related to Lico-Net on PGM. And the results are recorded in Tabel V.
+TABLE V: Reasoning Accuracies of models on PGM.
 Model Test Accuracy(%)
@@
 Lico-Net+D2C 98.5(98.53)
-D4C (with Lico-Net as the backbone) 98.6(98.64)
-D4C glove-train(with Lico-Net as the backbone) 98.7(98.73)
+D4C (with Lico-Net as the backbone) 98.7(98.70)
+D4C glove-train(with Lico-Net as the backbone) 98.8(98.82)
+RS-Tran+Tranclip[28] 99.0
+Meta Triple-CFN[32] 98.4
 Meta Triple-CFN+Re-space layer[32] 99.3
-Triple-CFN+Re-space layer+D3C 99.4(99.36)
-Triple-CFN+Re-space layer+D3C-cos 99.4(99.37)
-Meta Lico-Net 99.3(99.30)
+Lico-Net+D3C- 99.3(99.30)
 Lico-Net+D3C 99.4(99.35)
 Lico-Net+D3C-cos 99.4(99.38)
-In the experimental results recorded in Tabel IV, we can observe various phenomena.
-Lico-Net, in essence, addresses the imperfections within the Triple-CFN framework.
-In Triple-CFN, concepts and minimal reasoning units serve as queries and key-value pairs, respectively, for computing cross-attention outcomes.
-However, a fundamental flaw in its design lies in the logical hierarchy mismatch between concepts and minimal reasoning units.
-Specifically, concepts often require the combined expression of multiple minimal reasoning units.
-Despite this shortcoming, Triple-CFN remains a remarkable model.
-Lico-Net, born from the rectification of these imperfections, has achieved a little improvements in accuracy.
+From the experimental results presented in Table V, several notable phenomena emerge.
+In Triple-CFN, both progressive patterns and minimal reasoning units are positioned at the same logical level [32].
+However, a significant design flaw arises from the mismatch in the logical hierarchy between these patterns and units.
+More specifically, progressive patterns often require the combined expression of multiple minimal reasoning units.
+Despite this limitation, Triple-CFN still stands as an impressive model.
+Naturally, leveraging the inductive bias inherent to the RPM problem as the guideline in its design, Lico-Net has achieved even more remarkable results than Triple-CFN in the RPM challenges.
 The combination of Lico-Net and D2C, and the combination of Triple-CFN and D2C, did not bring significant improvement, with the enhancement only reflected in the percentile of reasoning accuracy.
@@
 In RPM problems, different cases share the same concept space and attribute recognition method.
+The D2C approach, designed to address issues of concept interference and erroneous coupling among cases, is not applicable to RPM problems.
 However, the combination of D4C and D4C glove-train significantly improved Lico-Net’s performance.
-Additionally, the glove-shaped generator proved to be more suitable for generating negative representations of RPM problems, challenging Lico-Net and enabling it to achieve further refinement.
+Additionally, the glove-shaped generator proved to be more suitable for generating negative representations of RPM problems for Lico-Net and enabling it to achieve further refinement.
+The core innovation of this paper, D4C, has yielded promising results on both the RPM and Bongard-Logo problems, demonstrating the significant contribution of this work.
 The combination of Lico-Net and D3C is noteworthy.
-To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted experiments using Meta Lico-Net.
-Unlike “Lico-Net+D3C”, which encodes the description of image progressive patterns into distributions using a Transformer-Encoder, Meta Lico-Net encodes them into vectors.
+To validate the theme of D3C, we applied both D3C and D3C-cos to Lico-Net.
+Through experimentation, we found that D3C and D3C-cos brought improvements to Lico-Net.
+To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted ablation experiments using “Lico-Net+D3C-”.
+Unlike “Lico-Net+D3C”, which encodes the description of image progressive patterns into distributions using a Transformer-Encoder, “Lico-Net+D3C-” encodes them into vectors.
 To simplify, instead of dividing the encoded representation into mu and logvar, we directly utilize the representation as a vector.
 Furthermore, the cosine distance between progressive pattern vectors $\{\overline{P}_{nm}\}$ and the vectorized pattern description in Lico-Net is constrained using infoNCE.
-In fact, the process mentioned above is an emulation of how Triple-CFN evolved into Meta Triple-CFN.
-The structure of Meta Lico-Net is illustrate in figure 19.
-The intention behind establishing Meta Lico-Net and Triple-CFN+D3C is to demonstrate that, compared to vectors, distributions are more suitable for describing abstract concepts.
-To further validate the theme of D3C, we applied both D3C and D3C-cos to Triple-CFN and Lico-Net.
-Through experimentation, we found that D3C and D3C-cos brought improvements to Triple-CFN, elevating its performance to the level of Lico-Net.
-Figure 19: Feedforward process of Meta Lico-Net
-TABLE V: Progressive Pattern Reasoning Accuracies and of models on PGM.
+In essence, the Lico-Net+D3C- can be interpreted as fixing the variance of the distributions established for all concepts in Lico-Net+D3C-cos to 0.
+The structure of Lico-Net+D3C- is illustrate in figure 19.
+Figure 19: Feedforward process of “Lico-Net+D3C-
+TABLE VI: Progressive Pattern Reasoning Accuracies and of models on PGM.
 Accuracy(%)
 Model shape line answer
+RS-Tran+Tranclip[28] - - 99.0
 Meta Triple-CFN[32] 99.5 99.9 98.4
 Meta Triple-CFN+Re-space layer[32] 99.7 99.9 99.3
-Meta Lico-Net 99.7 99.9 99.3
+Lico-Net+D3C- 99.7 99.9 99.3
 Lico-Net+D3C 99.7 99.9 99.4
 Lico-Net+D3C-cos 99.7 99.9 99.4
-The D3C’s reasoning accuracys of progressive pattern are recorded in Table V, highlighting the strong interpretability prior to training possessed by both D3C and Lico-Net.
-This exceeds that of Rs-Tran and its derivative, Tran-clip.
+The D3C’s reasoning accuracys of progressive pattern are recorded in Table VI, highlighting the interpretability expressed by Lico-Net+D3C(-cos).
 It must be acknowledged that the experiment of applying D3C, which is based on the Sinkhorn distance, to Lico-Net was not conducted to its fullest extent in this paper.
@@
 This was the intended purpose behind the design of D3C-cos as a replacement for D3C, as the power of Zeus needs to be conveyed by Hermes.
-VIII-D Ablation experiment on D2C
-In the final of this section, we devised an experiment to validate the D2C approach.
+Despite achieving remarkable results in terms of reasoning accuracy and interpretability of progressive patterns with Lico-Net+D3C or D3C-cos, it is noteworthy that the improvement over previous works involving Meta data as additional auxiliary labels is not particularly striking.
+The possible reason for this is that the introduction of Meta data significantly reduces the difficulty of the RPM problem Moreover, it is anticipated that a greater abundance of supervisory signals would correlate with enhanced intelligent performance.
+However, Lico-Net+D3C still demonstrates distinct advantages that surpass other methods involving metadata.
+Specifically, Tranclip [28] utilizes metadata to pre-train a set of warm-start parameters for RS-Tran [28].
+In other words, Tranclip fails to simultaneously endow RS-Tran with both interpretability of progressive patterns and high reasoning accuracy.
+Meta Triple-CFN [32] also relies on its specially designed Re-space Layer to enhance both its interpretability and reasoning accuracy.
+Similarly, the Re-Space Layer requires a fully trained Meta Triple-CFN to provide warm-start parameters [32].
+In contrast, Lico-Net operates effectively without the necessity of a Re-space layer or a warm-start process, thereby simplifying the training procedure to a certain extent.
+This indicates that Lico-Net and D3C are more direct and effective, eliminating the need for indirect approaches.
+This ”directness” reflects a more coherent and self-consistent logic of Lico-Net and D3C compared to other methods incorporating metadata.
+This paper argues that this is not an insignificant contribution, as warm-start behavior may introduce unexpected instabilities, which are not a concern for Lico-Net and D3C.
+For example, when Tranclip [28] uses Meta data to produce warm-start parameters for RS-Tran, early stopping is required [28].
+Neglecting this step may hinder Tranclip from improving or potentially worsen the reasoning performance of RS-Tran.
+The Rs-space Layer [32] relies on a well-trained Meta Triple-CFN to provide warm-start parameters [32], a process that demands considerable training time and epochs.
+Inadequate training of these warm-start parameters can expose Meta Triple-CFN to risks akin to those faced by RS-Tran.
+The stopping timing for the preparation of both warm-start parameters requires manual determination and control, which brings numerous potential issues.
+In contrast, Lico-Net effectively mitigates these potential issues.
+VIII-D Ablation experiment on D2C and D4C
+In the final of this section, we devised an experiment to validate the D2C and D4C approach.
 Utilizing a 500-step DDPM model, we generated a substantial quantity of Bongard-Logo data, specifically 3,180 Bongard-Logo images.
@@
 During the training of ResNet50 and Triple-CFN, for each feedforward pass, we introduced 14 randomly selected DDPM-generated images as auxiliary group examples for the Bongard-Logo cases.
-The experimental results were recorded in “D2C-” item of the Table VI.
-TABLE VI: Reasoning Accuracies of models on Bongard-logo.
+The experimental results were recorded in “D2C-” item of the Table VII.
+TABLE VII: Reasoning Accuracies of models on Bongard-logo.
 Accuracy(%)
 Model Train FF BA CM NV
-ResNet50 92.0 88.9 98.1 76.1 76.0
-Triple-CFN 93.2 92.0 99.2 80.8 79.1
-ResNet50+D2C 92.4 90.7 99.0 77.0 77.5
-Triple-CFN+D2C 96.5 94.8 99.4 79.5 81.6
-ResNet50+D2C- 93.5 84.2 99.5 77.0 76.2
-Triple-CFN+D2C- 95.7 93.2 98.9 79.3 80.9
-It can be observed from the experiment results recorded in Table VI that introducing generated samples as negative examples is difficult to fundamentally alter the network’s performance, as such behavior struggles to reorganize the concepts.
-In other words, D2C is restructuring the conceptual space, and the new concept space is highly coupled with the image style, with mutual exclusion between different cases of concepts.
-When the images generated by DDPM are introduced as auxiliary images into the Bongard-Logo case, they cannot achieve the above goal.
-Because D2C does not stop at the surface, that is, adding more auxiliary images to the Bongard-Logo case.
+ResNet50 (infoNCE baseline) 92.0 88.9 98.0 76.1 76.0
+Triple-CFN 93.2 92.0 98.2 78.0 78.1
+ResNet50+D2C- 91.6 90.0 97.9 76.3 76.1
+Triple-CFN+D2C- 93.0 92.1 98.2 77.9 78.2
+It can be observed from the experiment results recorded in Table VII that introducing generated samples as negative examples is difficult to fundamentally alter the network’s performance, as such behavior struggles to re-delineate the concepts.
+DDPM-generated data, sharing the same distribution as real data but with inherent uncertainties of concept and attribution, are limited in contributing to the delineation of concept boundaries.
+D4C posits that the accuracy achieved by the discriminator in abstract reasoning problems is closely related to the quality of the positive and negative examples set in the task.
+The established examples serve to delineate the concept boundaries for abstract reasoning challenges.
+D2C, while attempting to define new concept boundaries for such problems, represents an imperfect methodology.
+Conversely, D4C emerges as a relatively more comprehensive approach.
+This ablation experiment aims to verify that, within the context of Bongard-Logo problems, the provision of meaningless negative examples—examples that do not meaningfully contribute to defining the boundaries of reasoning tasks—fails to enhance the discriminator’s effectiveness.
+This finding echoes similar observations made in the RAVEN dataset, reinforcing the significance and rationale behind the design of D4C.
+Furthermore, it reaffirms D4C’s assertion regarding the importance of high-quality positive and negative examples in abstract reasoning tasks.
+It also provides indirect evidence that the approaches of D2C and D4C extend beyond merely introducing additional negative examples.
+Instead, they aim to introduce negative examples that better delineate the boundaries of the problem.
 IX Conclusion
@@
 The generator continually produces negative examples that push the boundaries of the problem, while the discriminator refines its skills through repeated defense.
-This iterative process not only enhances the discriminator’s ability to distinguish between genuine and generated examples but also improves the generator’s capacity to generate more sophisticated negative examples.
+This iterative process not only enhances the discriminator’s ability to distinguish between positive and negative examples but also improves the generator’s capacity to generate more sophisticated negative examples.
 Ultimately, this adversarial framework promotes the evolution of both components, leading to more robust and reliable solutions in abstract reasoning tasks.
+For future, the D4C method presents a novel approach for abstract reasoning problems and even other discriminative tasks.
+Its exploration of the identities of positive and negative examples may provide insights for the development of other works.

### v12

@@
-D4C Glove-train: Solving the RPM and Bongard-logo Problem by Circumscribing and Building Distribution for Concepts
+D4C Glove-train: Solving the Visual Abstract Reasoning Problems by Circumscribing and Building Distribution for Concepts
 Abstract
@@
 I Introduction
-Deep neural networks have achieved remarkable success in various domains, including computer vision[1, 2, 3], natural language processing[4, 5, 6], generative models[7, 8, 9], visual question answering[10, 11], and abstract reasoning[14, 12, 13].
-This success is primarily attributed to the remarkable capabilities of deep learning, which can recognize text, images, and sound data by learning the inherent patterns and representational hierarchies within sample data, thereby demonstrating analytical learning abilities similar to humans.
-Especially when addressing complex pattern recognition problems, the powerful capabilities of deep learning enable it to mimic human activities, further advancing the progress of artificial intelligence-related technologies.
-Its outstanding performance in speech and image recognition has surpassed previous technologies in terms of recognition accuracy.
-However, despite these impressive accomplishments in deep learning, there are still a series of unresolved issues and challenges that require further research and exploration.
-For instance, research is ongoing on how to utilize the results of upperlayer training as initialization parameters for lowerlayer training, aiming to enhance the efficiency of deep model training.
-Additionally, layer-wise initialization training with unsupervised learning constitutes a significant strategy within the field of deep learning.
-Especially in the domain of graphical abstract reasoning, the advancement of deep learning remains an active area of research.
-Although deep learning has achieved remarkable success in intelligent vision tasks, elevating machine intelligence to potentially higher levels, the academic community’s research on the abstract reasoning capabilities of deep learning has not ceased.
-Conversely, this has propelled the emergence of graphical reasoning problems.
-Graphical reasoning problems pose unique challenges for deep learning, including small and difficult-to-label datasets, the requirement for deep understanding of spatial relationships, shapes, and colors in graphics, demanding strong reasoning and innovation capabilities from models, necessitating a certain level of explanatory ability from models, and posing higher demands on models due to problem complexity and diversity.
-To address these challenges, we need to further explore novel model architectures, training methodologies, and optimization strategies.
-This may encompass directions such as designing more effective deep learning models, refining model training methods, and enhancing dataset quality.
-Simultaneously, we also need to develop a deeper understanding of the essence and characteristics of graphical reasoning problems so that we can better leverage deep learning techniques to tackle these issues.
-In summary, while deep learning has made certain strides in graphical abstract reasoning, numerous problems still await our investigation and resolution.
-The Ravens Progressive Matrices (RPM)[14] and Bongard problems[12, 13], for example, present learning requirements that span from perceptual processes to inferential capabilities.
+Deep neural networks have made remarkable achievements in various fields, including computer vision[1, 2, 3], natural language processing[4, 5, 6], generative model[7, 8, 9], visual question answering[10, 11], and visual abstract reasoning[14, 12, 13].
+However, deep learning faces challenges in visual abstract reasoning, as it requires models to capture complex patterns, handle compositionality, and possess abstraction capabilities.
+Models must also be flexible and scalable to adapt to diverse graphical structures.
+Existing deep learning models often lack the necessary architectural rigor and representational power to address these challenges adequately.
+Overcoming these obstacles demands innovative model designs, enhanced representation techniques, and efficient algorithms.
+Progress in this area can lead to significant advancements in graphical reasoning and practical solutions.
+Due to the higher demands of reasoning tasks compared to recognition, perception, and regression tasks, numerous prestigious research institutions are consecutively dedicating their efforts to studying abstract reasoning problems.
+They are focused on exploring diverse methodologies to address these challenges from multiple perspectives, with the expectation that the techniques developed in this field can be applied to benefit other areas of artificial intelligence.
+Notably, Ravens Progressive Matrices (RPM) challenges[14] and Bongard problems[12, 13] involve learning that ranges from basic perception to complex reasoning.
 I-A RAVEN and PGM
-The RAVEN database[15], a distinctive collection of Ravens Progressive Matrices (RPM) problems[14], stands out for its particular design elements.
-Each RAVEN problem consists of sixteen images, with eight forming the essence of the problem, termed as the stem, while the remaining eight constitute a pool of alternate choices.
-Subjects are tasked with selecting appropriate images from this pool, which, together with the stem images, form a 3×3 matrix of graphical patterns.
-The matrix is characterized by a progressive pattern of geometric images in the row direction, encompassing an underlying abstract concept.
-Figure 1 presents a concrete instantiation of a RAVEN case’s construction, shedding light on the methodology and general attributes of these problems.
+The RAVEN database[15]—a unique compilation of Ravens Progressive Matrices (RPM) problems[14]—stands out due to its distinctive design elements.
+In each RAVEN problem, sixteen images are presented, with eight images forming the core of the challenge, referred to as the “statement.” The remaining eight images serve as a selection pool for subjects to choose from.
+The task requires selecting appropriate images from this pool to create, alongside the stem images, a 3x3 matrix of graphical patterns.
+This matrix showcases a progressive geometric pattern along the rows, encapsulating an abstract concept known as the “rule.” Figure 1 illustrates a specific example of a RAVEN case, providing insight into the methodology and general characteristics of these problems.
+Akin to RAVEN, PGM utilizes a format featuring eight images for the stem and answer pool.
+Nevertheless, the “rule” concept in PGM encompasses both row and column progressions.
+Figure 2 showcases a representative example of a PGM case.
 Figure 1: RAVEN case
-Within the design of RAVEN problems, geometric images often encompass human-predefined concepts such as ”shape” and ”color”.
-However, these concepts are further abstracted in RAVEN problems into a finite, countable, and well-defined set of ”visual attributes”.
-The introduction of the term ”rule” facilitates the description of how these ”visual attributes” progress in a particular manner.
-It’s worth noting that not all visual attributes are constrained by this ”rule”; some may be assigned randomly, posing additional challenges for deep model reasoning.
-Constructing a comprehensive RAVEN problem entails initially randomly selecting a rule sample from a predefined rule set.
-Subsequently, corresponding visual attribute values are devised based on this chosen ”rule”.
-For attributes free from rules, their values are assigned randomly.
-Once these designs are in place, images are rendered according to the derived ”attribute” information.
-The RAVEN database not only encompasses problems with singular rules but also those with dual rules, elevating the level of difficulty and challenge.
-Specifically, the RAVEN database includes sub-databases with single rules such as center single, distribute four (G2$\times$2), distribute nine (G3×3), among others.
-In problems with a singular rule, the progressive variation of entity attributes within the images is governed by one set of rules.
-Conversely, in problems involving dual rules, like in center single out center single (O-IC), up center single down center single (U-D), etc., the attribute progression must satisfy two independent rule sets.
-The design approach of PGM[16] and RAVEN problems demonstrates a striking resemblance, with both frameworks employing a problem format consisting of a stem composed of eight images and an answer pool formed by another eight images.
-It’s noteworthy that within the context of PGM problems, the concept of ”rule” assumes a more profound significance.
-This concept not only delineates the progressive pattern of ”visual attributes” along the row direction within the matrix but also imposes constraints on the progressive trend along the column direction.
-A representative instance of a PGM case is illustrated in the accompanying figure 2.
 Figure 2: PGM case
-Given this complexity, the challenges associated with RPM problems extend beyond the mere exploration of visual attributes at various levels.
-They also encompass the crucial task of effectively inducing and learning the progressive patterns exhibited by these ”visual attributes”.
-Mastering and comprehending these progressive patterns hold profound implications for successfully tackling RPM problems, thereby underscoring their importance in this cognitive assessment domain.
 I-B Bongrad-logo
 The Bongard problem[12, 13] stands in stark contrast to the RPM problem, falling under the umbrella of small sample learning problems.
-Typically, Bongard problems consist of multiple images divided into two distinct groups: primary and secondary.
-Images within the primary group all convey an abstract concept that is bound by a specific rule, whereas images in the secondary group deviate from these rules in varying degrees.
+Typically, Bongard problems consist of multiple images divided into two distinct groups: primary and auxiliary.
+Images within the primary group all convey an abstract concept that is bound by a specific rule, whereas images in the auxiliary group deviate from these rules in varying degrees.
 The crux of solving Bongard problems lies in the ability of deep learning algorithms to accurately classify ungrouped images into their respective categories.
 Notably, Bongard-logo[13] problems, a concrete instantiation of Bongard problems within the realm of abstract reasoning and a combination of a lot of Bongard-logo cases, pose a formidable challenge in terms of inferential complexity.
-Each Bongard-logo case comprises 14 images, including six images from both the primary and secondary groups, along with two images awaiting classification.
+Each Bongard-logo case comprises 14 images, including six images from both the primary and auxiliary groups, along with two images awaiting classification.
 The grouping criterion primarily relies on the geometric shapes present within the images and their arrangement.
@@
 Figure 3: Bongard-Logo case
-Moreover, Bongard-logo problems can be further subclassified into three categories based on the concept domains they encompass: 1) Free-form problems (FF), where shapes are composed of randomly selected action strokes, with each image potentially housing one or two such shapes; 2) Basic shape problems (BA), which require recognition of a combination of one or two shape categories represented within the shape patterns; and 3) High-level concept problems (HD), designed to test a model’s ability to discover and reason about abstract concepts, such as convexity, symmetry, among others.
-Figure 1 exemplary presents a specific Bongard-logo case alongside its underlying principles.
+Bongard-logo problems can be categorized into three parts based on concept domains: 1) Free-form problems (FF), featuring shapes made of random action strokes; 2) Basic shape problems (BA), recognizing pattern of shape combinations; 3) High-level concept problems (HD), assessing abstract concept reasoning, like convexity and symmetry.
 II Related work
@@
 By combining effective methods, RS-CNN and RS-TRAN[28] have achieved impressive results on the RPM problem.
-Triple-CFN implicitly extracts information on concepts and reasoning units, and ”indexes” these reasoning unit information according to the concepts, achieving noteworthy reasoning accuracy.
-CRAB[29], based on a Bayesian modeling approach, has established a highly customized ”conservatory” that aligns seamlessly with its proprietary methodology, embodied in the creation of a new RAVEN database.
+Triple-CFN implicitly extracts information on concepts and reasoning units, and “indexes” these reasoning unit information according to the concepts, achieving noteworthy reasoning accuracy.
+CRAB[29], based on a Bayesian modeling approach, has established a highly customized “conservatory” that aligns seamlessly with its proprietary methodology, embodied in the creation of a new RAVEN database.
 This bespoke conservatory, despite forfeiting the inherent core challenges of RAVEN such as solution diversity and uncertainty, has facilitated remarkable breakthroughs for CRAB.
@@
 II-C Distribution distance measurement method
+II-C1 Wasserstein distance
 The formula for the Wasserstein distance[33] can be expressed as:
@@
 In the field of machine learning, the Wasserstein distance is widely used in generative models to evaluate the discrepancy between images generated by the generator and real images.
+II-C2 Sinkhorn distance
 Sinkhorn distance[34] is a metric based on optimal transport, which approximates the Wasserstein distance through the introduction of entropy regularization, leading to increased computational efficiency.
-This distance satisfies the triangle inequality and is symmetric, but it is not a true distance metric since it converges to the Wasserstein distance only when the regularization term tends to zero.
-In each iteration, the Sinkhorn algorithm computes a new transportation plan based on the current one and updates the cost matrix, gradually reducing the cost of the transportation plan and approximating the optimal transport solution.
-Under certain conditions, the Sinkhorn distance satisfies Lipschitz continuity.
-Specifically, if we consider two probability distributions with support sets in a compact metric space satisfying certain geometric conditions, the corresponding Sinkhorn distance can be shown to be Lipschitz continuous.
-This means that there exists a constant L such that for all probability distributions $\mu^{\prime}$ and $\nu^{\prime}$, as well as $\mu^{\prime}$ and $\nu^{\prime}$, the following holds:
-$S(\mu,\nu)\leq L\cdot d(\mu,\nu)+S(\mu^{\prime},\nu^{\prime})$ (2)
-where $S$ is the Sinkhorn distance based on the cost function $c$, and $d$ is an appropriate distance between these distributions (such as the Wasserstein distance).
-The Lipschitz constant $L$ typically depends on the choice of regularization parameters and cost functions.
-In practice, the Lipschitz constant may be difficult to compute directly but can be estimated through numerical experiments.
-It is worth noting that the Lipschitz constant often relates to the choice of cost function $c$.
-For instance, if $c$ is a metric in the metric space (such as Euclidean distance), and the geometric properties of the space (like diameter or curvature) meet certain criteria, determining bounds for the Lipschitz constant becomes easier.
-However, for general cost functions and regularization parameters, determining the Lipschitz constant can be a highly complex task.
-When the cost function for Sinkhorn distance is Euclidean distance, its Lipschitz constant depends on parameters of the Sinkhorn algorithm used and characteristics of the dataset, such as size, shape, and distribution.
-Additionally, Sinkhorn algorithm parameters, like the number of iterations and the weight of the regularization term, influence the continuity properties of the resulting Sinkhorn distance.
-Therefore, to determine the Lipschitz constant for Sinkhorn distance with Euclidean distance as the cost function, specific dataset experiments and analyses are necessary to evaluate the continuity properties of the algorithm under different parameter settings.
-In applications involving Sinkhorn distance, particularly in machine learning and optimization problems, the focus is often on numerical stability and computational efficiency rather than strictly proving Lipschitz continuity or determining the Lipschitz constant.
-If strict theoretical guarantees are indeed required, detailed analyses of specific problem settings may be necessary.
-II-D Spectral normalization
-Spectral normalization[35] is a stabilization technique for deep neural networks, particularly beneficial for generative adversarial networks (GANs).
+II-D Generative models and spectral normalization
+Wasserstein GAN[36] is a deep learning model that uses the Wasserstein distance as its loss function in Generative Adversarial Networks (GANs) to improve training stability and sample quality.
+Spectral normalization[35] is a stabilization technique for deep neural networks, particularly beneficial for GANs.
 It controls the Lipschitz constant by normalizing weight matrices, ensuring outputs change proportionally to inputs.
@@
 Spectral normalization is computationally efficient and offers a direct way to control network complexity, making it a valuable tool for deep learning practitioners.
-II-E generative model
-Wasserstein GAN[36] is a deep learning model that uses the Wasserstein distance as its loss function in Generative Adversarial Networks (GANs) to improve training stability and sample quality.
-By optimizing this distance, it effectively reduces mode collapse and enhances the realism and diversity of generated data.
-DDPM (Denoising Diffusion Probabilistic Model)[9] is a generative model that models the transition from a real image to random Gaussian noise and vice versa to create new images.
-It involves a forward (diffusion) process adding noise and a reverse (denoising) process restoring the original image.
-The aim is to generate images (x’) that are highly similar to real ones (x), emphasizing realism and authenticity.
 III Methodology
@@
 Bongard-logo and RPM problems employ different questioning approaches aimed at evaluating the solver’s proficiency in extracting and learning these abstract concepts.
-This paper presents Lico-Net, a novel baseline model specifically designed for the RPM problem, which achieves remarkable performance in addressing this cognitive challenge.
+This paper presents Lico-Net, a novel baseline specifically designed for the RPM problem, which achieves remarkable performance in addressing this cognitive challenge.
 Subsequently, we introduce a fundamental argument that representing abstract and complex concepts as distributions, rather than vectors, is a more suitable approach.
@@
 IV Discriminator of Case (DC)
-In this section, to enhance our understanding and tackle the RPM (Raven’s Progressive Matrices) problem more effectively, we introduce a new baseline architecture, named Lico-Net.
-This innovative network is designed to address the unique challenges posed by RPM, leveraging advanced techniques and methodologies to achieve superior performance.
-Through a rigorous evaluation, we demonstrate the effectiveness of Lico-Net in solving RPM problems, paving the way for future research in this domain.
-Furthermore, we delve deeper into the previously established baseline network for the Bongard-Logo problem, critically examining its performance and limitations.
+In this section, we introduce Lico-Net, a new baseline architecture designed to solve the RPM problem more effectively.
+This network utilizes advanced techniques for superior performance.
+Our evaluation shows Lico-Net’s effectiveness in RPM, opening new avenues for future research.
+Additionally, we critically assess the existing baseline for the Bongard-Logo problem, exploring its strengths and weaknesses.
 IV-A A novel baseline, Lico-Net, for RPM problem
@@
 If the completed matrix adheres to a specific image progressive pattern that can be inferred from the incomplete matrix (problem statement), it is considered a correct answer selection[39].
-Therefore, a successful RPM discriminator is expected to be able to identify the rationality of the progression pattern within a complete $3\times 3$ progressive matrix.
+Therefore, a successful RPM discriminator is expected to be able to identify the rationality of the progression patterns within a complete $3\times 3$ progressive matrix.
 We require an outstanding baseline, leading us to propose the Lico-Net (link-conception net) as a novel network for RPM solvers.
@@
 IV-A1 Perceptron module of Lico-Net
-Aligned with the current consensus among RPM solvers, which emphasizes the need for multi-scale or multi-viewpoint feature extraction from RPM images[28, 23, 24], Lico-Net incorporates the Vision Transformer (ViT) as a perceptron for feature extraction, while preserving the entire output vector, attention results, as multi-view point features.
-In this paper, we denote the number of viewpoints as $L$, a practice similar to that employed in Savirt[24], RS-Tran[28] and Triple-CFN.
-Lico-Net subsequently processes each viewpoint of extracted feature equally.
+Aligned with RPM solver consensus on the importance of multi-viewpoint feature extraction[28, 23, 24, 32], Lico-Net uses the Vision Transformer (ViT) for feature extraction, keeping the entire output vectors, which are attention results, as multi-viewpoint features.
+We denote the number of viewpoints as L.
+Lico-Net then equally processes each viewpoint of extracted feature.
 The perceptron of Lico-Net is illustrated in figure 4.
@@
 The detailed process is described as follows.
-For the isometric viewpoint of an completed $3\times 3$ progressive matrix, we employ an MLP (Multi-layers perceptron) with a bottleneck structure to extract information from all minimal reasoning units (three images within a row for RAVEN and three images within a row or column for PGM) and obtain unit vectors.
+For the equivalent viewpoint of an completed $3\times 3$ progressive matrix, we employ an MLP (Multi-layers perceptron) with a bottleneck structure to extract information from all minimal reasoning units (three images within a row for RAVEN and three images within a row or column for PGM) and obtain unit vectors.
 Specifically, we concatenate the representations of all images within a minimal reasoning unit in their inherent order and input them into the MLP.
@@
 Figure 5: The extraction process of unit vector
-Due to the design of RPM problems, typically multiple minimal reasoning units must be combined to arrive at the image progression pattern of a RPM matrix.
-For instance, a RAVEN case requires at least two row unites to discern the case’s pattern, while PGM necessitates at least two row unites and two column unites[39].
-It’s worth noting that determining the RPM matrix’s progression pattern doesn’t necessitate all minimal reasoning units; therefore, numerous combination methods of minimal reasoning units exist that are sufficient to identify the pattern.
+Due to the design of RPM problems, typically multiple minimal reasoning units must be combined to arrive at the image progression patterns of a RPM matrix.
+For instance, a RAVEN case requires at least two row units to discern the case’s pattern, while PGM necessitates at least two row unites and two column unites[39].
+It’s worth noting that determining the RPM matrix’s progression patterns don’t necessitate all minimal reasoning units; therefore, numerous combination methods of minimal reasoning units exist that are sufficient to identify the pattern.
 We refer to the combination requiring the least unit vectors as the minimal concept group.
-Given that RPM problems might be constrained by multiple independent and decouple patterns, we embed $N$ optimizable vectors within each minimal concept group (in this paper, we embed two optimizable vectors, with $N$ being a hyperparameter that can be freely set) and input them into a standard Transformer-Encoder as tokens to compute $N$ progression pattern vectors for the progressive matrix in the case.
+Given that RPM problems might be constrained by multiple independent and decouple patterns, we embed $N$ optimizable vectors within each minimal concept group (in this paper, we embed two optimizable vectors, with $N$ being a hyperparameter that can be freely set) and input them into a standard Transformer-Encoder as tokens to compute $N$ kinds of potential progression patterns for the progressive matrix in the case.
 Assuming there are $M$ minimal concept groups, we have thus obtained $N\times M$ progression pattern vectors under one viewpoint.
@@
 Figure 6: The extraction process of pattern vector
-This paper independently calculates the scores of each option in the problem under $N$ distinct pattern vectors.
+This paper independently calculates the scores of each option in the problem under $N$ distinct patterns.
 The $M$ minimal concept groups can be divided into two categories: those involving options and those not involving options.
 In this paper, we select the pattern vectors calculated from the concept groups not involving options as queries, and the pattern vectors involving options as keys and values.
-We then compute the multi-head cross-attention results between them and use a new MLP to score the attention outputs.
+We then compute the multi-head cross-attention results between them (using Transformer-Decoder) and use a new MLP to score the attention outputs.
 Since we intentionally introduced $N$ patterns, we obtain $N$ scores, which are summed to yield the final score for one viewpoint.
@@
 It’s worth noting that we treat each dimension of the unit vector as sample results from a distribution and use covariance matrices as additional loss function term to control the correlation between these distributions, aiming to encourage the unit vectors to contain more information.
-The loss function $\ell_{cov}$ attached to a batch of the minimal reasoning unit vectors $\{U_{ij}|\,i\in[1,b],j\in[1,a]\}$ can be expressed as:
-$M_{cov}$ ($(\{u_{ij}|\,i\in[1,n],j\in[1,m]\})$
-$=\frac{1}{n-1}\sum_{i=1}^{n}\sum_{j=1}^{m}({u}_{ij}-\bar{{u}})({u}_{ij}-\bar{{u}})^{\top},\,$
-$\text{where}\,\,\bar{{u}}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{m}{u_{i}}$ (3)
-$\ell_{cov}$ ($(\{U_{ij}|\,i\in[1,b],j\in[1,a]\})$
-$=\frac{1}{d}\sum M_{cov}(\{U_{ij}|\,i\in[1,b],j\in[1,a]\})^{2}\cdot(1-I)$ (4)
-Where $b$ represents the batchsize for training the Lico-Net, and $a$ denotes the count of minimal reasoning units within a single RPM case, which amounts to three in RAVEN and six in PGM.
-$I$ represents the identify matrix.
-And $d$ represents the dimension of unit vectors $U_{ij}$.
-The entire feedforward process of Lico-Net can be depicted in figure 8.
-Figure 8: Feedforward process of Lico-Net
+This approach bears resemblance to ViCReg[40].
 In summary, after completing the RPM problem matrix with any given option, Lico-Net employs its perception module to encode each image within the problem matrix from multiple viewpoints, resulting in $L$ sets of $3\times 3$ image representations.
@@
 For any given set of representations, Lico-Net combines them to form all possible $M$ minimal concept groups.
-Subsequently, Lico-Net extracts $N$ progressive patterns from each minimal concept group, totaling $N\times M$ progressive patterns.
+Subsequently, Lico-Net extracts $N$ progressive patterns from each minimal concept group, totaling $N\times M$ progressive pattern vector s.
 Finally, Lico-Net utilizes an cross-attention mechanism to measure the consistency of each progressive pattern expressed by the $M$ minimal concept groups and reflects this consistency through scores, resulting in $N$ consistency scores.
@@
 The baseline employs the InfoNCE loss function, which results in the $f_{\theta}(z|\,x)$ being able to solve Bongard-logo to a certain degree.
-The InfoNCE loss function[41] can be mathematically expressed as:
-${\ell_{\text{InfoNCE}}}({z_{pri}},{\tilde{z}_{pri}},\{{z_{au{x_{m}}}}\}_{m=1}^{M})$
-$=-\log\frac{{{e^{({z_{pri}}\cdot{\tilde{z}_{pri}})/t}}}}{{{e^{({z_{pri}}\cdot{\tilde{z}_{pri}})/t}}+\sum\nolimits_{m=1}^{M}{{e^{({z_{pri}}\cdot{z_{au{x_{m}}}})/t}}}}}$ (5)
-Where $z_{pri}$ and $\tilde{z}_{pri}$ represent distinct samples belonging to the representations in primary group, and $z_{au{x_{m}}}$ encompasses all $m$ samples within the auxiliary group.
-$m$ denotes the number of negative samples within a minimum learning unit, while $t$ represents the temperature coefficient, and $t$ is set $10^{-3}$.
-This approach fosters a higher degree of similarity between $z_{pri}$ and $\tilde{z}_{pri}$ compared to the similarity between $z_{pri}$ and $z_{au{x_{m}}}$, thereby emphasizing the distinction between primary and auxiliary representations.
+This approach fosters a higher degree of similarity between $z_{ij}$ and $z_{i\tilde{j}}$ (where $\tilde{j}\neq j$ and $\tilde{j}$, $j$ $\in[1,7]$), compared to the similarity between $z_{ij}$ and $\{z_{ij}|\,j\in[8,13]\}$, thereby emphasizing the distinction between primary and auxiliary representations.
 This process aligns with the underlying logic of Bongard-logo.
-InfoNCE can be construed as a loss function devised for the clustering of vector groups.
 This paper reproduces the baseline experiments in Triple-CFN [32], which were originally conducted on the Bongard-Logo using the ResNet18 model.
-The baseline focuses on learning four concepts in Bongard-Logo — ’FF’, ’BA’, ’NV’, and ’CM’ — separately or in combination [32].
+The baseline focuses on learning four concepts in Bongard-Logo — ‘FF’, ‘BA’, ‘NV’, and ‘CM’ — separately or in combination [32].
 To assess the impact of model size on the baseline experimental findings, we have replaced the backbone of the baseline experiments from ResNet18 to ResNet50 and repeated the experiments.
-The experimental results are presented in Table I, while the feedforward process is illustrated in Figure 9.
+The experimental results are presented in Table I.
 The findings indicate that when presented with a combination of databases, both ResNet18 and ResNet50 experience confusion among the four distinct human concepts.
-The varying degrees of decrease in the models’ reasoning accuracy on the FF and BA problems, coupled with the increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion.
+The varying degrees of decrease in the models’ reasoning accuracy on the FF and BA problems, accompanied by an increase in accuracy on the generalization problems NV and CM, both clearly demonstrate this confusion.
 Furthermore, this confusion cannot be attributed to insufficient network capacity, as even ResNet50, which is more capacious than the ResNet18 used in Triple-CFN’s baseline, is subject to the same challenge.
-Figure 9: Feedforward process baseline
 TABLE I: Reasoning Accuracies of DC based ResNet based ResNet50 on Bongard-logo.
@@
 V Distribution Distance-based Discriminator of Case (D3C)
-Using distributions to describe human concepts in abstract reasoning problems for the deep networks is preferable to using vectors.
-Distributions offer a more comprehensive and nuanced approach to capturing the diversity and uncertainty inherent in human concepts.
-Human concepts are often ambiguous, polysemous, and subject to varying degrees and fluctuations, which can be challenging to represent using static vectors.
-In contrast, distributions provide a dynamic and probabilistic framework that can accommodate this inherent variability.
-By leveraging distributions to represent human concepts, we can furnish deep networks with richer and more flexible information.
-This enhanced representational capacity enables the networks to better understand and address the complexity and uncertainty encountered in abstract reasoning problems.
-For instance, probability distributions can be utilized to depict the range of possibilities and variations associated with a concept, while Gaussian distributions can model the continuity, smoothness and even discreteness of a concept’s underlying structure.
+Using distributions to describe human concepts in abstract reasoning problems for deep neural networks is advantageous compared to the use of vectors.
+This is because distributions present a more comprehensive and nuanced means of capturing the inherent diversity and uncertainty in human concepts.
+Human concepts are frequently characterized by ambiguity, polysemy, intersectionality, and overlap, making their representation via static vectors challenging.
+Conversely, distributions afford a dynamic and probabilistic framework that can effectively accommodate such inherent variability.
+Employing distributions to represent human concepts endows deep neural networks with richer and more flexible information.
+This, in turn, enhances their capability to comprehend and resolve complex, uncertain abstract reasoning problems.
+Probability distributions illustrate the possibilities and variations associated with a concept, whereas Gaussian distributions specifically model the underlying structure of a concept.
 V-A D3C for Bongard-Logo
-D3C proposes that distributions are more suitable as abstract representations of concepts compared to vectors.
-Therefore, D3C decides to estimate the concept in each image of a Bongard-Logo case using a deep network in the form of $f(z|\,x)$ and represent the concept with a distribution.
-Finally, the distances between these distributions are measured and used as the distances between concepts.
-These distances can be naturally applied to solve the Bongard-Logo problem.
-D3C is a plug-and-play method targeted at solving Bongard-Logo and similar problems for any network in the form of $f(z|\,x)$.
-Notably, this brings improvements for models such as Triple-CFN and ResNet50.
-In practice, this paper hypothesizes that the concepts of the image samples in a Bongard-Logo case follows a Gaussian distribution.
-We utilize an arbitrary backbone network $f(z|\,x)$ to excract the concepts of images within Bongard-Logo cases in form of approximating distribution.
-To achieve this, we modify the network to output two vectors $\mu$ and $\sigma$ instead of a single vector $z$, effectively converting $f(z|\,x)$ into $f(\mu,\sigma|\,x)$.
-The encoded vector $z$ is then divided into two equal-length segments $\mu$ and $\log(\sigma^{2})$, which we interpret as the mean and the logarithm of the variance, respectively.
-This approach normalizes the output vectors into Gaussian distributions and represents concepts as distributions in Bongard-Logo cases.
-Our method bears resemblance to the Variational Autoencoder (VAE) framework[8].
-We argue that using distributions, rather than vectors, to describe image representations offers greater flexibility and accuracy in capturing the underlying complexity and uncertainty of abstract reasoning problems.
-With the estimated concepts for each image in the Bongard-Logo cases as distributions, we compute the Sinkhorn distance between these distributions using reparameterization technique[8].
-Following the logical structure of each Bongard-Logo case, we impose constraints on these distances for the backbone network $f(\mu,\sigma|\,x)$, aiming to train an effective discriminator.
-Specifically, we optimize the network by utilizing a binary cross-entropy loss function to enforce constraints on the exponential function of the negative Sinkhorn distance with base $e$ ($e^{-D}\in[0,1]$, where $D\in[0,+\infty]$ is the values of Sinkhorn distance), aligning with the case’s underlying logic mentioned in DC.
-This process can be expressed as figure 10.
-The D3C Method, applied to a batch of Bongard-Logo case distributions encoded by the backbone network $f(\mu,\sigma|\,x)$, can be expressed as follows:
+D3C argues that distributions are better abstract concepts than vectors.
+Therefore, it employs a deep network $f(z|\,x)$ to estimate concepts in Bongard-Logo images and represents them as distributions.
+By measuring concept distances as the distances between these distributions, the method becomes applicable to solving the Bongard-Logo problem.
+D3C is a plug-and-play method for any network with form of $f(z|\,x)$, improving models like Triple-CFN and ResNet50.
+In practice, this paper assumes the concepts in Bongard-Logo cases follow a Gaussian distribution.
+We use a backbone network $f(z|\,x)$ to extract these concepts as distributions.
+To do this, we adapt the network to output $\mu$ and $\sigma$ instead of a single vector $z$, turning $f(z|\,x)$ into $f(\mu,\sigma|\,x)$.
+The vector $z$ is split into two equal-length segments $\mu$ and $\log(\sigma^{2})$, representing mean and variance logarithm, respectively.
+This method normalizes original outputs $z$ to Gaussian distributions $\{\mu,\sigma^{2}\}$ and bears resemblance to the Variational Autoencoder (VAE) framework[8].
+Representing concepts as distributions, not vectors, provides more flexibility and accuracy in abstract reasoning.
+With the estimated concepts for each image in the Bongard-Logo dataset represented as distributions, we compute the Sinkhorn distance between these distributions using the reparameterization technique[8].
+Subsequently, we impose constraints on these distances to train and optimize the backbone network $f(z|\,x)$, aligning with the underlying logic of the cases.
+The optimization process involves minimizing a binary cross-entropy loss, which is calculated using the exponential of the negative Sinkhorn distance as a factor.
+The entirety of this process is illustrated in Figure 8.
+When applied to a batch of Bongard-Logo cases, the D3C method can be summarized as follows:
 ${\ell_{\text{D3C}}}(\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\})$
-$=\sum_{i=1}^{b}\sum_{j=1}^{7}\sum_{\tilde{j}=j+1}^{7}{\text{Sinkhorn}({z_{ijs}},\,{{z}_{i\tilde{j}s}})}$
-$+\sum_{i=1}^{b}\sum_{{j}=1,}^{7}\sum_{\tilde{j}=7}^{14}\log({1-{e^{-\text{Sinkhorn}({z_{ijs}},\,{{z}_{i\tilde{j}s}})}}})$ (6)
-Where $b$ is the batchsize of training.
-The $\{z_{ijs}|\,i\in[1,b],j\in[1,14],s\in[1,S]\}$ represents $S$ samples drawn from the distribution $\{\mu_{ij},\sigma_{ij}|\,i\in[1,b],j\in[1,14]\}$ using the reparameterization trick, where the distribution $\{\mu_{ij},\sigma_{ij}|\,i\in[1,b],j\in[1,14]\}$ is obtained by mapping the Bongard-Logo image $\{x_{ij}|\,i\in[1,b],j\in[1,14]\}$ through the network $f(\mu,\sigma|\,x)$.
-The network $f(\mu,\sigma|x)$ is derived from the network $f(z|x)$ by D3C Method.
-$S$ is set to 20, and $S$ represents the sampling frequency of the reparameterization technique.
+$=\sum_{i=1}^{b}\sum_{j=1}^{7}\sum_{\tilde{j}=j+1}^{7}{\text{S}(\{z_{ijs}|\,s\in[1,S]\},\,\{z_{i\tilde{j}s}|\,s\in[1,S]\})}$
+$+\sum_{i=1}^{b}\sum_{{j}=1,}^{7}\sum_{\tilde{j}=7}^{14}\log({1-{e^{-\text{S}(\{z_{ijs}|\,s\in[1,S]\},\,\{z_{i\tilde{j}s}|\,s\in[1,S]\})}}})$ (2)
+[…]
 Figure 8: Feedforward process of D3C
-The design of D3C allows for its application to any image encoder that can be abstracted as $f(z|x)$.
-This requirement is met by a wide range of encoders, including those from the ResNet family, ViT, and specialized ones like Triple-CFN, which is particularly suitable for Bongard-Logo problems.
+The design of D3C allows it to be applied to any image encoder that can be abstracted as $f(z|x)$.
+A wide range of encoders, including those from the ResNet family, ViT, and specialized encoders like Triple-CFN—which is particularly suitable for Bongard-Logo problems—meet the abstraction requirement and are thus compatible with D3C.
 V-B D3C on Lico-Net for RPM
-Notably, in contrast to the Bongard-Logo database, both RAVEN and PGM possess well-defined and comprehensive Meta data in the form of (auxiliary) rule labels.
-These rule descriptions facilitate the construction of a rational concept space, with boundary concepts within the rule labels.
-This paper introduces the rule labels from the RPM database into the Lico-Net, aiming to enhance the reasoning accuracy and human-perspective rule interpretability of Lico-Net.
+Notably, in contrast to the Bongard-Logo database, each case in RAVEN and PGM is accompanied by well-defined and comprehensive rule labels in the form of Meta data.
+These auxiliary labels provide descriptions of the image progression patterns within the cases, facilitating the delineation of concept boundaries.
+Therefore, this paper decides to introduce the Meta data from the RPM database into the Lico-Net, aiming to organize the distribution of the encoded progressive pattern vector $\{P_{nm}\}$ in Lico-Net.
+This enhancement is expected to improve both the reasoning accuracy and human-perspective interpretability of Lico-Net.
 This process will be implemented by the D3C method.
-Within the frameworks of RAVEN and PGM, there is a many-to-one mapping between image progression patterns and their corresponding rules.
-For instance, both single-step and double-step increments, as well as single-step and double-step decays in image attributes, are all categorized under the ‘progressive’ rule.
-Therefore, this paper proposes the concept of rules from the RPM database and represent them as distributions while also vectorizing the image progression patterns.
-This approach is precisely the original idea behind D3C, which posits that describing concepts using distributions is more appropriate than using vectors.
-We can leverage this approach to naturally constrain the image progression patterns within their corresponding rule distributions, thereby optimizing the deep model’s learning of rules and clarifying the boundaries among various concepts within the data.
-In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual descriptions of Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data.
-Specifically, the initial portion of the feature vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the rule descriptions and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$.
+Within the frameworks of RAVEN and PGM, there exists a many-to-one mapping between image progression patterns and their corresponding concepts.
+For instance, both single-step and double-step increments, as well as single-step and double-step decrements in image attributes, are categorized under the umbrella of the “progressive” concept according to Meta data.
+Therefore, this paper proposes extracting the concepts described by the Meta data and representing them as distributions, which aligns with the original idea behind D3C.
+D3C posits that using distributions to describe concepts is more suitable than using vectors.
+As Lico-Net has already extracted the progression patterns of images for the RPM matrix and encoded them as vectors, we can naturally employ concepts with distribution forms to constrain and delimit these image progression pattern vectors, ensuring that these progression pattern vectors belong to the distribution of corresponding concept.
+This approach will optimize Lico-Net’s learning of rules and clarify the boundaries among various concepts.
+The aforementioned idea can be represented in the framework illustrated in the Figure 9.
+Figure 9: The framework of D3C
+In practice for D3C, we utilized a standard Transformer-Encoder to convert the textual Meta data into concept vectors $\{{r}_{k}|\,k\in[1,K]\}$, where $K$ represents the count of distinct rule description in Meta data.
+Specifically, the initial portion of the concept vector is designated as the mean $\{\mu_{k}|\,k\in[1,K]\}$, whereas the latter half signifies the logarithm of variance $\{\log({\sigma_{k}}^{2})|\,k\in[1,K]\}$, effectively distributing the concepts described in Meta data and yielding a set of distributions $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$.
 Lico-Net has inherently achieved the vectorization of image progression patterns, yielding vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ termed as progression pattern vectors.
-Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns should exhibit a high likelihood of adhering to the distribution outlined by their corresponding textual rules, and conversely, demonstrate a low probability of deviation.
-A specific RPM case is encoded into $N\times M$ progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ from $L$ viewpoints by Lico-Net as aforementioned.
-To align Lico-Net with the D3C framework, the number of encoding patterns $N$ in Lico-Net is set according to the instance of the RPM problem, depending on the number of decoupled concepts present in the Meta data of that instance.
+Lico-Net extracts N types of progression patterns (where N can be freely set) for each complete RPM matrix, and encodes M progression pattern vectors (where M is determined by the number of existing minimal concept groups) for each progression pattern, forming the vector group $\{P_{nm}\}$.
+This process can be reviewed in Figure 6.
+To integrate Lico-Net with the D3C framework, the number of encoding patterns $N$ in Lico-Net, which can no longer be set arbitrarily, is set according to the instance of the RPM problem, depending on the number of decoupled concepts present in the Meta data of that instance.
 For example, in PGM, there are two such concepts: shape pattern and line pattern, and $N$ is set to 2.
 The special setting of $N$ in Lico-Net can be viewed as $N$ slots, which can be aligned with the $N$ decoupled concepts described in the Meta data, respectively.
+Consequently, we integrated a constraint term into Lico-Net, dictating that vectors representing progression patterns $\{P_{nm}\}$ should exhibit a high likelihood of adhering to the distribution outlined by their corresponding concepts, and conversely, demonstrate a low probability of deviation.
+To impose constraints on the distributions of the extracted $N$ progression patterns, we have devised the following process.
 By synthesizing all perspectives—specifically, by calculating the mean of the progression pattern vectors $\{P_{nm}|\,n\in[1,N],m\in[1,M]\}$ across all $L$ viewpoints—Lico-Net encodes $N\times M$ vectors, denoted as $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$, representing the full-viewpoint progression patterns.
-This process can be expressed as:
-$\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$
-$=\frac{1}{L}\sum_{l=1}^{L}\{P_{nm}|\,n\in[1,N],m\in[1,M]\}_{l}$ (7)
-The $l$ represents the index of viewpoints.
-The $\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\}$ can be treated as $N$ sets of $M$ sampling results.
-This enables the optimization of distribution distances based on the Sinkhorn distance, thereby enhancing the overall performance through D3C framework.
-Apparently, this paper utilizes the Sinkhorn distance to constrain the distance between $M$ progression pattern vectors and their corresponding distributions’ sampling results.
-The $N$ decoupling concept presented in Meta data is employed to concurrently impose constraints on the $N$ sets of $M$ progression pattern vectors, and the process of aligning $\{\overline{P}_{nm}\}$ with the distributional concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ can be illustrated in figure 11.
+The set $\{\overline{P}_{nm}\}$ can be interpreted as the outcome of sampling $M$ times from each of the $N$ distributions of progression patterns across all viewpoints generated by Lico-Net.
+This interpreting perspective enables us to utilize Meta data to enforce constraints on the $N$ distributions.
+Next, we sample $S$ times from the distribution of concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ to obtain sampling results $\{r_{ks}|\,k\in[1,K],s\in[1,S]\}$.
+Consequently, this paper employs the Sinkhorn distance to constrain the distance between the distributions of the $N$ kinds of progression patterns encoded by Lico-Net and their corresponding concept distributions in the Meta data.
+The process of aligning $\{\overline{P}_{nm}\}$ with the distribution of concepts $\{{\mu}_{k},{\sigma^{2}}_{k}|\,k\in[1,K]\}$ can be illustrated in figure 10.
 Figure 10: Feedforward process of D3C
@@
 ${\ell_{\text{D3C}}}(\{\overline{P}_{nm}|\,n\in[1,N],m\in[1,M]\},\{r_{ks}|\,k\in[1,K],s\in[1,S]\})$
-$=\sum_{n=1,\tilde{k}|\,n}^{N}{\text{Sinkhorn}({\overline{P}_{nm}},\,{r_{\tilde{k}s}})}$
-$+\sum_{k=1,k\neq\tilde{k}}^{K}\sum_{n=1,\tilde{k}|\,n}^{N}\log({1-{e^{-\text{Sinkhorn}({\overline{P}_{nm}},\,{r_{{k}s}})}}})$ (8)
-The $\overline{P}_{nm}$ represents the full-viewpoint progressive pattern vectors encoded by Lico-Net, and $\{r_{ks}\}$ represents the $S$ sampling results from the distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of the rule description.
-It is evident that the specific value of $\tilde{k}$ is contingent upon the description of Meta data corresponding to the particular case of the RPM problem.
-symbol “$\tilde{k}|\,n$” represents that the specific value of $\tilde{k}$ is depended on $n$.
-This dependency realizes the fact that each slots in Lico-Net are aligned with their corresponding concepts.
-The $\text{Sinkhorn}(\cdot,\cdot)$ represents the Sinkhorn distance operator.
-Other logic of symbol in function (V-B) has mentioned.
-It becomes evident that we are employing the Binary Cross Entropy (BCE) loss function to impose constraints on the exponential negative Sinkhorn distance result.
-The feedforward process is illustrated in the figure 12.
-Figure 12: Feedforward process of Lico-Net aligned with D3C
+$=\sum_{n=1,\tilde{k}|\,n}^{N}{\text{S}(\{\overline{P}_{nm}|\,m\in[1,M]\},\,\{r_{\tilde{k}s}|\,s\in[1,S]\})}$
+$+\sum_{k=1,k\neq\tilde{k},\tilde{k}|\,n}^{K}\sum_{n=1}^{N}\log({1-{e^{-\text{S}((\{\overline{P}_{nm}|\,m\in[1,M]\},\,\{r_{{k}s}|\,s\in[1,S]\}))}}})$ (3)
+[…]
+In summary, Lico-Net is capable of extracting $N$ types of progression patterns for each minimal concept group in the RPM problem.
+Given the existence of $M$ minimal concept groups, we obtain a total of $N\times M$ progression pattern vectors.
+To constrain and clarify the distribution of vectors under each of the $N$ progression patterns, we consider the vectors as the outcome of $M$ samplings from each of the $N$ distributions.
+This perspective enables us to implement distributional constraints based on concepts in Meta data, thereby enhancing the alignment and interpretability of the progression patterns extracted by Lico-Net.
 V-C A low-cost loss function for the D3C, the Hermes for the Zeus
 Sampling $S$ times (where $S$ is arbitrarily set, and $S$ is set to $20$ for Sinkhorn distance) from the constructed distribution $\{\mu_{k},{\sigma^{2}}_{k}\}$ of concepts in Meta data and yielding $\{r_{ks}|\,k\in[1,K],s\in[1,S]\}$, we calculate the Sinkhorn distance between the progression pattern vectors and these sampling results $\{r_{ks}\}$.
-This Sinkhorn distance approximates the Wasserstein distance, which is currently a linear programming algorithm also known as the Earth Mover’s Distance (EMD) algorithm.
 Typically, this approach requires a substantial number of samples and incurs a considerable computational cost.
-Even if we treat the $M$ progression pattern vectors as $M$ sampling results, it is still difficult to satisfy the prerequisite of extensive sampling in the context of the RPM problem, where $M$ is set to 9 in PGM and 3 in RAVEN.
-To reduce computational cost and sampling frequency, this paper employs cosine similarity between vectors as an alternative distance metric norm to the Sinkhorn distance.
-Obviously, the cosine similarity satisfacte with the $Lipschitz$ continuity condition.
-For the inner product, treating it as a function with one fixed vector, we observe that it exhibits $Lipschitz$ continuity.
-Specifically, given a fixed vector $v$, the function $f(u)=u\cdot v$ is $Lipschitz$ continuous with respect to $u$, with the $Lipschitz$ constant $L$ being equivalent to the norm of $v$.
-This property is attributed to the inner product’s adherence to the Cauchy-Schwarz inequality: $|u\cdot v|\leq||u||\cdot||v||$.
-Leveraging this inequality, we are able to deduce the underlying inequality:
-$f(u_{1})-f(u_{2})=|u_{1}\cdot v-u_{2}\cdot v|=|(u_{1}-u_{2})\cdot v|\leq||u_{1}-u_{2}||\cdot||v||$ (9)
-In our study, we opted for cosine similarity, considering normalized vectors $u$ and $v$, where $||v||$ and $||u||$ are equal to 1, and Lipschitz constant $L=||v||=1$.
-It is reasonable to choose the mean of cosine similarity between sampling vectors as the distance metric function instead of Sinkhorn distance.
+To reduce computational cost and sampling frequency, this paper employs the mean of cosine similarity between sampling vectors as an alternative distance metric norm to the Sinkhorn distance.
 The calculate process of new method applied on Bongard-Logo can be expressed as:
@@
 It is evident that, compared to function (V-A) and (V-B), the demand for sampling frequency $S$ in D3C-cos has been significantly reduced.
-From a certain perspective, estimating the distribution distance based on cosine similarity between sampling results can be regarded as a variant of the Wasserstein distance based on cosine similarity that relinquishes the lower bound estimation and possesses a relatively low, fixed Lipschitz constant.
-Therefore, this method, we called “D3C-cos”, cannot precisely quantify the distance between distributions.
-It serves as a low-cost loss function for constrained inter-distribution distance, serving as a surrogate for the Sinkhorn distance, the Hermes for the Zeus.
+Estimating distribution distance via cosine similarity between samples can be seen as a modified Wasserstein distance, relinquishing lower bound estimation and having a fixed Lipschitz constant.
+This ”D3C-cos” method, while not exact, provides a cost-effective loss function for distribution distance, acting as a substitute for the Sinkhorn distance.
+We still use Sinkhorn distance for performance testing.
 The success of such simplification can be attributed, in part, to our modeling of the distribution of pattern descriptions as a Gaussian distribution.
@@
 In the Bongard-Logo problem, the determination of image attributes is influenced by the specific contextual concepts associated with each individual case [32].
-Distinct concepts require the network to encode specific pixel configurations into differentiated attribute representations, potentially compelling the network to maintain mutually exclusive representation encoding strategies and methods [32].
+Distinct concepts require the network to encode specific pixel configurations into different attribute representations, potentially compelling the network to maintain mutually exclusive representation encoding strategies and methods [32].
 This scenario often leads to concept confusion within the network, rendering it uncertain about which concept to rely on when processing a Bongard-Logo case [32].
-In a nutshell, from the perspective of network, there exists mutual infringement and erroneous coupling among the high-dimensional human concepts in the Bongard-Logo.
+In a nutshell, from the perspective of network, there exists mutual infringement and erroneous coupling among the high-dimensional human concepts in the Bongard-Logo cases.
 Therefore, this study aims to develop a methodology that effectively decouples the concepts across different cases within the Bongard-Logo problem.
 We posit that by successfully decoupling such that distinct cases rely on distinct concepts we may alleviate, to a certain extent, the issue of the network losing concept matching when confronted with a Bongard-Logo problem.
-To achieve this objective, we have redefined the concepts within the Bongard-Logo database, representing them as conditional probability distributions $p_{i}(x|\,Y_{i})$ and $q_{i}(x|\,Y_{i})$.
-Here, $i$ ranges from $1$ to $n$, denoting each index of the Bongard-Logo cases.
-Additionally, we introduce the notation $\{Y^{\prime}_{i}|\,i\in[1,n]\}$ to represent the set of reorganized concepts.
-It is evident that the number of reorganized concepts corresponds to the number of Bongard-Logo cases.
-Through this concept reorganization, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases.
-In other words, each case is categorized and solved based on its distinct concepts.
-By establishing a one-to-one correspondence between problem cases and these redefined concepts, we aim to enable the network to avoid uncertainty in determining which concept to rely on when solving a specific case.
+Therefore, we strive for mutually exclusive expressions of concepts across all Bongard-Logo cases.
+By establishing a one-to-one correspondence between problem cases and concepts, we aim to enable the network to avoid uncertainty in determining which concept to rely on when solving a specific case.
 This approach aims to mitigate the impact of concept confusion on the solution for the Bongard-Logo problem.
@@
 VI-A1 D2C phase 1: The concept delineation method based on a novel dynamic mini-batch sampling technique
-By incorporating an additional ancestral sampling procedure into the traditional mini-batch training sampling methods, this study constructs a Bongard-logo problem with reorganized concept space for training purposes.
-This constitutes a hot-swappable approach for training the Bongard-Logo solver, which is designated as phase 1 of the D2C method in this paper.
-Specifically, we denote a batch of Bongard-logo cases as $\{x_{ij}|\,i\in[1,b],j\in[1,14]\}$, where $i$ denotes the index of the Bongard-logo case within a batch, the batch size is $b$, and $j$ represents the sample index within a single Bongard-logo case, with $j\in[1,7]$ being the primary samples and $j\in[8,14]$ being the auxiliary samples.
-From this batch of cases $\{x_{ij}\}$, we deliberately choose two batches of cases $\{x_{tj}|\,j\in[1,14]\}$ and $\{x_{\tilde{t}j}|\,j\in[1,14]\}$, ensuring that $t$, $\tilde{t}$, and $i$ are all distinct.
-This process can be viewed as ancestral sampling.
-The extracted primary samples $\{x_{tj}|\,j\in[1,7]\}$ and $\{x_{\tilde{t}j}|\,j\in[1,7]\}$ are then incorporated into $\{x_{ij}\}$ as auxiliary samples, yielding an expanded set $\{x_{i\tilde{j}}\}$ where $\tilde{j}\in[1,28]$.
-Subsequently, The new sample groups $\{x_{i\tilde{j}}\}$, which oppose the concepts from different cases compared to their original form, can be fed into any discriminator used to solve Bongard-Logo problems, such as Triple-CFN[32] and PMoC[31].
-The feedforward process, as illustrated in the accompanying diagram, involves a strategic misalignment of Bongard-logo problem batches to facilitate ancestral sampling during implementation.
-The alignment of the D2C phase 1 process with the baseline in DC is illustrated in figure 13.
+The design idea of D2C phase 1 is quite simple, which involves introducing primary group samples from other cases as additional samples into the auxiliary group of the current case.
+During the implementation process, for each case under consideration, we introduce two distinct cases into it.
+This process can be regarded as ancestral sampling.
+This constitutes a plug-and-play approach for training most Bongard-Logo solver such as PMoC[31] and Triple-CFN[32], which is designated as phase 1 of the D2C method in this paper.
+The process of D2C phase 1 is in Figure 11.
+As evident from the figure, the ancestral sampling process can be accomplished by performing a misalignment stacking of the sampled Bongard-Logo cases.
+By incorporating this additional ancestral sampling procedure into the traditional mini-batch training sampling methods, this study constructs a Bongard-logo problem with reorganized concept space for training purposes.
+This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts.
+This paper aims to utilize this strategy to mitigate concept infringement and erroneous coupling among different Bongard-Logo cases.
 Figure 11: Feedforward process of D2C phase 1
-It is worth emphasizing that the improvement in model performance observed through the aforementioned approach is not solely attributed to an increase in the number of negative examples within each contrastive group.
-This paper posits that the incorporation of negative examples, which assist in delineating concept boundaries, holds the key to enhancing a model’s proficiency in abstract reasoning tasks.
-Merely augmenting the dataset with insignificant negative examples does not lead to substantial gains in model performance.
-In an attempt to enrich the dataset, this study generated a substantial amount of Bongard-logo virtual data using DDPM and integrated these as auxiliary data points within the original problems.
-However, subsequent analysis revealed that this methodology did not tangibly enhance model accuracy.
-In summary, the phase 1 of the D2C approach embodies a straightforward yet effective concept: it introduces positive examples from other cases as negative samples in the current case.
-This strategy sends a clear and additional signal to Bongard-Logo discriminators that distinct problem cases should rely on differing concepts.
-As a result, the first phase of D2C is widely applicable to various Bongard-Logo discriminators, including Triple-CFN and PMoC.
+It is worth emphasizing that the improvement in model performance provided by D2C is not solely attributed to an increase in the number of negative examples within each contrastive group.
+This paper posits that the incorporation of negative examples, which assist in delineating concept boundaries among cases, holds the key to enhancing a model’s proficiency in abstract reasoning tasks.
+Simply adding unhelpful or irrelevant negative examples to the dataset does not significantly improve model performance.
+In an attempt to enrich the dataset, this study generated a substantial amount of Bongard-logo image data using DDPM [9] and integrated these as auxiliary group sample within the original problems.
+However, this methodology based on DDPM [9] did not significantly enhance model accuracy.
 VI-A2 D2C phase 2: The concept delineation method based on a new loss function term
-This paper posits that, in addition to altering the manner of sampling training data, establishing an extra loss function to constrain the similarity between the representations of the primary group samples across different cases within a batch of Bongard-Logo training data can implicitly achieve the fundamental objective of this paper.
-Specifically, in this paper, we exclusively retain the representation of primary group within the set $\{z_{ij}|\,i\in[1,b],j\in[1,14]\}$, which specifically correspond to $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$.
-Subsequently, we randomly partition the primary representation of each case within $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ into two distinct subsets: $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, where $\{z_{ih}|\,i\in[1,b]\}\cup\{z_{i\tilde{h}}|\,i\in[1,b]\}=\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ and $\{z_{ih}\}\cap\{{z}_{i\tilde{h}}\}=\emptyset$.
+This paper proposes that, besides changing training data sampling, adding a loss function to limit the similarity of primary group samples’ representations across cases in a Bongard-Logo training batch implicitly fulfills the D2C’s main target.
+Specifically, in this paper, we exclusively retain the representation of primary group within the set $\{z_{ij}|\,i\in[1,b],j\in[1,14]\}$, which is $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$.
+Subsequently, we randomly partition the primary representation of each case within $\{z_{ij}|\,i\in[1,b],j\in[1,7]\}$ into two distinct subsets: $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$.
 By taking the mean of the vectors within $\{z_{ih}|\,i\in[1,b]\}$ and $\{z_{i\tilde{h}}|\,i\in[1,b]\}$, we derive the corresponding vectors $\{v_{i}|\,i\in[1,b]\}$ and $\{\tilde{v}_{i}|\,i\in[1,b]\}$.
@@
 ${\ell_{\text{D2C}}}(\{z_{ij}|\,\,i\in[1,b],j\in[1,7]\})$
-$=-\sum^{b}_{i=1}\log\frac{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}}}{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}+\sum\nolimits_{\tilde{i}=1,\tilde{i}\neq i}^{b}{{e^{({v_{i}}\cdot{v_{\tilde{i}}})/t}}}}}$
+$=-\sum^{b}_{i=1}\log\frac{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}}}{{{e^{(v_{i}\cdot\tilde{v}_{i})/t}}+\sum\nolimits_{\tilde{i}=1,\tilde{i}\neq i}^{b}{{e^{({v_{i}}\cdot{v_{\tilde{i}}})/t}}}}}$ (6)
 $\text{where}\,\,v_{i}=\sum\{z_{ih}\},\,\tilde{v}_{i}=\sum\{z_{i\tilde{h}}\}\,\,$
-$\text{and}\,\,\{z_{ih}\}\cup\{z_{i\tilde{h}}\}=\{z_{ij}|\,j\in[1,7]\}$ (12)
-This loss term is to decrease the similarity of primary group representations across different cases. The schematic representation of the feedforward procedure is provided in the figure 14.
-Figure 14: Feedforward process of D2C phase 2
-In essence, the phase 2 of D2C attempts to achieve a similar goal as its first phase by formulating a new loss function, approaching the task from a different angle.
-Nevertheless, this strategy renders the phase 2 of D2C primarily compatible with backbone that can be abstracted as $f(z|x)$.
-This adaptability enables a majority of image encoders, such as ResNet, ViT, and Triple-CFN[32], to be compatible with the complete D2C methodology.
-However, the application of PMoC[31], which relies on probability-based measures to assess image identity, is restricted to the first phase of D2C.
+$\text{and}\,\,\{z_{ih}\}\cup\{z_{i\tilde{h}}\}=\{z_{ij}|\,j\in[1,7]\}$
+This loss term is to decrease the similarity of primary group representations across different cases.
+Evidently, D2C employs a ”rather be safe than sorry” approach, which may lead to some overcorrection.
+Essentially, D2C’s phase 2 aims to achieve a similar goal as the phase 1 by introducing a new loss function.
+However, this makes phase 2 mainly compatible with backbones abstractable as $f(z|x)$.
+This allows most image encoders like ResNet and Triple-CFN [32] to be compatible with full D2C.
+But PMoC[31], relying on probability measures, is limited to D2C’s phase 1.
 VI-B D2C on RPM probelm
-Furthermore, the spirit of D2C method can be integrated with Lico-Net by introducing images from other case statements as negative examples within the cases.
-However, the application of D2C in RPM discriminators is limited to its phase 1, as it becomes challenging to forcibly apply the phase 2 to these discriminators.
-This process is illustrated in figure 15.
+Lico-Net can achieve the integration with D2C by introducing images from other RPM cases as incorrect options within the option pool of the current RPM case.
+It’s worth noting that the application of D2C in RPM discriminators is confined to phase 1 due to the complexity arising in the implementation of phase 2.
+This process is shown in Figure 12.
 Figure 12: Feedforward process of Lico-Net aligned with D2C
-In summary, D2C is a straightforward approach.
-In the context of Bongard-Logo, where the determination of image attributes heavily relies on the concepts outlined by the primary and auxiliary groups within each case, D2C addresses this challenge by decoupling the concepts across different cases.
-Nevertheless, regarding the RPM problem, which emphasizes consistency in image attribute identification, this study maintains a pessimistic outlook on the efficacy of D2C in tackling RPM.
-In a nutshell, D2C is suitable for addressing the Bongard-Logo problem, where there exists mutual infringement and erroneous coupling of concepts across cases, but it is challenging to apply D2C to the RPM problem, which involves sharing a common set of concepts across cases.
+VI-C The summary of D2C
 It is evident that D2C employs a precautionary approach, aiming to minimize mutual infringement and erroneous coupling among concepts, even at the cost of potential overcorrection.
-Finally, the D2C and D3C methodologies are not mutually exclusive.
-Specifically, D2C endeavors to redefine concept boundaries for reasoning problems.
-According to D3C, representing concepts in the form of distributions is a more plausible approach compared to utilizing vectors.
+However, we posit that D2C is suitable for addressing the Bongard-Logo problem, where there exists mutual infringement and erroneous coupling of concepts across cases, but it is challenging to apply D2C to the RPM problem, which involves sharing a common set of concepts across cases and focuses on consistent attribute identification.
 VII Defense-Driven Discrimination of Distributed Concepts (D4C)
-This paper aims to delineate more appropriate concept boundaries for the RPM and Bongard-Logo problems.
+The performances of D2C and D3C on both Bongard-Logo and RPM tasks have stimulated deeper reflections on abstract reasoning challenges.
+This raises some fascinating questions: Why does modifying negative examples have a profound impact on the solver’s performance specifically in the Bongard-Logo problem?
+Beyond posing difficulties for the discriminator, what are the additional roles played by positive and negative examples in abstract reasoning challenges?
+After thoughtful consideration of these aspects, this section introduces the D4C approach.
 VII-A Negative samples are not merely intended to confound the discriminator
 This paper posits that it is worthwhile to consider the significance of positive and negative examples in abstract reasoning problems.
-Suppose that the correct solutions to each case in a reasoning problem follow a Gaussian distribution.
-Given that some reasoning problems may have multiple correct solutions, it is also possible that the distribution is a mixed Gaussian distribution.
-This is exemplified in datasets such as OIG and D9 from Raven database.
-One of the core challenges in RPM problems lies in the existence of multiple correct answers.
-While CRAB has eliminated this core difficulty through data cleaning in the RAVEN dataset, it has achieved noteworthy results on the cleaned version.
-Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern where positive examples yield high probabilities when substituted into the case’s problem statement, while negative examples yield low probabilities.
-In a nutshell, they are models with probabilistic labeling scheme.
+Currently, existing discriminative models for solving RPM and Bongard-Logo problems often adhere to a pattern known as the probability labeling scheme.
+As the name suggests, when a positive example is submitted to the reasoning problem’s statement, the discriminator will respond with a high probability, whereas the negative examples will be labeled with a low one.
 We can view this pattern in the following way: the discriminator constructs a solution distribution for the cases presented in abstract reasoning problems.
@@
 The endeavor to train the discriminator can be conceptualized as a process of optimizing and fine-tuning the solution distribution, with the ultimate objective of aligning it more closely with the underlying distribution of correct solutions for the given cases.
-Using positive and negative examples to estimate a Gaussian distribution through the cross-entropy loss function can be seen as leveraging them to optimize the following function:
+To further explain our approach in a clearer and intuitive manner, we assume that the distribution of correct solutions for each case in a reasoning problem follow a Gaussian distribution.
+Given that some reasoning problems may have multiple correct solutions, it is also possible that the distribution is a mixed Gaussian distribution.
+Using positive and negative examples to estimate the parameters of a Gaussian distribution through the cross-entropy loss function can be seen as leveraging them to optimize the following function:
 $-\log(p(x,\mu,\sigma))=\frac{{(x_{i}-\mu_{i})^{2}}}{{2\sigma^{2}_{i}}}-\log(|\,\sigma_{i}|\,)-\frac{1}{2}\log(2\pi)$ (7)
 Where $x$ represents the positive or negative samples, $\mu$ and $\sigma$ represent the mean and standard deviation of the Gaussian distribution separately.
-It can be observed that when positive samples are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the mean squared loss between sample $x$ and the mean of the correct solution $\mu$.
-Conversely, when negative samples are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$ of the respective segment.
-Therefore, this paper posits that in reasoning problems, the establishment of positive examples aids deep models in estimating the mean of the correct solution distribution, while negative examples serve to estimate the variance.
+It can be observed that when positive samples are substituted into the formula, we aim for the smallest possible result, which corresponds to minimizing the Euclidean distance between sample $x$ and the mean of the correct solution $\mu$.
+Conversely, when negative samples are substituted, we desire the largest possible result, equivalent to minimizing the standard deviation $\sigma$.
+Therefore, this paper posits that in reasoning problems, the establishment of positive examples aids deep models in estimating the mean of the correct solution distribution, while negative examples serve to limit the variance.
 The variance represents the distance between the boundaries of the distribution and its center.
-Correspondingly, we posit that negative examples play a crucial role in abstract reasoning tasks, serving to delimit the concepts presented in each case.
+Correspondingly, we posit that both positive and negative examples play a crucial role in abstract reasoning tasks.
+Positive examples assist in determining the centroid of the concepts presented in each case, whereas negative examples contribute to delineating the boundaries of these concepts.
+Thus, whether negative examples can minimize the variance $\sigma^{2}$ to the size of the correct distribution becomes crucial.
+If the quality of negative examples is not high enough, the variance cannot be minimized effectively, in other words, it cannot delineate a more precise concept boundary.
 This underscores the significance of establishing both positive and negative examples in abstract reasoning problems.
@@
 VII-B High-quality negative samples lead to more rigorous concept boundaries
-The D2C methodology proposed in this study has established a novel problem boundary for deep networks by means of resampling negative samples and incorporating a fresh loss function.
-Experimental findings tentatively indicate that this boundary is better suited for scenarios characterized by the absence of concept supervision signals, as compared to the boundaries delineated by the original negative samples in Bongard-Logo.
-Nonetheless, the precise tightness of the new boundary remains indeterminate.
-In essence, D2C introduces a new variance to the solution distribution of the network, the magnitude of which relative to the variance of the correct solution distribution is indeterminate.
-In the context of the Bongard-Logo challenge, the variance provided by D2C appears closer to that of the correct solution distribution in the absence of concept supervision signals, as compared to the variance described by the original negative samples.
-Evidently, the efficacy of D2C is contingent upon specific dataset attributes.
-As anticipated, empirical evidence suggests that this approach may be insufficient for delineating precise boundaries within the framework of the RPM problem.
-Contrary to the Bongard-Logo paradigm, RPM problems postulate a shared set of core concepts across distinct cases, ensuring discreteness and stability.
-This shared nature precludes the utility of methods like D2C in addressing RPM challenges.
-Specifically, image properties within RPM cases are rigidly governed by definitions of “rules” or “progressive patterns,” maintaining consistent attributions across all cases.
-The Bongard-Logo problem exhibits varying identifications and determinations of image attribute due to differing clustering bases among cases, thereby augmenting problem complexity and difficulty.
-Incorporating positive samples from other cases as negative examples can, to some extent, delineate concept boundaries between distinct cases, adhering to a cautious approach aimed at minimizing concept infringement.
-However, this strategy proves inapplicable in addressing challenges presented by RAVEN and PGM due to substantial disparities across multiple dimensions.
-In summary, the success of D2C in addressing the Bongard-Logo problem can be attributed to its indirect enhancement of the quality of negative samples within Bongard-Logo cases.
-Conversely, its limitations in tackling the RPM problem can also be ascribed to its insignificant improvement in the quality of negative examples within RPM scenarios.
-To address the limitations inherent in the D2C approach, this paper introduces the D4C (Defense-Driven Discrimination of Distributed Concepts) method.
-The underlying idea of D4C is to generate higher-quality negative samples and samples within abstract reasoning problems.
-This approach aims to delineate a more appropriate problem boundary for the discriminator as compared to the original problem setup in abstract reasoning.
-By leveraging novel mechanisms and strategies, D4C aims to deliver more precise and reliable definitions of concept boundaries and foster further advancements in the realm of abstract reasoning.
+The aforementioned perspective is the core contribution of this paper.
+Since negative examples delineate the boundaries of concepts, we posit that enhancing the quality of negative examples equates to delineating more precise concept boundaries.
+A more precise boundary imply superior discriminator performance.
+Therefore, this paper aims to device a method that can enhance the quality of negative examples in reasoning problems, and D2C is one such approach.
+From the novel perspective on negative samples in reasoning problems, the D2C methodology presented in this study essentially defines a new problem boundary for deep networks.
+However, we suggest that this approach of defining a new concept boundary, specifically the introduction of images from other cases as negative examples for the current case, can make a significant contribution to problems like Bongard-Logo, where there exists concept infringement and erroneous coupling among different cases.
+Conversely, for problems like RPM, where different cases share a common concept set, this method may not contribute effectively.
+In other words, due to the specific characteristics of the Bongard-Logo problem, D2C’s approach has, to some extent, enhanced the quality of negative examples and delineated more precise concept boundaries.
+However, it cannot replicate these improvements in RPM cases, owing to the distinct problem characteristics of RPM.
+In a nutshell, while D2C introduces new negative examples and concept boundaries for reasoning problems, it cannot guarantee the superior quality of these examples or the precision of the new concept boundaries.
+Hence, this paper proposes the D4C (Defense-Driven Discrimination of Distributed Concepts) method, which deliberately seeks out higher-quality negative examples for reasoning problems, thereby defining more accurate concept boundaries.
+The underlying idea of D4C is to generate higher-quality negative samples for abstract reasoning problems, aiming to delineate a more appropriate problem boundary for the discriminator compared to the original problem setup in abstract reasoning.
+By leveraging novel mechanisms and strategies, D4C strives to deliver more precise and reliable definitions of concept boundaries, fostering further advancements in the realm of abstract reasoning.
 VII-C Impelmention of D4C method
@@
 VII-C1 discriminator for Bongard-logo
-Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC and Triple-CFN, achieving notable successes.
-These discriminators seamlessly integrate with the D2C and D3C methodologies proposed in this study, and pertinent experiments have been conducted to explore their combinations.
-These discriminators excel in accurately identifying and classifying Bongard-Logo cases, significantly contributing to the advancement of research in this domain.
-Regarding discriminators capable of handling adversarial tasks, the Wasserstein GAN, a prominent adversarial generative network, poses a noteworthy requirement.
-The discriminator must undertake the task of measuring the distance between samples generated by the generator and real samples.
-This necessitates that the discriminator satisfies Lipschitz continuity to a certain extent.
+Effective Bongard-Logo discriminators abound, with our preceding works, such as PMoC[31] and Triple-CFN[32], achieving notable successes.
+Regarding discriminators capable of handling adversarial tasks, the discriminator must undertake the task of measuring the distance between samples generated by the generator and real samples.
+This necessitates that the discriminator satisfies Lipschitz continuity to a certain extent[36].
+Therefore, in this paper, we choose the spectrum-normalized PMoC and Triple-CFN as discriminators for our D4C, and design generators to refine them.
 In this study, the choice of PMoC, constrained by spectral normalization, as the discriminator for the D4C method on Bongard-logo is motivated by its explicit measurement of the distance between auxiliary and primary group sample distributions.
@@
 Specifically, we employ a standard Transformer encoder to encode a Gaussian distribution based on the representations of the primary samples in a Bongard-Logo case, along with an optimizable vector.
-This encoding approach bears similarity to D3C, wherein we extract the optimizable vector from the attention results and process it through a linear feed-forward layer to obtain the mean and logarithm of variance for the Gaussian distribution.
-Subsequently, we sample multiple representations from this distribution as confusing representations to challenge the discriminator, which must exert maximum effort to counter them.
+This encoding approach bears similarity to D3C, wherein we extract the optimizable vector from the attention results and process it through a linear feed-forward layer to obtain the mean and logarithm of variance for a Gaussian distribution.
+Subsequently, we sample multiple representations from this distribution as confusing representations to challenge the discriminator, which must strive to counter them.
 The generator’s training objective is to generate representations for Bongard-Logo cases that achieve higher scores in the discriminator compared to the representations of the original negative examples in the case.
 Conversely, the discriminator’s training objective is unambiguous: to classify these generated representations as negative examples.
-The aforementioned approach can be seamlessly applied to the PMoC model, providing a direct and effective means of enhancing its performance.
-The alignment between the generator and PMoC is illustrated in Figure 16.
+The aforementioned approach can be seamlessly applied to the PMoC model, providing a straightforward and effective means of enhancing its performance.
+The integration method between the generator and PMoC is illustrated in Figure 13.
 Figure 13: Feedforward process of D4C aligned with PMoC on Bongard-Logo
-In this figure, the “$f_{\theta}(z_{ij}|x_{ij})$” and “$g_{\omega}(p^{\prime}_{i}(z_{ij}|y)|\{z_{ij}\}_{j=1}^{6},z_{ij^{\prime}})$” represent the perceptron and logic calculation modules in PMoC, respectively[31].
-For the generator in Triple-CFN, it is essential to decay the generated representations.
+For the generator in Triple-CFN[32], it is essential to decay the generated representations.
 Specifically, when the generator with the aforementioned structure produces a confusing representation for Triple-CFN, this confusing representation must be weighted and summed with the representations in the auxiliary group.
 After taking the average of the representations in the auxiliary group, it is weighted and summed with the confusing representation in a ratio of $0.9:0.1$ to decay the confusing representation.
-This decay process is necessary due to the mechanism of Triple-CFN, which defines the identity of an image based on the direction of its representation.
-This mechanism makes it easier for a generator based on the Transformer encoder to generate realistic confusing representations, potentially causing Triple-CFN to collapse.
+Due to the mechanism of Triple-CFN, this decay process is deemed necessary, which defines an image’s reasoning identity based on the direction of its representation.
+This mechanism makes it easier for a Transformer Encoder-based generator to generate realistic confusing representations, potentially causing Triple-CFN to collapse.
 Therefore, we have designed the aforementioned decay process.
-The alignment between the generator and Triple-CFN is illustrated in Figure 17.
-In this figure, “$f_{\theta}(z_{ij}|x_{ij})$” represents the encoder in Triple-CFN.
+The integration method between the generator and Triple-CFN is illustrated in Figure 14.
 Figure 14: Feedforward process of D4C aligned with Triple-CFN on Bongard-Logo
@@
 It is anticipated that the application of spectral normalization will enhance the performance of Lico-Net, positioning it as a highly suitable discriminator for the D4C methodology.
-However, Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix.
+Lico-Net will encounter challenges posed by the generator, which produces confusing representations aimed at completing the $3\times 3$ reasoning matrix.
 To address these challenges, Lico-Net will undertake an assessment of the internal consistency within the progressive patterns of the $3\times 3$ reasoning matrix, which are completed by these confusing representations.
 Any inconsistencies or irregularities identified within these representations will be flagged as negative samples, indicating incorrect responses.
-Through this rigorous process, Lico-Net strives to accurately identify and reject perplexing representations, thereby augmenting its overall precision and reliability.
+Through this rigorous process, Lico-Net strives to accurately identify and reject confusing representations, thereby augmenting its overall precision and reliability.
 VII-C4 generator of RPM
@@
 In this paper, we employ a glove-shaped network to simulate the process of sampling from such mixed Gaussian distributions.
-Specifically, we utilize a standard Transformer-Encoder to process the representations of all problem images in an RPM case, alongside $W$ optimizable vectors (where $W$ is a hyperparameter) and a stochastic promoter.
-Notably, when these $W+1$ vectors are arranged together, they resemble the shape of a glove.
+Specifically, we utilize a standard Transformer-Encoder to process the representations of all images in the statement of an RPM case and $N$ pattern vectors from the statement encoded by Lico-Net, alongside $W$ optimizable vectors (where $W$ is a hyperparameter) and a stochastic promoter.
 The promoter is a vector sampled from a multivariate Gaussian distribution, wherein the mean is optimizable while the variance is fixed at 1.
+The Transformer-Encoder needs to process $N+W+1+8$ vectors in total.
+Notably, when these $W$ vectors and the promoter are arranged together, they resemble the shape of a glove.
 It is worth noting that we have devised a novel positional encoding embedding approach for the Transformer-Encoder, which serves as the backbone of our RPM generator.
 This encoding aims to restore the $3\times 3$ image arrangement characteristic of the RPM problem matrix.
-For the PGM (Progressive Graph Matching) problem, our positional encoding scheme assigns unique encodings to images located on the diagonal and upper triangular regions within the problem matrix of a given PGM case.
+For the PGM problem, our positional encoding scheme assigns unique encodings to images located on the diagonal and upper triangular regions within the problem matrix of a given PGM case.
 Correspondingly, the positional encodings for images in the lower triangular region are symmetrical with respect to the diagonal, mirroring the encodings of their counterparts in the upper triangular region.
 However, it should be acknowledged that this paper does not demonstrate the superiority of D4C on the relatively less challenging RAVEN dataset.
-Moving on to the main topic, after obtaining attention results from the Transformer-Encoder, we extract $W$ optimizable vectors from these results and process them into $W$ distributions using a linear layer, as mentioned earlier.
-Following the extraction of the promoter, it is processed into a probability output using a linear layer and hard-gumbel-softmax activation, yielding highlighted probabilities for the $W$ distributions.
+Moving on to the main topic, after obtaining attention results from the Transformer-Encoder, we extract $W$ optimizable vectors from these results and process them into $W$ distributions using a linear layer, like D3C method.
+Following the extraction of the promoter, it is processed into a probability output using a linear layer and hard Gumbel-Softmax activation, yielding highlighted probabilities for the $W$ distributions.
 Subsequently, we sample representations, referred to as confusing representations, from the highlighted distributions each time.
 In the technical implementation, we employ the reparameterization trick to draw samples from each distribution, resulting in $W$ sampling outcomes.
-We then perform a weighted sum of these sampling results using the hard-highlighted probabilities processed by the hard-gumbel-softmax, thus completing the entire mixed-Gaussian sampling process.
-The methodology outlined in this paper for modeling and training the network is coined as ”D4C glove-train”.
-A more detailed illustration of the forward process is provided in the accompanying figure referenced as 18.
+We then perform a weighted sum of these sampling results using the hard-highlighted probabilities processed by the hard Gumbel-Softmax, thus completing the entire mixed-Gaussian sampling process.
+The methodology outlined in this section is coined as “D4C glove-train”.
+The forward process of D4C on Lico-Net is shown in the figure 15.
+In figure 15, “pattern extractor” is illustrated in figure 6 and the process of “pattern scoring” is shown in figure 7.
 The training objective of the glove-shaped generator remains consistent: to produce confusing representations that achieve higher scores in the discriminator compared to the original negative examples provided by the dataset.
@@
 Figure 15: Feedforward process of D4C on RPM problem
-Finally, in the realm of abstract reasoning problems, particularly those involving generating answer samples, this paper believes that the D4C approach can also make valuable contributions.
-This opens up new avenues of exploration for future work, presenting challenges that are yet to be tackled.
+Overall, compared to traditional methods, D4C is a fundamentally different approach.
+Previous studies on solving abstract reasoning problems[23, 21, 28, 32] have primarily focused on constructing solvers that structurally align with the inductive bias in the problems, including our Lico-Net.
+However, our D4C reconsiders the significance of negative examples in reasoning problems, and designs a solution based on this introspection.
+As a bonus, the glove-shaped generator in D4C provides suggestions for simulating the random sampling process of a mixed Gaussian distribution.
 VIII Experiment
@@
 VIII-A Experiment on RAVEN
-Initially, we conducted experiments with Lico-Net on both RAVEN[15] and I-RAVEN[37] datasets to demonstrate its inherent potential.
-To ensure a fair comparison between excellent models, all experimental settings and equipment in this study were kept consistent with those of RS-Tran [28] and Triple-CFN [32], including data volume, optimizer parameters, data augmentation methods (such as reducing the RAVEN image size to $80\times 80$), batch size, and other relevant hyperparameters.
+Initially, we tested Lico-Net on RAVEN and I-RAVEN datasets to show its potential.
+For a fair comparison, we used the same settings and equipment as RS-Tran[28] and Triple-CFN[32], including data volume, optimizer parameters, data augmentation, batch size, and other hyperparameters.
 TABLE II: Reasoning Accuracies on RAVEN and I-RAVEN.
@@
 Lico-Net 99.5/99.7 100.0/100.0 99.7/99.8 98.2/99.2 99.9/100.0 99.9/100.0 99.9/99.9 98.8/99.2
-An interesting observation from the Table II indicates a consensus among existing excellent RPM solvers, including Lico-Net.
-Specifically, these solvers exhibit differing values of reasoning accuracy when tackling the RAVEN [15] and I-RAVEN [37] datasets.
+An interesting observation from the Table II indicates a commonality among existing excellent RPM solvers, including Lico-Net.
+Specifically, these solvers exhibit varying reasoning accuracy when tackling the RAVEN [15] and I-RAVEN [37] datasets.
 The I-RAVEN dataset was designed to address certain limitations in the option pool configuration of RAVEN by introducing improvements to its negative example setup.
 These enhancements include, but are not limited to, increasing the number of attributes that are altered in the negative examples and augmenting the magnitude of attribute deviations within them.
-However, from a human perspective, such additional setups do not alter the difficulty level of reasoning, and thus do not make I-RAVEN any harder or easier compared to RAVEN.
-This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the superior quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN.
-Such phenomenons echo the D4C’s spirit.
+However, from a human perspective, since the setup of positive examples remains unchanged, such additional setups do not alter the difficulty of reasoning, making I-RAVEN neither more nor less challenging than RAVEN.
+This paper speculates that the disparate outcomes achieved by these models on RAVEN and I-RAVEN could potentially be attributed to the higher quality of negative examples present in the I-RAVEN dataset compared to those in RAVEN.
+Such phenomena echo the D4C’s spirit.
 VIII-B Experiment on Bongard-Logo
-We conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN [32] as backbones for D2C, D3C(-cos), and D4C and using PMoC [31] as backbone for D2C and D4C.
+We conducted experiments on Bongard-Logo using ResNet50 and Triple-CFN [32] as backbone for D2C, D3C(-cos), and D4C and using PMoC [31] as backbone for D2C and D4C.
 PMoC, a probabilistic labeling model, does not meet the requirements of D3C(-cos), as D3C(-cos) demands a backbone in the form of $f(z|x)$.
 The experimental results are presented in the table III.
-Upon observing the results presented in the table III, we found that D2C, D3C, and D3C-cos made considerable contributions in the Bongard-Logo task.
-As can be seen in the table III, replacing D3C with D3C-cos to integrate with ResNet50 would not significantly affect performance.
-We argue that, from a results-oriented perspective, D3C-cos can be considered as a effective alternative to D3C.
-Notably, we did not conduct an experiment with Triple-CFN+D3C and Triple-CFN+D2C+D3C.
-This is because the large number of Triple-CFN parameters makes integrating the D3C approach impractical on our current experimental equipment.
-This was the driving force that led us to propose D3C-cos as an alternative to D3C in this paper.
+Observing table III, D2C and D3C(-cos) notably contributed to the Bongard-Logo task.Replacing D3C with D3C-cos when integrated with ResNet50 has minimal impact on performance.
+From the result-oriented perspective, D3C-cos is a qualified alternative for D3C.
+We didn’t test Triple-CFN with D3C or D2C+D3C due to practical limitations from our equipment, which led to the proposal of D3C-cos.
+Furthermore, D4C has significantly improved Triple-CFN, ResNet50, and PMoC, often matching the combined effects of D2C and D3C.
+However, on Bongard-Logo’s FF concept, D4C’s improvements on Triple-CFN and ResNet50 were less than ideal.
+This might be because D4C was designed for probabilistic labeling models, whereas Triple-CFN and ResNet50 do not follow this probabilistic approach.
+A decay mechanism, as a compromise for non-probabilistic labeling models, was added when integrating D4C with these models.
+Experimental results suggest this mechanism needs further refinement.
+Future work should explore better designs to adapt D4C to non-probabilistic labeling models like Triple-CFN.
+Conversely, PMoC [31], a probabilistic labeling model for Bongard-Logo, benefits more from D4C due to its scheme alignment.
 TABLE III: Reasoning Accuracies of models on Bongard-logo.
@@
 PMoC+D4C 93.0 92.6 98.2 78.0 77.5
-Additionally, D4C has imparted varying degrees of improvement to Triple-CFN, ResNet50, and PMoC, with its effectiveness rivaling that of the combined application of D2C and D3C.
-However, D4C did not yield an ideal improvement on the FF concept for the Bongard-logo problem when applied to Triple-CFN and ResNet50.
-This study hypothesizes that the suboptimal performance may be attributed to the fact that D4C was specifically designed and tailored for discriminative models with a probabilistic labeling scheme.
-Conversely, Triple-CFN and ResNet50 do not inherently adhere to such a framework.
-It has been mentioned that when integrating D4C with Triple-CFN and ResNet50, a decay mechanism was incorporated.
-This mechanism is a compromise made by D4C to accommodate the non-probabilistic labeling scheme of these models.
-From the perspective of experimental results, this decay mechanism does not appear to be a fully developed mechanism.
-Exploring more advanced designs to facilitate the adaptation of D4C to non-probabilistic labeling models, such as Triple-CFN, deserves attention as future work.
-In contrast to Triple-CFN, PMoC [31] serves as an intuitive probabilistic labeling model for the Bongard-Logo problem.
-Correspondingly, D4C provides a more significant enhancement when applied to PMoC, owing to its scheme.
-D4C demonstrated improvement and contribution, validating the effectiveness of defining tighter problem boundaries for the model with a probabilistic labeling scheme.
-To integrate D4C with Triple-CFN, we devised a representation decay mechanism that, despite its imperfections, is sufficiently effective.
-This approach warrants further investigation and development in future research.
 Finally, this paper notes that augmentation experiments were conducted for both PMoC [31] and Triple-CFN [32].
 These experiments involved expanding the data volume of the Bongard-Logo dataset using data augmentation techniques, specifically including rotation and flipping.
-It should be emphasized that these data augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [31].
+It should be emphasized that these augmentation techniques do not explicitly alter the diversity of concepts or the proportion between concepts inherent in the Bongard-Logo dataset [31].
 Rather, the aim of these experiments was to investigate the potential of the models with a larger dataset.
-Since PMoC posits that Transformer-based visual processing models require substantial data as a foundation, a point also echoed by ViT [17] and RS-Tran [28].
+Since PMoC posits that Transformer-based visual processing models require substantial data as a foundation, a point that is also echoed by ViT [17] and RS-Tran [28].
 Furthermore, both PMoC and Triple-CFN agree that the utilization of these data augmentation methods serves to enrich the expressions of concepts at the pixel level, thereby challenging the network’s abilities to summarize, derive, accommodate, and invert concepts [31, 32].
-This paper conducted experiments to verify the effectiveness of D4C with the expanded dataset.
-The results are recorded in Table IV.
+To verify the effectiveness of D4C on the augmented dataset, we conducted experiments, and the results are presented in Table IV.
 TABLE IV: Reasoning Accuracies of D4C on Augmented Bongard-logo.
@@
 From the experimental results presented in Table V, several notable phenomena emerge.
-In Triple-CFN, both progressive patterns and minimal reasoning units are positioned at the same logical level [32].
-However, a significant design flaw arises from the mismatch in the logical hierarchy between these patterns and units.
-More specifically, progressive patterns often require the combined expression of multiple minimal reasoning units.
-Despite this limitation, Triple-CFN still stands as an impressive model.
-Naturally, leveraging the inductive bias inherent to the RPM problem as the guideline in its design, Lico-Net has achieved even more remarkable results than Triple-CFN in the RPM challenges.
-The combination of Lico-Net and D2C, and the combination of Triple-CFN and D2C, did not bring significant improvement, with the enhancement only reflected in the percentile of reasoning accuracy.
-This phenomenon was anticipated in this paper, as we believe it is due to the difference between RPM and Bongard-Logo problems.
-In RPM problems, different cases share the same concept space and attribute recognition method.
-The D2C approach, designed to address issues of concept interference and erroneous coupling among cases, is not applicable to RPM problems.
-However, the combination of D4C and D4C glove-train significantly improved Lico-Net’s performance.
-Additionally, the glove-shaped generator proved to be more suitable for generating negative representations of RPM problems for Lico-Net and enabling it to achieve further refinement.
-The core innovation of this paper, D4C, has yielded promising results on both the RPM and Bongard-Logo problems, demonstrating the significant contribution of this work.
+Triple-CFN[32] positions progressive patterns and minimal reasoning units at the same logical level, leading to a design flaw due to their mismatched logic.
+Despite this, Triple-CFN remains an impressive model.
+By leveraging the RPM problem’s accurate inductive bias in its design, Lico-Net has surpassed Triple-CFN in RPM challenges.
+The combinations of Lico-Net+D2C and Triple-CFN+D2C showed minimal improvement, limited to reasoning accuracy percentile.
+This was expected, as RPM and Bongard-Logo differ.
+RPM cases share a concept space and attribute recognition, rendering D2C, designed for concept interference and erroneous coupling issues, ineffective for RPM.
+The combination of D4C and D4C glove-train notably enhanced Lico-Net’s performance.
+D4C, our core contribution, showed promising results on both RPM and Bongard-Logo problems, highlighting the significance of this work.
 The combination of Lico-Net and D3C is noteworthy.
-To validate the theme of D3C, we applied both D3C and D3C-cos to Lico-Net.
+To validate D3C method, we applied both D3C and D3C-cos to Lico-Net.
 Through experimentation, we found that D3C and D3C-cos brought improvements to Lico-Net.
 To demonstrate the effectiveness of D3C’s approach to describing abstract concepts through distributions, this paper conducted ablation experiments using “Lico-Net+D3C-”.
-Unlike “Lico-Net+D3C”, which encodes the description of image progressive patterns into distributions using a Transformer-Encoder, “Lico-Net+D3C-” encodes them into vectors.
-To simplify, instead of dividing the encoded representation into mu and logvar, we directly utilize the representation as a vector.
-Furthermore, the cosine distance between progressive pattern vectors $\{\overline{P}_{nm}\}$ and the vectorized pattern description in Lico-Net is constrained using infoNCE.
+Unlike “Lico-Net+D3C”, “Lico-Net+D3C-” encodes Meta data into vectors.
 In essence, the Lico-Net+D3C- can be interpreted as fixing the variance of the distributions established for all concepts in Lico-Net+D3C-cos to 0.
-The structure of Lico-Net+D3C- is illustrate in figure 19.
-Figure 19: Feedforward process of “Lico-Net+D3C-
 TABLE VI: Progressive Pattern Reasoning Accuracies and of models on PGM.
@@
 Lico-Net+D3C-cos 99.7 99.9 99.4
-The D3C’s reasoning accuracys of progressive pattern are recorded in Table VI, highlighting the interpretability expressed by Lico-Net+D3C(-cos).
-It must be acknowledged that the experiment of applying D3C, which is based on the Sinkhorn distance, to Lico-Net was not conducted to its fullest extent in this paper.
-Once the experimental results of Lico-Net+D3C were sufficient to illustrate our point, we halted the model training.
-This was due to the significant computational resources required, even with the experimental equipment used in this paper: four A100s graphics cards.
-The training involved numerous epochs and tedious individual training epochs.
-This was the intended purpose behind the design of D3C-cos as a replacement for D3C, as the power of Zeus needs to be conveyed by Hermes.
-Despite achieving remarkable results in terms of reasoning accuracy and interpretability of progressive patterns with Lico-Net+D3C or D3C-cos, it is noteworthy that the improvement over previous works involving Meta data as additional auxiliary labels is not particularly striking.
-The possible reason for this is that the introduction of Meta data significantly reduces the difficulty of the RPM problem Moreover, it is anticipated that a greater abundance of supervisory signals would correlate with enhanced intelligent performance.
-However, Lico-Net+D3C still demonstrates distinct advantages that surpass other methods involving metadata.
-Specifically, Tranclip [28] utilizes metadata to pre-train a set of warm-start parameters for RS-Tran [28].
-In other words, Tranclip fails to simultaneously endow RS-Tran with both interpretability of progressive patterns and high reasoning accuracy.
-Meta Triple-CFN [32] also relies on its specially designed Re-space Layer to enhance both its interpretability and reasoning accuracy.
+Table VI records D3C’s reasoning accuracies for progressive patterns, emphasizing Lico-Net+D3C(-cos)’s interpretability.
+Due to resource constraints, this paper did not fully explore the application of Sinkhorn-based D3C to Lico-Net.
+We stopped training once results illustrated our point, given the significant computational demands even with four A100s.
+In terms of results, D3C-cos can serve as a stand of D3C.
+Despite impressive reasoning accuracy and interpretability with Lico-Net+D3C(-cos), the improvement over previous works using Meta data as auxiliary labels isn’t significant.
+This might be because the inclusion of Meta data lowers the RPM problem’s reasoning difficulty.
+More supervisory signals typically enhance the intelligent performance of the system.
+Lico-Net+D3C still shows distinct advantages over other Meta data-based methods.
+Specifically, Tranclip[28] utilizes Meta data to pre-train a set of warm-start parameters for RS-Tran[28], achieving promising reasoning accuracy for RS-Tran.
+However, Tranclip fails to simultaneously endow RS-Tran with both the the interpretability of progressive patterns and high reasoning accuracy.
+Meta Triple-CFN[32] also relies on a specially designed Re-space Layer[32] to enhance both its interpretability and reasoning accuracy.
 Similarly, the Re-Space Layer requires a fully trained Meta Triple-CFN to provide warm-start parameters [32].
-In contrast, Lico-Net operates effectively without the necessity of a Re-space layer or a warm-start process, thereby simplifying the training procedure to a certain extent.
-This indicates that Lico-Net and D3C are more direct and effective, eliminating the need for indirect approaches.
-This ”directness” reflects a more coherent and self-consistent logic of Lico-Net and D3C compared to other methods incorporating metadata.
+In contrast, Lico-Net+D3C operates effectively without these extras, showing a more direct, coherent and self-consistent logic.
 This paper argues that this is not an insignificant contribution, as warm-start behavior may introduce unexpected instabilities, which are not a concern for Lico-Net and D3C.
-For example, when Tranclip [28] uses Meta data to produce warm-start parameters for RS-Tran, early stopping is required [28].
-Neglecting this step may hinder Tranclip from improving or potentially worsen the reasoning performance of RS-Tran.
-The Rs-space Layer [32] relies on a well-trained Meta Triple-CFN to provide warm-start parameters [32], a process that demands considerable training time and epochs.
-Inadequate training of these warm-start parameters can expose Meta Triple-CFN to risks akin to those faced by RS-Tran.
-The stopping timing for the preparation of both warm-start parameters requires manual determination and control, which brings numerous potential issues.
-In contrast, Lico-Net effectively mitigates these potential issues.
-VIII-D Ablation experiment on D2C and D4C
-In the final of this section, we devised an experiment to validate the D2C and D4C approach.
-Utilizing a 500-step DDPM model, we generated a substantial quantity of Bongard-Logo data, specifically 3,180 Bongard-Logo images.
-According to the DDPM theory, these images belong to the same distribution as the real data, albeit with unknown semantics and concepts.
-During the training of ResNet50 and Triple-CFN, for each feedforward pass, we introduced 14 randomly selected DDPM-generated images as auxiliary group examples for the Bongard-Logo cases.
-The experimental results were recorded in “D2C-” item of the Table VII.
-TABLE VII: Reasoning Accuracies of models on Bongard-logo.
-Accuracy(%)
-Model Train FF BA CM NV
-ResNet50 (infoNCE baseline) 92.0 88.9 98.0 76.1 76.0
-Triple-CFN 93.2 92.0 98.2 78.0 78.1
-ResNet50+D2C- 91.6 90.0 97.9 76.3 76.1
-Triple-CFN+D2C- 93.0 92.1 98.2 77.9 78.2
-It can be observed from the experiment results recorded in Table VII that introducing generated samples as negative examples is difficult to fundamentally alter the network’s performance, as such behavior struggles to re-delineate the concepts.
-DDPM-generated data, sharing the same distribution as real data but with inherent uncertainties of concept and attribution, are limited in contributing to the delineation of concept boundaries.
-D4C posits that the accuracy achieved by the discriminator in abstract reasoning problems is closely related to the quality of the positive and negative examples set in the task.
-The established examples serve to delineate the concept boundaries for abstract reasoning challenges.
-D2C, while attempting to define new concept boundaries for such problems, represents an imperfect methodology.
-Conversely, D4C emerges as a relatively more comprehensive approach.
-This ablation experiment aims to verify that, within the context of Bongard-Logo problems, the provision of meaningless negative examples—examples that do not meaningfully contribute to defining the boundaries of reasoning tasks—fails to enhance the discriminator’s effectiveness.
-This finding echoes similar observations made in the RAVEN dataset, reinforcing the significance and rationale behind the design of D4C.
-Furthermore, it reaffirms D4C’s assertion regarding the importance of high-quality positive and negative examples in abstract reasoning tasks.
-It also provides indirect evidence that the approaches of D2C and D4C extend beyond merely introducing additional negative examples.
-Instead, they aim to introduce negative examples that better delineate the boundaries of the problem.
+For instance, Tranclip requires early stopping when using Meta data for training RS-Tran’s warm-start parameters[28], and neglecting this can hinder RS-Tran’s performance.
+Similarly, The Re-space Layer [32] relies on a well-trained Meta Triple-CFN to provide warm-start parameters [32], a process that demands considerable training time and epochs, with inadequate training posing risks.
+Both methods require manual control, whereas Lico-Net+D3C mitigates these issues.
 IX Conclusion
-In summary, D3C advocates for representing concepts in abstract reasoning problems using distributions.
-It posits that, compared to vectors, distributions are better suited to capture the high-dimensionality, abstractness, and complexity inherent in human concepts, and experimental evidence underscores the superiority of this approach.
-Additionally, D2C presents a novel approach to tackling Bongard-Logo and RPM problems by resampling training data and introducing new loss function terms, effectively outlining new boundaries for solutions to these challenges.
-However, the extent to which these boundaries adapt to the specific problems remains uncertain.
-D4C builds upon the principles of both D2C and D3C, exploring the roles of positive and negative examples in abstract reasoning problems from a perspective on distribution.
-Specifically, positive examples are utilized to depict the mean for deep networks, while negative examples contribute to describing the variance.
-Based on this understanding, D4C employs an adversarial approach, pitting the generator against the discriminator in repeated challenges.
-The generator continually produces negative examples that push the boundaries of the problem, while the discriminator refines its skills through repeated defense.
-This iterative process not only enhances the discriminator’s ability to distinguish between positive and negative examples but also improves the generator’s capacity to generate more sophisticated negative examples.
-Ultimately, this adversarial framework promotes the evolution of both components, leading to more robust and reliable solutions in abstract reasoning tasks.
-For future, the D4C method presents a novel approach for abstract reasoning problems and even other discriminative tasks.
+D4C, building on D2C and D3C principles, explores the roles of positive and negative examples in abstract reasoning from a distribution perspective.
+Positive examples depict the mean for solution distribution, while negative ones limit variance.
+This seems to reflect a despairing fact: the upper limit of accuracy achieved by the discriminator may be constrained by the quality of positive and negative examples.
+Therefore, D4C emerged to address this phenomenon.
+D4C adopts an adversarial approach, where the generator creates challenging negative examples and the discriminator refines its skills through defense.
+This iterative process enhances both components, resulting in remarkable solutions for abstract reasoning tasks.
+This paper does not advocate for the interpretability of the samples generated by the D4C generator; they are solely used to improve the performance of the discriminator.
+For future, the D4C method presents a unique perspective and approach for abstract reasoning problems and even other discriminative tasks.
 Its exploration of the identities of positive and negative examples may provide insights for the development of other works.